A Complete Bibliography of Publications in Computer Systems Science and Engineering and International

Journal of Computer Systems Science and Engineering

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 April 2017 Version 2.03

Title word cross-reference

(t, n) [WLY^{+}14]. 3 [TC05]. Δ [BPRS98]. k [ASDOK03, KC10, SR03, YLK11, ZHZ^{+}15]. l [YLK11]. N [CT93, Li98, KWL07]. N^{2/3} × N^{2/3} [PS95]. r [Rav92, Rav95].

-ary [ASDOK03]. -causal [BPRS98]. -coteries [SR03]. -dimensional [Rav92, Rav95]. -gram [KWL07, KWL07]. -Gram/2L-approximation [KWL07]. -nearest [ZHZ^{+}15]. -NN [KC10, YLK11]. -stage [CT93]. -version [Li98].

.NET [TP09].

1 [HMF93, IMA14, RMDF91].

2.0 [RFMP07]. 2006 [FHFL07]. 2014 [LL15]. 2015 [LK16]. 264 [BDPV90]. 2L-approximation [KWL07].

4G [DD11, LCWC14]. 4G/LTE [LCWC14].

5 [YWD00].

802.11 [MN14]. 88 [HS88].

95 [JPMAB00].

A* [IST^{+}13]. A-star [IST^{+}13]. AAM [LKK14]. AAM-based [LKK14]. abilities [GZ06]. ABR [ACA99, TZL96]. abstract [TH94]. Access [Gai87, SSS01, AD04, CMRR02, Cra06, DKP95, GS01, Hsu08, KFW04, KA97, LY92, LWHS06, MCMM95, QYCG10, SA03, WD90, WWH95, WC01]. accesses [JSM96]. accessible [BGRV99]. accessing [HC90]. Accountable [LdOS13]. accuracy [KL00, MK11]. Accurate [PRPS13]. achieve [WWDL09]. Achieving [CD94, SN09, Sah95]. acoustic [GFH^{+}10, LCWC12]. Acquisition [MFU02]. across [MOEMK^{+}15]. Act [LBLB13]. action [BWB^{+}01, RZ01]. Active [PV96, GVCV13, LSR^{+}92, XFZ09]. activeshape [CGR^{+}09]. activities [SN07]. activity [Bed88, GCJP03, KS97]. Ad [WCLC04, KTV14, LHC08, PYHO04, SSB^{+}07, WC08b, WWS16]. ad-hoc [KTV14, WWS16]. Ada [BST98, BWB^{+}01, JPMAB00, NCMH00, VG00]. Adaptability [SV13]. adaptable [SARAL05]. Adapter [Bow86]. Adaptive [Cha08, CH06, HS96, LY92, SP16, WC08a, ZYM96, CCZ03, FCGC13, GFH^{+}10, MJRIV14, OSZ95, QKSN09, ZCG^{+}15]. address [HC90, SVL02, SL02]. Adjacency [CZ09]. adjudication [Rom02]. administration [CM04]. admission [BBMC98, CH98, RW11]. adopting [LLHG06]. AdOrBAC [CM04]. Advanced [FVD13, LM09, HOGS11]. affected [SED^{+}99]. affine [Che93]. Affordance [BSR11]. against [SF02, TV08]. age [HABJ05]. Agent [BDKD12, GL06, TC06, GKK03, HC08, HBL16, KHA06b, NCB06, RJS06, XQZ11, YC13]. Agent-based [BDKD12, XQZ11, YC13]. Agent-oriented [TC06]. Agents [GZ06, CCSS06, Hog04, LP03, QKZ^{+}07, SIGC15, YHLC13]. aggregate [KW10, MK10]. aggregation [Kam93, LSBW14, SFLP99, XLSM10]. aggregator [SPB11, SPBK12]. agile [CCSS06, CCSS06]. Agreement [DGGS88, DGGS89, AJTT15, VK08, WL91]. agreements [Hog04]. AHS [vRB11]. aided [AK92]. alarm [BLY10]. algebra [CMM91, ZT03]. algebra-based [CMM91].

Algorithm

[DGGS88, SA86, SE85, Agu03, ADMB15, Ang16, BYX16, CSXC11, CHCL90, CY94, Cha02, CL07, CPCB11, DD90, Din04, DG01b, GDK88, HN94, HA08, Hui00, IST^{+}13, JDG92, JM99, KS00, KW02, Lee04, Lee09, LNS^{+}07, OSZ95, OSZZ97, PXQ08, PN09, PCB10, PA10, PRB13, QKZ^{+}07, QY15, RHS^{+}14, SKR08, Sri91, SBK^{+}92, SB96a, TZC09, WL91, WLXD15, WW01, WF93, XQZ11, Yan02, YSS^{+}16, YET94, ZZWD15]. Algorithmic [Amm90, FH13, SEuH96]. Algorithms [DGGS89, GET10, Kap92, KP90, ASDOK03, ABRW93, BXST12, BSS90, DFT97, DS97, FCGC13, GFH^{+}10, Hof93, JIB03, LS97, MS96, MZ03, MCMM95, ML09a, ML09b, Pap94, PTX^{+}09, PC13, RMB91, RP90, RD92, SG90, SSC97, SM90, SS93, TR02b, TRT04, VKAR15, Yan00, YA97]. Allocating [QAS91]. allocation [CCHL03, CHE12, Hsu94, HWL97, IS97, JWH^{+}15, KC94, jKsJdChK12, LM91, LJL^{+}12, LYL^{+}12b, SKR08, TX16, WL01, ZZZZ15, Zhu96]. alternating [CP90, MVA09]. alternative [SLPK02]. ALU [RC98]. ambient [Jon03, Kri08]. ambiguity [NXG^{+}16]. among [MGOB15]. AMS [PONA11]. analyses [Tho97]. Analysis [ASDOK03, BBKT87, DYB91, GDK88, IYD87, KS97, LC96b, SHGA13, SE85, VB87, Wan97, AHH^{+}16, BqQHj^{+}15, Bat00, BSMB14, BST98, BCL^{+}01, BNSM01, Car14, CI94, DD98, FG97, FSA97, FM10, GLT05, GCM14, GQW^{+}14, GS01, HWL97, HZS12, KNP16, KTK14, KKEG^{+}09, KVG08, KBB09, LMI90, LP93, LY96, LCH92, MM15, MOK03, Mor07, MN98, RJS06, Red88, SMSJ10, SE07, SW12, SVN96, SSC91, SBK^{+}92, SPB11, SDN14, TTD10, TR06, Vla07, Wan03, XQZ11, XC15, YQ95, YWS06, YSS^{+}16, ZCL03]. Analytic [Lam09, QA93]. Analytical [FSA97, CQ96, Li99]. Angle [IB10].

Annealing [BBKT87, Din04, RP90]. annotation [CC15, CC16, WKT07]. anomalies [KKP11]. anomalous [CP15]. Anomaly [HFS^{+}08]. anonymity [LC00a]. anonymous [JC02, LHC08, WH97]. Answering [HF08]. Anti [Nan94]. Anti-code-disjoint [Nan94]. Anticipatory [TF06]. Application [AD13, AD14, AM00, BBKT87, TM94, TH94, AR04, CSXC11, CS15, Kof05, MLLS94, PHPd98, RS88, The94, ZCK11]. application-layer [CSXC11]. Applications [FCHJ05, SKV12, XSZ16, ALP99, CL92, CFLZ06, DGGE88, GMC^{+}13, GD06, JPMAB00, JOR^{+}00, KH96, KS04, LS97, MC95, NMMS02, PMO16, SN09, SSB^{+}07, TJS15, TX16, Toh94, WW01, Yan93, Ye15, ZEZF11]. applied [BqQHj^{+}15, HA08]. Applying [Cra06, LB03]. Approach [EGH^{+}86, PN09, WDF87, AST99, AR92, Amm90, And98, AdPT06, CZT^{+}16, CPV90, CH06, CY90, DD11, FM14, GW01, HK98, JIB03, KKK^{+}02, KKLK02, LASS00, LYA06, LYC02, MM15, MMD99, NK03, NKWK07, PKC07, PV96, QA93, RMDF91, RDRLG^{+}15, RFMP07, RSDD10, SHGA13, Sha09, SF11, SSC97, SLA^{+}11, SD99, WBA94, WCL15, WWDL09, YWL^{+}03, vdA01]. approaches [AP16, HM04, KLK01, KI09]. Approximate [Tho97, Dan07, KWL07, MLK15, PVRM11]. approximating [Li99]. Approximation [PTX^{+}09, CO91, KWL07, MK10]. Arabic [HHCASEQ11]. Arbitrary [IB10]. Architectural [CQ96, KS99, KBB09]. Architecture [HP87, HM87, LBLB13, SZ91, CCdF^{+}07, CPT08, DDL91, GH11, HSS^{+}11, KG07, KA97, LYW93, zLTC95, LS97, MG95, NPR10, PXY93, RSP91, SVN96, SS99, TR02a, TC06]. architecture-level [GH11]. Architectures [IYD87, DMP98, GD06, HC90, LS94, MRSW98, WMFK95]. Area [BG87, LMI90]. arecanut [SDN14]. arm [TF06]. arrangement [Wan03]. Array [CLO92, Sch88, TX16]. Arrays [BSK85, BSS90, Che93, KR92, KP90, OSZ95, SG90, YWD00]. art [THDC08, FM14]. Artificial [ELG00, ZL15]. ary [ASDOK03]. Ascending [XLH16]. ASLM [MHHP92]. aspects [Pos98]. Assessing [ÇG11, CPCB11, KTJ14, PCB10]. assessment [ADSP12, GP90, LH98, NK03]. assignment [ELG00, jKsJdChK12, OSZZ97, WC01]. assignments [BD96, WF92]. assisted [YDR^{+}09]. assistive [RCNL15]. associated [CHL^{+}12, KSH96, ZLL^{+}15]. Association [TRLD10, SL14, XLW11]. associations [YLQW13]. associative [MHH93]. associativity [ZS06]. assumption [WF92]. assumptions [LLHG06]. assurance [War07]. assured [SLPK02]. astronomical [VSZMCV08]. Asymmetrically [WSL93]. Asymmetrically-initiated [WSL93]. asynchronous [ADSP12, DAFG95]. Atlas [GPR10]. ATM [And98, BBMC98, BGM98, BVCW96, CCZ03, CH97, DD98, FG97, FSA97, GMF96, KS97, KA97, LL96, LY96, LH97, LH02, LD96, MA97, RM98, SB96b, TH97, TZL96, WL01, XM97, ZYM96]. Atomic [JSM96, BWB^{+}01, Ray00, RZ01]. attack [GKK03]. Attacks [ZXG08, Bul11, SF02, TV08]. Attempto [Gun93]. attenuation [GFZ^{+}16]. attributes [ADSP12, SHGA13]. auctions [Mor07, XC07]. audio [KKLK02]. augmented [BDKD12]. Augmenting [SK92]. authenticated [HTJ08, LHC08, TJ01]. authentication [AAKD09, Cer04, FM14, HDPC13, LCX08, PA10, SN07, VK08]. authorisation [SN07]. Authorization [SW06]. authorship [FMS^{+}11]. Auto [ZEZF11]. Auto-scaled [ZEZF11]. Automata [YWL^{+}03, Ang16, Buc00]. Automated [Hal09, BST98, WKT07]. Automatic [BBKT87, EG97, JTN95, Vla07, YLZ11, CC15, CC16, GLT05]. Automation [EGH^{+}86]. Autonomous [MMI85, ASH12, Hog04, Joh96]. Availability [Ano87, CY94]. avionics [Joh96]. Avoidance [DGHE88, Sri91, WYT01]. aware [BLY10, BWG06, Dam04, GMN98, GKS06, MMH16, PYHO04]. axioms [FL16].

B2B [YC13]. B2C [TBEH04]. BAC [CM04]. background [QY15, SB96b]. backhauling [Car15]. backup [WCLC04]. backward [BDMV97, HCD08a, HCD08b]. backward-and-forward [HCD08a, HCD08b]. balancing [BM97, DMP98, Kar91, LWL07, LJ97, RSG96, SCL97, Tan97, WSL93]. bandwidth [FCGC13, jKsJdChK12, KW10, WL01]. banks [HA08]. banyan [PLL00]. Based [HCY87, HYY88, VB87, AST99, Agu03, AGP16, AR04, AD04, AAKD09, AGT98, BCHR01, BLY10, BZLX16, BYX16, BCL^{+}01, BDKD12, CSQW14, CSB16, CS97, CL00, CMZ10, CHZ11, CCL11, CHL^{+}12, CZT^{+}16, Cho11, CPV90, Cra06, CMM91, DR09, DJ96, DDL91, DG01b, ET10, FAL^{+}01, FXZ16, FM14, GVCV13, GH11, GR94, GMC^{+}13, GA14, GP90, HC08, HCD08a, HCD08b, HGR12, Hel13, HFS^{+}08, HWL97, HBL16, HZS12, HCL^{+}06, HDCH10, IST^{+}13, IMA14, JM99, JWH^{+}15, JLL01, KTJ14, KKY04, KTK14, Kat05, KL00, Kat92, KBG^{+}94, KKLK02, KJKK07, KHL^{+}08, KSL12, KVR03, KVR04, LHW04, LKK14, LJL^{+}12, LYA06, LYL^{+}12b, LWL93, MS96, MM15, Mor07, NKWK07, NHK14, Pap98, PKC07, PRSM01, QWK14, QLL14, QL13, QL15, QKZ^{+}07, RCNL15, RES^{+}16a, RMR00, Ray00, Ray02, SF11, SP16, SD99, SCL97, SS13, SPBK12, TJS15]. based [Tam08, TY97, THT98, TZC09, TRAR02, TMNL08, TZL96, WD90, WBA94, Wan97, Wan03, WCLC04, WLXD15, WCL15, WYLW16, WWS16, WBY^{+}13, WWH95, WY02, XZ11, XQZ11, YQ95, YSS^{+}16, YC13, YDR^{+}09, YLS12, YWW96, YLZ11, YLQW13, ZD07, ZYM96, ZEZF11, ZLL^{+}15, ZLZL16, Zhe03, ZJ02, ZL15, ZCG^{+}15, vdA01]. bases [BT92]. basic [Kha06a, KLK01]. basis [RDRLG^{+}15]. battlefield [DMP98]. BDCloud [LK16]. be [Kra97]. beaconless [WH11]. Behavior [ZD07, HHD98]. behavioral [CPCB11, PCB10, YSAMB96]. behaviors [XC07]. behaviour [Kri01]. behavioural [ND07]. behaviours [CCL11]. being [LBLB13]. benchmark [Duj91, SY04]. best [AGP16]. Between [Sch88, AST99, CT93, DYB91, Hog04, SVL02, SL02, You01, YLQW13]. bi [GVBVME13, GVBVME14]. bi-dimensional [GVBVME13, GVBVME14]. bidding [XC07]. big [KNP16, LJW^{+}15, XC15]. bilinear [VK08]. binary [CC98a, CC95b, IMA14, MFU02, SG90, TJY^{+}11]. bio [VKAR15]. biomedical [PMO16]. biometric [MCSV14]. bipanpath [FH13]. bipartite [Yan02]. bisectioning [LSM96]. bit [CZY^{+}03, TR06]. bitmap [SP16]. bitmap-based [SP16]. bits [SVL02, SL02]. blasting [GFZ^{+}16]. blind [JLL01]. Block [BI86, CC97]. Block-switch [CC97]. blocking [LT97, Ray00]. blocks [Rom98]. Bluetooth [RSDD10]. board [VG00]. Book [Ano01a, Ano02]. bounded [BD96, KB02]. Bounding [Sto90]. bounds [KI09]. BPEL [KVG08]. BPMN [RDRLG^{+}15]. Branch [HCK90]. break [KNP16]. Bridging [AST99, CC15, CC16, MS91]. broadband [KA97]. broadcast [MCMM95, SP16, WF93]. Broadcasting [CW91, CWY98, DD97, GGDG01, BPRS98, CC95a, HLLC96, PS95, WY93, LC96a]. broker [QAY09]. BTRIMER [TJY^{+}11]. bubblesort [Hsu99]. Buffer [ALP99, And98, CY90, FSA97, KA97, LH97, LWD04]. buffering [TR06, YWD00]. bug [FM10]. Building [War07, WZWD15, JPMAB00, KNEDK92, NTI^{+}07, XLW11]. Bulk [KW02]. Bulk-loading [KW02]. bulldozer [WSZG15]. burst [WD90]. burstiness [CH97]. bursty [GMF96, SE07]. Bus [WK85, GS01, KH96, LMI90, OSZ95, SVL02, SL02, Zhe03]. bus-type [Zhe03]. business [GCJP03, HDCH10, RFMP07, YHLC13]. business-to-business [YHLC13]. Byzantine [DGGS88, DGGS89].

C [PXQ08, PXY93]. C-means [PXQ08]. C2C [HTJ08]. cache [CZ09, IMA14, LNS^{+}07, RW00, SC10, SC11, WD90, WCD99, WWD04]. caching [FCGC13, LPY03, LYW93, LS04, LNS^{+}07, WY02]. calculus [MMO00]. calibrated [MRSW98]. call [CH98]. Cambridge [Bow86]. Campfire [CKJ10]. campus [WYLW16]. can [Kra97]. capability [MC95, QAY09, XM97]. capacitor [HA08, SKR08]. capacity [ELG00, QYCG10]. capital [KS04]. capture [vdA01]. car [CHZ11]. carbon [Zha16]. card [Hsu08, VK08]. Carlo [SSP03]. carrier [Tam08]. cascading [TH97]. Case [CDFP90, EKA06, KKEG^{+}09, Kof05, PRB13]. cases [AGP16, Kri08]. catalogue [ABNY12]. categorization [HC13]. Categorizing [EQ11]. category [CMZ10]. causal [BPRS98]. CBR [KSL12, SB96b]. CCCI [HCD04b]. ccNUMA [KKEG^{+}09]. CD [HJS87, HYY88, LMI90]. CDMA [MG02, WC08a]. ceiling [SED^{+}99]. cell [BVCW96, CH98, GMF96, KS97, YS99]. cellular [Ang16, WYO12]. centered [You01]. centralized [CDY92, DYB91, RD92]. centric [Kat05, PN10, PN07, TLB06]. Certification [BCPS10, DBBA10, NPR10]. chain [CO91, Wan03]. chain-dependent [CO91]. chains [TN93]. Chang [HLLC96]. change [vdAJ00a, vdA01]. changes [DSS00, LCWY11]. channel [OSZZ97, SA03]. Characterisation [HPTC11]. characteristic [WYLW16]. Characteristics [SLPK02, WC08b]. characterization [MCD^{+}15, SFLP99]. Characterizing [QYCG10]. cheater [WLY^{+}14]. Cheating [WLY^{+}14]. checking [KVG08, MHPS96, Nan94, XC07]. checkpoint [DG01b]. checkpointing [MZ03]. checkpoints [BCHR01, THT98]. Chinese [WWDL09, WWH95]. chip [Pos98]. choreographies [RDRLG^{+}15, SW12]. Chronolog [LOZ01]. Circle [Lee85]. Circuits [BBKT87, ML86, JTN95, MLLS94, MNK14]. class [CC97, LY95, PM94, Rom02, SG90, SSC97, XS02]. class-level [XS02]. Classes [BG87, DDL^{+}90]. classical [RS14]. Classification [HCD04a, KJI13, UPSL^{+}13, ABP09, DS97, HK90, PRB13, QLL14, VKAR15]. classifier [ZHZ^{+}15]. Classifying [BT92]. Clearance [HGR12]. Clearance-based [HGR12]. Client [FH02a, FH02b, HTJ08]. Client-side [FH02a, FH02b]. client-to-client [HTJ08]. clinical [WWDL09]. clock [YET94]. Clos [LT97]. Closed [Kar87, RS88, CQN^{+}16]. Closed-network [RS88]. Closest [CCC05]. Cloud [BV13, FVD13, GMP13, LdOS13, SV13, BSR11, BDKD12, HWH^{+}16, JWH^{+}15, LCWC14, WLXD15]. clouds [FBZS12]. Cluster [LM09, Mor07, CKJ10, Men99, NKWK07, TR02a, Wan97, WLG^{+}11, ZLZL16]. Cluster-based [Mor07, NKWK07, Wan97]. clustered [LB96]. clustered-star [LB96]. Clustering [Duj91, WYLW16, Cho11, PXQ08, YSS^{+}16]. clusters [KKEG^{+}09, TTD10]. CMH [GDK88]. co [PVRM11]. co-simulation [PVRM11]. Code [KKP11, FMS^{+}11, MG02, Nan94, PA10, WYO12]. coded [HA08, KSH96]. coding [CZY^{+}03, SL09]. cognitive [CP15, GQW^{+}14, ZZZZ15]. coherence [MD03, MD04, SC10, SC11, WD90]. collaborative [BBO08, ÇG11, GZ06, ZCK11]. collection [TR02b]. Collective [UPSL^{+}13]. colored [RES^{+}16b]. COMA [ZS06]. combination [FM14, YUM10]. Combinational [ML86, MLLS94]. command [Amm90]. Comment [LC96a]. commerce [LASS00, QZK09, SN07, YHLC13, YC13]. commercial [PC13]. commitment [Ray00]. common [Hui00]. Communicating [MYU10]. Communication [GLK89, SR03, BqQHj^{+}15, CWY98, DKP95, FAL^{+}01, GP90, Hsu94, HCD04a, HCD04b, KK96, MZ03, YA97]. communication-induced [MZ03]. communications [Ang16, SDC99]. community [MGOB15, RPD^{+}13, RPDH15, War07]. compact [DR09]. compaction [MS91]. Comparative [GS01, RD92, AMuRKK14]. Comparing [MHHP92, PC13]. Comparison [WF92, AD04, Duj91, IDY88, KNP16, KK96, RM98, WD90, WBA94]. comparison-based [WBA94]. compatible [QY15, SL97, WK85]. compensation [KVG08]. competence [SSH99]. complete [BGNP01, CC98a, CC95b, GGDG01, JC02, Kap92, Kap92]. completion [HWL97]. Complex [CF86, Meg90, BNSM01, MMO00, PN09, TYW09, ZD07, dA99]. complexity [CWS^{+}00]. Component [SMSJ10]. components [Che07, Wei95]. composition [CFLZ06, LLHG06, RDRLG^{+}15]. compositions [AdPT06, LV08]. compression [RMB91]. compromise [GS10]. compromise-resilient [GS10]. Computation [HP87, Buc00, HGR12, HN94, HC90, LCWC12, LC00b, RHS^{+}14, Sto90]. computations [CC98b, QA93, Sta93]. computed [YLK11]. Computer [AK92, Kar86, Kar87, Lee85, MMI85, PBS86, YDF87, CW91, CD93, CL16, DJ96, GKK03, HS88, Hof93, HLLC96, LC00a, zLTC95, LC96a, SF02, SM90, TN93, Vla07]. computers [DG01b, JSS^{+}98, LP93, PS95, SL97, Yan93]. Computing [BG87, FVD13, GMP13, HKL99, HRG87, LM09, LSG09, SKV12, BSR11, BDKD12, CMMA03, CMMA04, Fin90, FMP94, IS97, KT95, KKEG^{+}09, KI09, KSL12, LH98, NTI^{+}07, RS14, Sha96, Toh94, WLXD15, WCL15, Ye15, ZL15, ZL16, Zom98]. Concentrators [LL96]. Concept [Sha88, WZWD15]. conceptual [FCD03, FCD04, PN09, YW98b]. Concurrency [Kim04, SA86, DGGE88, RD92, SED^{+}99, WL91]. Concurrent [GP90, KH96, KA93, TDGNH97, WDF87, DJ02, HCK90, Pap98, Rom94, WYT01, XC07, BST98]. Concurrently [GET10]. conference [WH97, YDR^{+}09]. Configurable [SW12, IDY88]. Configuration [Sha88, KBG^{+}94, SVL02, SL02]. configurations [HA08, MHHP92]. conflict [MCMM95]. conflicts [PV96]. congestion [ACA99, ZYM96]. connected [PS95, WF93]. connection [BBMC98, WF92, KK96]. connections [FBZS12]. connectivity [FH13]. connector [MGOB15]. conquer [LSM96, LC00b, LPTH99]. consecutive [KS97]. consensus [HCL^{+}06]. Considerations [DKO86]. considering [YOM^{+}12]. consistency [KDP05, Ray02, TRAR02]. consistent [BCHR01, NMMS02, ZMMS04]. constant [LJL^{+}12]. constrained [JWZ14, Lee04]. constraint [CL07, CHL^{+}12, GR94, Pap98]. constraint-based [GR94]. constraints [Hsu94, HMPF91]. CONSTRICTOR [GR94]. construct [Amm90, BGRV99, CL92]. Constructing [Wei95, CSXC11, HABJ05]. construction [MKK15, Vla07]. Constructive [YUM10]. constructs [KT95]. content [DR09, YDR^{+}09, YLZ11]. content-based [DR09, YLZ11]. contention [MN14]. context [AR92, FL16, MCD^{+}15, MOEMK^{+}15, TYW09]. context-free [AR92]. continual [LPTH99]. continuous [ZXTS10]. contourlet [ZHZ^{+}15]. contract [BBMC98]. contracts [OO06]. Control [CKT94, ET10, LdOS13, SA86, Son87, AD04, ACA99, BBMC98, BNSM01, Cra06, Hal09, HJXY12, Hsu08, Kim04, KFW04, LWHS06, LJL^{+}12, LWD04, MN14, PONA11, QY15, RW11, RD92, SSS01, SED^{+}99, TEQ11, TZL96, WL91, WWH95, WC01, WYT01, ZYM96, ZEZF11]. Controlled [WY02, BVCW96, CF06]. Controller [LWD04]. convergence [WW01]. conversation [KB02]. conversions [RS96]. convertibility [LW10]. convex [OSZ95]. convolutional [ZKWM14]. cooperative [KB02, ON12]. coordinated [LS04, RZ01]. Coordination [RXR99, RMR00]. Coprocessor [HC90, XFA99]. CORBA [CMRR02, GMN98, KLMS02, NMMS02, OSN02, PM99, PRSM01, Zhe03]. CORBA-based [PRSM01]. core [IMA14, WGK^{+}11, Yan02]. cores [Yan00]. correcting [TZC09]. correlated [DD98, Gut01]. correlation [Cho11]. Cost [BDX93, CC97, Lee04, OSZZ97, SR03]. Cost-effective [BDX93, CC97]. cost-optimal [OSZZ97]. costs [QAS91, QA93]. coteries [SR03]. coupled [SG98]. Coupling [KHA06b]. Coverage [Kra97]. crack [GFZ^{+}16]. crash [AJTT15]. crash-recovery [AJTT15]. crashes [LSD95]. criteria [BBMC98, KNEDK92]. critical [GMC^{+}13, Toh94]. criticality [SMSJ10]. Cross [GAHL00, LCX08, NPR10, NXG^{+}16]. cross-authentication [LCX08]. Cross-organizational [GAHL00]. cross-platform [NPR10]. crossbar [Wan97]. CrossFlow [GAHL00]. Cryptanalysis [Hsu08]. Cryptographic [WC01]. Cryptosystem [WY93, CW91, CS97, Hsu08, HLLC96, LC96a, TJ01, WLC07]. CSMA [HJS87, HYY88, LMI90]. CSMA-CD [LMI90]. CSMA/CD [HJS87, HYY88]. CSP [KSL12]. cube [ASDOK03, WF93]. cube-connected [WF93]. current [SWJF11]. curve [Hsu08]. customized [JTN95, The94]. cut [LSM96, RSP91]. Cyber [XSZ16]. Cyber-Physical [XSZ16]. cycle [BGNP01, HL16, LWD04, MYYW06, SF11]. cycles [AL03, Cha02, WF93].

D [CSXC11, TC05]. daemon [KKK^{+}02]. DAP [BDG^{+}94]. Data [ABNY12, BBKT87, CKT94, DM06, Gai87, Ham08, HJS87, MPC91, MH10, MS91, TX16, ADSP12, AM00, BBO08, BNP05, BGRV99, CF06, Cer04, CC14, Cha08, CH11, CI94, CPT08, DZ15, HOGS11, IT07, Jon03, JIB03, KNP16, Kat05, KKHN13, LJW^{+}15, LS94, LSBW14, zLTC95, MKK15, ML09a, ML09b, RMB91, Sha96, SL14, Spi99, SB96a, SDN14, TD92, THDC08, VSZMCV08, WWS16, WCJ09, XC15]. data-driven [zLTC95]. data-parallel [Sha96]. Database [Elm87, HM87, LDF86, CHL^{+}12, CQN^{+}16, CL92, CDY92, CMRR02, Gab06, Hin99, HMPF91, JOR^{+}00, KHL00, LY92, LQX^{+}14, MHHP92, MYYW06, PV96, RSG96, RD92, SHGA13, SY90, SK92, TRT04, TX16, TH94, WL91, WZWD15, XFA99]. Databases [SA86, CCC05, DKP95, GE97, KS08, LWHS06, PDK95, PKC07, PC13, San90, SS99, WST05]. Dataflow [HP87]. date [LY96]. DB [OO12]. DBMISD [CSXC11]. DBMISD-D [CSXC11]. DBMS [Kim03, Kim04]. DDoS [TV08]. Deadline [KLK01, KHL^{+}08]. Deadlock [DGHE88, Elm87, GDK88, JDG92, Rom00, Sri91]. Deadlock-Avoidance [DGHE88]. deadlock-detection [GDK88]. Dealing [vdAJ00a]. DeBruijn [KR95]. debugging [TM94]. Decentralized [MS96, MMI85, Joe00, YA97]. Decision [MYU10, SDN14, GCM14, HDCH10, Vla07]. decomposition [CHZ11, CWS^{+}00, ZD07]. Decoupling [Tan97]. Deduplicated [PRPS13]. deep [ZKWM14]. defect [RAF15]. Defense [ZXG08]. defined [FMS^{+}11]. defining [AD04]. definition [CH11, ZZWD15]. deflection [BGM98]. Degradable [DGGS88]. Delay [SB96b, Kra97, Lee04, SE07]. delays [GDK88, TN93]. delegation [AHBU07]. delivery [EQ11, YDR^{+}09]. demand [WCLC04]. denial [SF02]. Departure [LY96]. Dependability [BCL^{+}01, BNSM01, HHD98]. Dependable [CKT94, TK00, SCKN11]. dependence [AL03]. dependencies [HMPF91]. dependency [BCHR01]. dependency-based [BCHR01]. Dependent [BI86, Mae10, Apo00, CO91, Hel13, WST05]. Deploying [LV08]. derivation [KI09]. deriving [QA93]. describing [MLLS94]. Description [CF86, CMMA03, CMMA04, GD06, ZD07]. descriptions [YSAMB96]. descriptors [MCSV14]. Design [Ano87, BqQHj^{+}15, BSK85, Bow86, CH97, CCdF^{+}07, DKO86, GA14, Gun93, IYD87, KTK14, LTKK09, zLTC95, LYL95, MRDF91, Omo91, SL97, SVN96, SK86, VB87, YH97, ZMMS04, Abi00, AdPT06, BCL^{+}01, CD93, Che93, CCDL09, CFLZ06, CPV90, CPS85, EKA06, HC08, Ito94, KNU12, Lov96, MLLS94, MNK14, PN09, RMDF91, Rom98, SHGA13, SY04, SL09, SSH99, SS13, The94, WH97, XS02, YS99, ZZWD15]. Designing [Bul11, CCSS06, MHH93]. designs [HF94, MNK14, MN98]. desired [KL00]. Desktop [LBLB13, KHL^{+}08]. Desktop-as-a-Service [LBLB13]. Detecting [Lov96, KNP16, WOE06]. Detection [Elm87, FVD13, AGP16, AP16, AR04, CCL11, CP15, GKK03, GDK88, HFS^{+}08, HL16, JDG92, KA93, LJW^{+}15, LDZP08, TDGNH97, WLY^{+}14]. detector [AJTT15]. deterioration [YUM10]. determination [AD13, AD14, BCHR01]. determining [KL00]. deterministic [HMF93]. develop [RFMP07]. developer [TJS15]. Development [KNU12, Kim03, XS02, BGFL08, BGNP01, KiMKT94, MJRIV14, NCB06, PS10, Tam08, TYY^{+}12]. devices [CCdF^{+}07, DD11, MCSV10, ML09a, ML09b]. diagnosability [SS93]. Diagnosis [Car14, PLL00, SS93, TC06, WBA94, WF92]. diagrams [GCJP03]. diamonds [AD13, AD14]. Different [IB10, DDL^{+}90]. differential [ADMB15]. differentially [XLH16]. Differentiated [jKsJdChK12, SLPK02]. Diffracted [IB10]. Diffusion [GMF96]. Digital [Wu01, YLS12, AD13, AD14, CL00, CCdF^{+}07, FH12, JTN95]. dimensional [ACR00, GVBVME13, GVBVME14, HK90, JKC16, KR92, OSZ95, Rav92, Rav95, WBY^{+}13]. dimensionality [JKC16]. Dimensioning [BBMC98]. dimensions [DGL00]. Direct [BCHR01, KA97]. direct-access [KA97]. direction [CP90, SWJF11]. Disc [HS85, Bed88]. discipline [LY96, Yan02]. disciplines [You01]. Discovering [CQN^{+}16, YLQW13, FBZS12]. Discovery [GPR10, ASH12, BWG06, HCL^{+}06, SL14, WWDL09]. Discrete [Sha09, ELG00, HJXY12, JLL01]. Discussions [LSBW14]. disjoint [Nan94]. Disk [TX16, CHKW99, TF06, Tho13, YWD00]. display [TJS15]. dissemination [WC08b]. dissimilarity [AP16]. distance [BD96, CPCB11, MM15, PCB10]. Distributed [Cho11, DGHE88, DKO86, Elm87, FAL^{+}01, HKL99, JC97, Men99, SZ91, Sha88, SL88, Son87, Apo92, BCLG98, BGNP01, BYX16, BM97, Bun98, CSXC11, Cha02, CG00, CDY92, DKP95, DYB91, DG01b, EG97, FC99, Fin90, FMP94, GW01, GST03, HPF02, HSS^{+}11, HS96, Hof93, IS97, JDG92, JM99, Jen01, JSM96, JTL01, KT95, KS00, Kar91, KLK01, KLMS02, KB02, KI09, KC10, KSK99, LC00a, LH98, LSD95, LYL95, MP88, MW90, OSN02, OESHK07, OESHK08, OO12, PV96, Ray00, Red88, Red91, RXR99, RZ01, San90, SL97, SSC91, SDC99, SM90, Sri91, SS99, SC10, SC11, THT98, TRAR02, WL91, WSL93, WLC93, Wan03, WY96, WW01, WF93, XS02, Yan02, YW98a, YW98b, YZP97, YH97, ZMMS04, Zom98, CLO92]. distributed-centralized [CDY92]. distributed-object [XS02]. Distribution [DG87, Bed88, HA08, KS97, LC00a, QY15, SKR08, TM94, WH97]. distributions [KC94, LQLL11]. Diversity [CKT94, BGFL08, RSBW14, Lov96, Rom02]. Divide [LSM96, LC00b]. Divide-and-conquer [LSM96, LC00b]. Division [HJS87]. DNA [KNP16, RHS^{+}14]. document [YWW96]. documents [Cra06, GKS06, HHCASEQ11, HC13, MdlFD03, YLZ11]. domain [EKA06, Kof05, LTLH08]. domains [MOEMK^{+}15]. domestic [Kat14]. domination [Yen08]. domino [Rom00]. DOT [MC95, WBY^{+}13]. DOT-based [WBY^{+}13]. dotted [WYO12]. double [LY12, LQLL11, MC95, MNK14]. double-layered [LY12]. double-sampled [LQLL11]. down [CPV90]. DQDB [MOZR03]. Drago [MÁGA00]. DRB [KBG^{+}94]. Driven [SV13, SE85, AKT12, BGFL08, DZ15, HOGS11, LJL^{+}12, LTLH08, zLTC95, RDRLG^{+}15, SYNB99, SE07, SLA^{+}11, YS99, ZCK11]. Drivers [Ano87]. drives [LYL12a]. driving [ZHZ^{+}15]. DS [MG02]. DS/CDMA [MG02]. DSL [GH11]. DSP [ALP99]. DSSH [YLS12]. DTN [WYLW16]. Dual [LCWC14, SVL02, SL02, Wu16]. duals [RS88]. due [LY96]. duplex [IDY88]. duty [HL16]. Dynamic [CDY92, JWH^{+}15, LJ97, MLK15, RS96, TEQ11, WWH95, AGP16, Fin90, GAHL00, HOGS11, HSS^{+}11, Joh96, Kha06a, KW01, Lee09, LWL93, LWD04, PV96, Pos98, SZ03, SCL97, TJY^{+}11, WWD04, YWD00, Zhu96, vdA01]. dynamically [Ric90, TCD07]. dysarthria [Car14].

e-business [HDCH10]. e-commerce [QZK09, SN07, YHLC13, YC13]. e-governance [ZL15]. e-government [SL12]. e-GrOV [VSZMCV08]. e-health [UI06]. e-science [SWJF11, ZCK11]. e-transactions [HCHD06]. E-TTM [PONA11]. earliest [LY96]. earliest-due-date [LY96]. Early [PVRM11, TC05, AP16, BCL^{+}01, KT95]. early-receives [KT95]. ease [MGOB15]. Easily [KR92]. easy [YW98b]. ecologic [ZL16]. economy [ZL16]. edge [MNK14]. edges [IB10]. Editor [Pot14]. Editorial [Ano90, Ano92, CDF03, FP13, Ham05, Ham08, MMD99, PB02, PS01, RXR99, YBT12, YS12, ZG05, vdAJ00b]. education [BZLX16]. EFCI [TZL96]. Effect [KC94, YDF87, BXST12, NXG^{+}16, Rom00]. Effective [BSS90, JKC16, WL01, BDX93, CC97, TMNL08, Wan03]. Effectiveness [SY90, HHCASEQ11, Lam09]. effects [GFZ^{+}16, SN07, WCD99]. Efficiency [DGGS89, NK03]. Efficient [BPRS98, BGM98, Buc00, CY94, DG01a, GFH^{+}10, HN94, HP87, HWL97, Li98, LQX^{+}14, MK90, ML09a, ML09b, QWK14, TH97, Yan00, CZ09, Cha08, CCHL03, DG01b, GKS06, HC90, KKHN13, LNS^{+}07, zLTC95, LC00b, LYC02, SSP03, TYY^{+}12, WH11, Wu16, XLSM10, Yan02, YWD00, YA97, ZZZZ15]. efficiently [WY96]. Eigenproblem [Meg90]. Eight [TMNL08]. electric [WSZG15]. Electromagnetic [IB10]. electronic [AAKD09, LASS00]. electrostatic [YLQW13]. eliminate [LWD04]. elliptic [Hsu08]. Email [KJI13, EQ11]. Embedability [Hsu99]. embedded [CC95a, GMC^{+}13, HSS^{+}11, KI09, MCSV10, SN09, SCKN11]. Embedding [AA99, Koc94, CC98a, CC95b]. Emergency [IST^{+}13, SP16, ZL15]. emerging [Sel02]. emotion [LKK14]. Empirical [FM10, KJI13, CHZ11]. emptive [CCZ03, LPY03]. emulation [MRSW98]. en-route [LS04]. enabling [KLMS02]. enactment [Joe00]. encrypted [FJ04]. encryption [Ang16, HCD08a, HCD08b, HCC98]. encryption/multisignature [HCC98]. Energy [WH11, Wu16, YOM^{+}12, ZZZZ15, JWH^{+}15, KKHN13, XLSM10]. Energy-efficient [WH11, ZZZZ15, KKHN13]. Energy-saving [YOM^{+}12]. enforce [HMPF91]. Engine [IYD87]. Engineering [CFLZ06, Dow87, GL06, GMPR08, AST99, AK92, CSB16, Sha09, SA03, TLB06, You01]. enhance [SIGC15]. enhanced [AGP16]. enhancement [CHZ11]. enhancements [FH02a, FH02b]. Enhancing [CPCB11, FCGC13, PCB10, CU07, WW01]. ensemble [CHZ11]. enterprise [BSMB14, FH12, HSS^{+}11]. enterprises [GAHL00]. entity [SHGA13]. environment [Abi00, Apo00, BGNP01, BDKD12, DKP95, JWH^{+}15, KSL12, MPC91, Men99, PVRM11, SPB11, TYW09, WLC07, YQ95]. Environments [SV13, BBO08, HMPF91, JC97, LTLH08, LYL^{+}12b, MCD^{+}15, NTI^{+}07, RCNL15, RMR00, SP16, Zom98]. EPON [jKsJdChK12]. equal [SL02]. Equivalence [CT93]. equivalent [AHH^{+}16, Buc00, HWL97]. Error [CKT94, LDF86, BDMV97, KA93, MMH16, TDGNH97]. error-aware [MMH16]. errors [SMSJ10]. ESGIA [SPBK12]. establishment [HL99]. Estelle [LJ96]. estimate [GMF96]. Estimates [BG87]. estimating [LQLL11]. Estimation [MMH16, SBK^{+}92]. Eulerian [SF11]. EV [HTJ08]. EV-C2C-PAKE [HTJ08]. Evaluating [DGGS89, HHCASEQ11, KS04, CY94, MRSW98, vRB11]. Evaluation [KJI13, PRPS13, SF11, SK86, BVCW96, CHCL90, Hof93, Ito94, KNU12, KNP16, KTV14, KW02, Lam09, LH97, MN14, MRDF91, RKZ99, SM90, TK00, WMFK95, YSS^{+}16, YW98a, ZL16]. evaluations [NHK14]. event [Bun98, DD11, LJL^{+}12, OSN02, VT95]. event-driven [LJL^{+}12]. events [YOM^{+}12]. evolution [ADMB15, Hin99]. evolving [BT92]. excellence [CKJ10]. Exception [dLR01, Nan94, RXR99]. exceptions [MLGC94, TD92]. exchange [HTJ08, KR95, LHC08, SB96a]. exclusion [CG00, Hof93, JM99, MS96, SM90]. Executable [UPSL^{+}13, PONA11]. Execution [SV13, EG97, GE97, LOZ01, SN09, Sto90, WLC93, WGK^{+}11, Wu16, YQ95]. Exhaustive [ML86]. existing [MKK15]. exits [LY12]. Expandable [DHH96]. expansion [AKT12, HC13]. experience [CPS85]. Experimental [RKZ99, YWS06, Lov96]. experiments [KTK14, OO12]. Expert [EGH^{+}86, KNEDK92]. Expert-Systems [EGH^{+}86]. Exploiting [AHH^{+}16, KHR^{+}09, WY96]. Exploring [CL16]. exponential [BT98]. exponentiation [LC00b]. expression [Met16]. expressions [KHL00]. Extended [CF86, HMF93, SSS01]. Extending [PMO16, YMST98, RW11, SD99]. Extensible [SPBK12, YMST98]. Extension [GCJP03, KRSS16, RFMP07, SD99]. extracted [QWK14]. Extracting [Cer04]. extraction [CGR^{+}09, CMZ10, MdlFD03, Met16, PMO16, YLZ11].

Fabric [RAF15]. fabrics [BGM98]. face [FM14]. Facial [CGR^{+}09, LKK14]. facilitate [AKT12, LWL07]. facilities [PDK95]. facility [YH97]. factorization [JWZ14, WLC07]. factors [ACB09]. failure [AJTT15]. failures [Gut01, TK00]. Fair [JLL01]. Fairness [MOZR03, TZL96]. family [MGOB15]. farming [SL97]. Fast [CCHL03, CC98b, OSZ95, SB96a, YLK11, Zhu96, LC98, MCSV14, TDGNH97]. Fault [ACR00, BDMV97, CH03, CH04, CC95b, Elm87, GMPR08, IYD87, MG95, NCMH00, PL88, RMDF91, WF93, YDF87, YS90, ASDOK03, Apo00, AGT98, BSS90, BDX93, CD94, CC98a, CMRR02, DG01a, FMP94, Gun93, HS88, HS96, JM99, KKY04, KBG^{+}94, KiMKT94, LV08, LY95, LYL95, MRDF91, NMMS02, PLL00, PHPd98, RC98, RW00, Toh94, TH94, WBA94, WF92, YET94, YA97, ZMMS04]. fault-injection [AGT98]. Fault-tolerance [MG95, NCMH00, TH94]. Fault-Tolerant [Elm87, YDF87, ACR00, BDMV97, CH03, CH04, CC95b, RMDF91, WF93, YS90, ASDOK03, BSS90, CC98a, Gun93, HS88, HS96, JM99, KKY04, KBG^{+}94, LY95, MRDF91, NMMS02, Toh94, YET94, ZMMS04]. Faults [YDF87, Kra97, Lov96, MS91, Rom98, VG00]. faulty [AA99]. FBASE [Mul92]. FCFS [CO91]. FCST [LK16]. FDTD [Mae10]. feasible [Kat14]. Feature [KJI13, CGR^{+}09, CCL11, DR09, LYL12a, LKK14, VKAR15, WOE06]. feature-based [CCL11]. features [BKBM10, BKBM13, QWK14, YW98a]. fed [DD98]. federated [AKT12, Mul92]. Fermat [TDGNH97]. field [LCWC12]. Fields [IB10, QWK14]. FIFO [ACA99]. File [Bed88, JSM96, KW01, KW02, LCWY11]. files [JSM96]. filtering [BCLG98, MLK15, QWK14]. find [Hui00]. finding [Cha02, Din04, KS00, Yan00]. fine [ADMB15]. fingerprint [MCSV14]. finite [AM00, OESHK07, OESHK08, WCD99]. first [NK03]. fixing [FM10]. flash [LYL12a]. Flexibility [DGL00, Sch88]. flexible [BDX93, GCJP03, Joe00, ZD07]. Flight [CKT94]. flip [MNK14]. flip-flops [MNK14]. flops [MNK14]. FLOSS [PS10]. Flow [ET10, AHH^{+}16, LS94, LQLL11, TZL96, ZLL^{+}15]. flow.net [Joe00]. flows [Cer04]. folding [YCFX09]. footprint [Zha16]. force [YLQW13]. fork [Tho97, CT93]. Fork-Join [CT93]. fork/join [Tho97]. Formal [KVG08, MN98, NC03, NC04, PHPd98, SEuH96, ZCL03, BCPS10, Che07, FL16, KRSS16]. formalization [Gab06]. formation [Hog04, ON12]. Formulae [MB88]. forward [HCD08a, HCD08b]. forwarding [CS15]. Four [MLLS94]. Four-valued [MLLS94]. Fourier [DZG92]. FPGA [Ang16, IST^{+}13]. FPS [BDPV90]. FPS/264 [BDPV90]. fractal [MVA09]. Framework [FVD13, TY97, AHBU07, ASH12, AKT12, BLY10, BM97, CZY^{+}03, DZ15, FAL^{+}01, JPMAB00, KTJ14, KKY04, KJKK07, KSL12, KDP05, LTLH08, OO06, QY15, RPD^{+}13, RPDH15, SSB^{+}07, SC09, SLA^{+}11, TCD07, Tan97, TLB06, WKT07, WY96, YW98b]. free [AR92, TR06]. Frequency [Mae10]. frequent [BYX16, CQN^{+}16]. frontal [CGR^{+}09]. full [HL16]. Function [HYY88, KA93]. Functional [HMPF91, JTN95, WOE06]. fundamental [Cha02]. fusion [MCSV14, WWS16]. future [SWJF11]. fuzzy [CL07, FM14, GA14, KTJ14, KNEDK92, KTV14, TEQ11]. fuzzy-labview-based [GA14].

gain [SVL02, SL02]. game [LYL^{+}12b]. games [MOEMK^{+}15, ON12, Ye15]. gap [AST99, CC15, CC16]. gateway [SZM^{+}01]. gathering [KKHN13]. gaze [CCL11]. General [RM02, AA99, BST98, KKK^{+}02, SSP03, SK92]. generalization [MK11]. Generalized [TJ01, Kri08]. Generating [CJV01, ZLL^{+}15, SSP03]. generation [BGRV99, JTN95, MS96, PRB13]. generator [Wal90]. Generic [BPX06, CC14, AR04, Ray00, SY04, WLXD15, vdA01]. Genetic [YUM10, Agu03, BXST12, CL07, HA08, JIB03, PN09, SYNB99, SKR08]. genetiv [AdPT06]. geographic [MKK15]. geographically [DYB91]. geometric [HJXY12, TM94]. Geometry [YUM10]. German [CKJ10]. GIS [CL92, KHA06b]. Give [HCY87]. Give-Up-Based [HCY87]. giving [Wu01]. Global [DJ96, BCHR01, IST^{+}13, KJKK07, YS99]. Goal [RCNL15, RJS06]. goal-oriented [RJS06]. Goals [CSB16]. Google [CU07]. governance [ZL15]. government [SL12]. GPS [IST^{+}13]. Gracefully [DGGS88]. grade [Tam08]. gradient [ERY^{+}16, RES^{+}16a, RES^{+}16b]. gram [KWL07]. Gram/2L [KWL07]. grammars [GD06]. grammatical [Amm90]. granularity [GP90]. graph [Cha02, GD06, Hsu99, JWZ14, LB96, Pap99, TZC09]. graph-partitioning [TZC09]. graphical [BKBM10, BKBM13, NHK14]. graphics [DGGE88]. Graphs [HP87, Bun98, CLO92, KR95, LSM96, SS93, Yen08]. grey [YSS^{+}16]. Grid [Ham08, LM09, TYY^{+}12, ABNY12, ASH12, CWBZ11, CHE12, KC10, RPD^{+}13, RPDH15, SPB11, SPBK12, VSZMCV08, ZL15]. grids [KHL^{+}08, RSDD10, SPB11]. Group [LW10, TR02a, TJ01]. group-by [TR02a]. group-oriented [TJ01]. grouping [YWD00]. groups [HCC98, RPD^{+}13, RPDH15]. GrOV [VSZMCV08]. GTOOLS [LSR^{+}92]. guaranteed [CH97, KS99]. Guest [Pot14]. GUI [LSR^{+}92, NPR10].

Hadoop [ZLZL16]. Hammerstein [ERY^{+}16, RES^{+}16b]. hand [LH02]. hand-off [LH02]. handheld [ML09a, ML09b]. handle [vdA01]. handling [KLK01, MMO00, Nan94, RXR99, TL95, dLR01]. handover [YLS12]. hard [VM98, WGK^{+}11]. hardware [Abi00, HF94, KHR^{+}09, Lov96, RSP91, TM94, XZ11]. hardware/software [Abi00]. Harmony [VKAR15]. hash [WY02]. hashing [KK96]. HBP [LWD04]. HDA [TX16]. header [QWK14]. health [UI06]. healthcare [MOEMK^{+}15]. Heavy [ZCG^{+}15, Kat92]. Heights [IB10]. heterogeneous [IS97, LYL^{+}12b, LP03, RM98, Rom94, TX16]. Heuristic [Agu03, MS97, Sri91, Lee04, SYY^{+}10]. heuristics [GH11]. HHT [GFZ^{+}16]. hierarchic [YSS^{+}16]. Hierarchical [BNSM01, OO12, RES^{+}16a, RES^{+}16b, TRLD10, CSQW14, Cra06, GS10, QLL14, SDC99, SP16, WC01]. hierarchically [LSD95]. hierarchies [Cer04, dA99]. hierarchy [LWL93, Nan94]. High [Ano87, JTL01, SV13, YSAMB96, CSXC11, CCdF^{+}07, HK90, JKC16, KLMS02, Kim03, KS04, Omo91, Sah95, SN09, SB96b, TTD10, ZZWD15]. High-Availability [Ano87]. high-definition [ZZWD15]. high-dimensional [HK90, JKC16]. High-level [YSAMB96, KLMS02]. high-performance [Kim03]. high-precision [KLMS02]. high-priority [SB96b]. high-speed [Sah95, TTD10]. highly [FM14, The94]. hints [TKCR04]. Hippocratic [LWHS06]. histogram [MKK15, XLH16]. hitter [ZCG^{+}15]. Hoare [MP88]. hoc [KTV14, LHC08, PYHO04, SSB^{+}07, WCLC04, WC08b, WWS16]. holders [Hel13]. homogeneity [RAF15]. homogeneous [Kar91]. homotopic [HGR12]. Horizontal [IB10]. hull [OSZ95]. Human [MYU10, CD93]. humanoid [TEQ11]. Hungarian [DD90]. Hybrid [DE00, KI09, CWBZ11, CDY92, DG01a, NC03, NC04, RSDD10, TV08, WSZG15]. hyper [TM94]. hyper-geometric [TM94]. Hypercube [HP87, Agu03, DG01a, FH13, MOK03]. hypercubes [AA99, BD96, CC98a, CC95b, CC98b, SS93]. hypercycles [DD97]. hypermedia [SA03, YWW96]. HyperMonitor [XZ11]. hypervisor [GMC^{+}13].

I/O [Ano87, SIGC15]. IBM [SL97]. ICA3PP [LL15]. ICA3PP/U [LL15]. ICA3PP/U-Science [LL15]. ICTree [FBZS12]. ID [CS97, CL00]. ID-based [CS97, CL00]. Identification [UPSL^{+}13, ERY^{+}16, EG97, FMS^{+}11, LC00a, PN09, QL13, QL15, RES^{+}16a, RES^{+}16b, WLY^{+}14, ZKWM14, ZCG^{+}15, vdAJ00a]. identifiers [FMS^{+}11]. identifying [YW98a]. Identity [PN10, PN07]. IEEE [MN14]. II [MRDF91]. Image [WYO12, AD13, AD14, CC15, CC16, CMM91, DR09, DGGE88, FXZ16, HN94, NHK14, QLL14, WKT07, YUM10, YZP97]. image-processing [DGGE88]. images [RAF15]. IMI [PN10, PN07]. Impact [ACA99, ZS06]. impaired [CP15]. Implementation [CHCL90, NHK14, ASH12, Ang16, CCdF^{+}07, CMM91, Gun93, JTN95, JSS^{+}98, KBG^{+}94, LTKK09, LB03, LCX08, LYL95, MK90, MRDF91, RC98, SL97, The94, VG00, YH97, ZMMS04]. implementations [BPX06, Men99]. implemented [BWB^{+}01, MD03, MD04]. Implementing [Lee85, MJRIV14, vRB11]. implicants [RS14]. implicates [RS14]. Implications [GZ06, MD03, MD04]. implicit [CP90]. improve [EQ11]. Improved [Hof93, Sah95, Tho13, HTJ08, LY12, PXQ08, SHGA13]. Improvement [KiMKT94, RP90]. improves [NXG^{+}16]. Improving [ADSP12, CS15, HL16, LHC08, QAY09, WP95, YET94]. In-car [CHZ11]. Incomplete [HM87, DG01a, GGDG01]. incorporated [OO06]. Incorporating [ABRW93, TD92, KT95]. Increasing [MK11]. Incremental [XLW11, RW00, RKZ99]. independent [Yen08, ZJ02]. Index [Ano97, Ano98, Ano01b, Ano03, Ano06, Ano07, Kim03, KWL07, SP16, TRT04, WBY^{+}13]. indexing [CZ09, JKC16, QL13, QL15]. Indian [KTJ14]. indifference [CLO92]. induced [MZ03]. industry [Kat14]. Information [CL16, ET10, GMP13, HM87, LQ14b, BSMB14, CC95a, KRSS16, LNR04, MdlFD03, MM15, NKWK07, ND07, PMO16, QZK09, Red88, SPBK12, THDC08, UI06, WP95, WC08b, War07, WWD04, vdA01]. informative [QWK14]. informed [FCHJ05]. Infrastructure [PN10, PN07, CWBZ11, KHA06b, SCKN11, TYW09, ZEZF11]. infrastructures [NTI^{+}07, SN07]. initial [Jen01]. initiated [WSL93]. injection [AGT98]. Inlining [BA12]. inner [PTX^{+}09]. inner-node [PTX^{+}09]. InOrder [CU07]. input [FG97]. inrush [BqQHj^{+}15]. inspired [VKAR15]. instance [XS02]. instance-level [XS02]. Instruction [Sch88, Omo91, YCFX09]. instrumentation [HSS^{+}11]. Integer [WLC07]. Integrated [CY90, HJS87, Joh96, SL88, SS13, BGNP01, DJ96, JTN95, YSS^{+}16]. Integrating [CH98, LP03, MKK15, RJS06, VSZMCV08]. Integration [KKHN12, KKHN13, ABNY12, MG02, MPC91, SL12, YMST98]. integrity [Hal09, HMPF91, JOR^{+}00]. intelligence [ELG00, WWDL09]. intelligent [MFU02, TEQ11]. intention [RCNL15]. inter [DSS00]. inter-organizational [DSS00]. Interaction [Kri08, GD06, MGOB15, RJS06, XCHY11, You01, ZLL^{+}15]. interactions [HDCH10, WOE06]. Interactive [GLK89, HRG87, SL14, Amm90, BCLG98, KNU12, MJRIV14, OSN02, SS13]. interceptors [FH02a, FH02b]. Interconnection [PL88, CC97, DG01a, DS97, GGDG01, Hsu99, HL99, JSS^{+}98, Kam93, PLL00, PM94, SVN96, YS90]. interconnections [TH97]. Interface [Bow86, TCD07, BBMC98, CD93, KHE01, YMST98]. interfaces [MJRIV14]. interlocking [BNSM01]. Internet [BqQHj^{+}15, CL16, HZS12, LWL07, WP95, WWD04, XC07, YC13]. interoperability [MFU02]. interoperable [PDK95]. interoperation [LTLH08, YMST98]. interplay [You01]. interpolating [CW91, LC96a]. interprocess [SDC99]. interprocessor [KK96]. interrupt [SE07]. interrupt-driven [SE07]. interval [LPY03]. intractability [Nem96]. Intrusion [FVD13, AR04, GKK03, LDZP08]. invalidate [WCD99]. invariants [AGP16]. inverse [HN94]. inverted [KWL07]. Investigation [WSZG15]. involving [DG01b, Pos98, TRT04]. IP [Apo00, CCDL09, DD11, KTK14, PC13, PA10]. IP-SEC [PA10]. iris [MCSV14, MCSV10]. ISL [ZEZF11]. isomorphic [Kap92]. isomorphism [PN09]. Issue [Ham08, LQ14b, MA97, SK98, ZXG08, Zom98, CFLZ06, DGM06, DBBA10, DM06, FHFL07, GL06, GMPR08, Hal09, HKL99, KG07, LWHS06, LM09, LSG09, LK16, LL10, MH10, SKV12, SWJF11, TS99, XSZ16]. Issues [Men99, KLK01, LYW93, Tan97, TMNL08, TZL96, vdAJ00a]. itemsets [CQN^{+}16]. iterative [HJXY12, TZC09].

J2EE [HBL16]. Jacobi [Meg90]. JADE [NCB06]. Japan [Toh94]. Java [BA12, FMS^{+}11, GMC^{+}13, Jen01, YCFX09]. jitter [TR06]. jitter-free [TR06]. jitters [KKLK02]. jobs [CSQW14, GHW15, YQ95]. join [KW01, LHW04, TL95, TR02b, Tho97, WLC93, WBY^{+}13, CT93]. joins [TY97].

KBMS [LSR^{+}92]. Kernel [GLK89, CL92, QLL14]. kernels [KS99, SE07]. key [HCD08a, HCD08b, HTJ08, LC00a, LHC08, SSS01, TMNL08, VK08, WH97, WC01, Wu01]. key/lock [SSS01]. keyword [YLZ11]. kit [TYY^{+}12]. knapsack [LWL93]. KNN [ZXTS10]. Knowledge [GPR10, BZLX16, BT92, SS13, XC15, ZLL^{+}15]. knowledge-based [SS13].

labelling [GKS06]. laborious [Spi99]. labview [GA14]. LAN [BVCW96, KBG^{+}94]. language [BKBM10, BKBM13, BST98, CMM91, GR94, GLT05, Kof05, Lam09, LB03, PXY93, KHL00]. languages [AD04, LYA06]. LANs [CCZ03, GS01, Sah95]. Large [CI94, ML86, CPT08, LQX^{+}14, LCWY11, OSN02, Red91, WZWD15, ZLL^{+}15]. large-scale [CPT08, OSN02, ZLL^{+}15]. layer [BqQHj^{+}15, CSXC11, CS15, SCL97]. layered [LY12, QLL14, SL09]. layers [KSK99]. Layout [BBKT87]. layouts [SEuH96]. learned [VG00]. learning [BSR11, BZLX16, CC15, CC16, GKK03, HJXY12, ZKWM14]. least [CH03, CH04]. leel [SZM^{+}01]. legacy [CMRR02]. legal [MdlFD03]. length [KL00, LDZP08, LQLL11]. less [Wei95]. Lessons [VG00]. Level [DG87, SV13, CH03, CH04, CG00, CMM91, GH11, IMA14, JIB03, KTJ14, KLMS02, KWL07, LWD04, LP03, SSH99, XS02, YSAMB96, ZT03]. levels [Hal09, TX16]. Levenshtein [CPCB11, PCB10]. Leveraging [SA03]. libraries [TTD10]. library [GH11]. lifecycle [dLR01]. lifecycles [RCNL15]. lifetime [QKSN09]. lifetime-adaptive [QKSN09]. lightweight [XZ11]. LimeVI [WLG^{+}11]. limited [GZ06]. LIMITS [MN98]. line [RAF15]. Linear [SL02, Hui00, SG90]. Linguistic [GPR10]. link [ELG00, LL94, MM15, SVN96, TK00, Wan03, XLW11]. linked [QL13, QL15]. Linux [PVRM11]. Lisp [FC99]. list [BKBM10, BKBM13, KHL^{+}08]. live [WLG^{+}11]. living [RCNL15]. Load [BM97, DMP98, FMP94, RSG96, Cha08, CDY92, Fin90, Kar91, LCH92, LWL07, LJ97, SCL97, Tan97, Vla07, WSL93]. load-balancing [Tan97]. load-sharing [Fin90, LCH92]. loading [ABNY12, KW02]. Local [BG87, CL07, DMP98, LMI90, LJW^{+}15, RAF15]. local- [DMP98]. localization [RAF15]. location [CHL^{+}12, Kat05, NKWK07, WST05]. location-based [CHL^{+}12, Kat05]. location-dependent [WST05]. Lock [IYD87, SSS01]. locks [CC95a]. log [QL13, QL15, Spi99]. logarithm [JLL01]. Logic [BSK85, KTJ14, MP88, MLLS94, NC03, NC04, Pos98, VT95]. Logical [KSK99, Gab06]. logics [HN94, RS14]. LogN [IMA14]. loneliness [AJTT15]. Long [HWH^{+}16]. Long-term [HWH^{+}16]. longest [Hui00]. LookAhead [ZXTS10]. looking [DE00]. lookup [CZ09]. loop [HWL97, Lee04]. loops [AL03, HWL97]. loss [GMF96]. losses [KS97]. Low [Lee09, CZY^{+}03, CMM91, Zha16]. low-carbon [Zha16]. low-level [CMM91]. LRD [MOK03]. LTE [LCWC14].

M QUAD [KBB09]. MAC [MG02]. Machine [HM87, DZG92, GHW15, HN94, HFS^{+}08, HMPF91, MHHP92, TH94, KK96]. Machines [SV13, FBZS12, JWH^{+}15, OESHK07, OESHK08, SG90]. magics [YC13]. Magio [SIGC15]. main [Kim03, Kim04]. main-memory [Kim04]. maintaining [CSXC11, LC00a]. maintenance [LM91, RKZ99]. making [HDCH10, GMN98]. malicious [LMPJ12]. malleable [CSQW14]. Management [DM06, Ham08, MH10, PN10, PN07, ACB09, And98, Apo00, AM00, BZLX16, BBO08, BNP05, CKJ10, CHL^{+}12, CY90, DGL00, DMP98, GVCV13, GAHL00, HZS12, Jon03, Koc94, KW10, LH97, MRDF91, TYW09, TYY^{+}12, YLS12, ZL15, vdA01]. manager [Kim03]. managing [ZJ02]. MANET [LY12]. manifold [CC15, CC16]. Manipulator [SL88]. Manufacturing [SL88, ZD07]. map [AKT12, XC15]. map-reduce [AKT12]. Mapping [SG90, Agu03, ABNY12, LTLH08, Nan94]. mapreduce [ZCG^{+}15]. maps [HPTC11]. markets [KS04, TBEH04]. Markov [TN93, Wan03, ZLZL16]. Markov-chain [Wan03]. mass [AD13, AD14]. Massive [Pos98]. massively [DZG92, Nem96, RSG96]. master [TJY^{+}11]. master-slave [TJY^{+}11]. Match [YWL^{+}03, KA93]. matcher [DZ15]. matching [ABNY12, AP16, DZ15, KS00, KWL07, KS08, MK90, MFU02, ML09a, ML09b, MK10, YZP97, ZZWD15]. Matchmaking [SC09]. materialized [RKZ99]. mathematical [LS94]. matrices [GZG95]. matrix [JWZ14, LS97, SBK^{+}92]. maturity [KTJ14]. maximization [HWH^{+}16, JWH^{+}15]. Maximum [SVL02, KS00]. MBR [MK10]. MBR-safe [MK10]. MCFS [KNEDK92]. MDA [RFMP07]. MDDT [KT95]. MDE [GH11]. MDE-based [GH11]. Means [LdOS13, MdlFD03, PXQ08]. measure [Met16, SF02]. measurement [CWS^{+}00, HCHD06]. Measures [SK86, RG88, YLZ11]. mechanical [GA14]. mechanism [CCZ03, HZS12, LWHS06, PN09]. Mechanisms [DGHE88, GMP13, MN14, RMR00, RW00]. mechatronic [HOGS11]. medicine [WWDL09]. Meeting [The94]. Mellor [PHPd98]. members [MGOB15]. memories [MHH93]. Memory [Red91, ALP99, DAFG95, DMP98, HC90, KHR^{+}09, Kim03, Kim04, LYL95, Red88, SG98, SC10, SC11]. MEMSY [DHH96]. MERASA [WGK^{+}11]. merge [TR02b]. mergesorts [WYT01]. merging [BDG^{+}94, PRB13]. mesh [BXST12, CCHL03, CL96, PS95, Zhu96]. mesh-connected [PS95]. meshed [HA08]. meshed/ [HA08]. meshes [ACR00, MS97].

Message [HYY88, PA10, CL96, DAFG95, Kat92, KJKK07, QWK14, TTD10, Wu01]. Message-Based [HYY88, Kat92]. message-passing [TTD10]. message-triggered [KJKK07]. Messaging [GST03]. messy [You01]. Metadata [LCWY11, FJ04, LP03]. Metamodel [ZCK11]. Metamodel-driven [ZCK11]. metamodels [SLA^{+}11]. Method [Mae10, WOE06, ABNY12, CH03, CH04, Che07, CP90, GFZ^{+}16, KNU12, KTK14, jKsJdChK12, LQLL11, LC98, MK11, Meg90, QKSN09, QZK09, VT95, WYLW16, WWS16, XLH16, YUM10, YSS^{+}16]. methodological [TLB06]. Methodologies [MH10]. Methodology [CF86, BSMB14, CD93, HC08, HCHD06, MPC91, NCB06]. Methods [BBKT87, KJI13, BCPS10, BT98, HM04, MKK15, MS97, SSH99]. metrics [HCD04b]. microcomputer [Red88]. microcomputers [Red91]. middleware [GVBVME13, GVBVME14, KLMS02, PDK95, vRB11]. migration [KKY04, LCWC14, TP09, WLG^{+}11]. MIMD [CC98b]. MIMO [ZZZZ15]. min [RSP91]. min-cut [RSP91]. mine [BqQHj^{+}15]. mini [Red88]. mini/microcomputer [Red88]. Minimal [Lee04, CH03, CH04, Din04, LSM96]. minimal-cut [LSM96]. minimize [QAS91, SMSJ10]. Minimizing [LS94]. minimum [MYYW06, PTX^{+}09]. Mining [TRLD10, BYX16, Spi99, THDC08]. minisupercomputer [Rav92, Rav95]. misleading [Kra97]. Mobilaction [KDP05]. Mobile [CKJ10, MH10, AMuRKK14, CHL^{+}12, CCdF^{+}07, DD11, DG01b, GVCV13, HL16, KC10, KDP05, LCWC14, LKK14, MCSV10, PYHO04, QKZ^{+}07, RSDD10, SN09, SARAL05, SSB^{+}07, SIGC15, WST05, WCJ09, XQZ11, YOM^{+}12, YC13, ZXTS10]. Mobility [HZS12, GQW^{+}14, LWL07, YLS12]. mode [CHZ11, LCWC12, Wu16]. mode-WKBZ [LCWC12]. Model [BGFL08, HOGS11, LDF86, MYU10, XC07, AR04, Apo00, BDPV90, CHKW99, CMMA03, CMMA04, CHL^{+}12, CM04, DJ96, DJ02, FXZ16, FSA97, GE97, GVBVME13, GVCV13, GVBVME14, Gut01, HHCASEQ11, IMA14, IT07, Kar91, KBG^{+}94, KHMK98, KVG08, LWHS06, Li99, LCX08, LOZ01, LWL07, MJRIV14, MLGC94, MYYW06, NXG^{+}16, Pap98, PONA11, RDRLG^{+}15, SLA^{+}11, SD99, SS99, Sto90, TM94, TV08, WD90, Zha16, ZLZL16, Zhe03]. Model-driven [HOGS11, RDRLG^{+}15]. Modeling [GQW^{+}14, HHD98, Joe00, KHMK98, OESHK07, OESHK08, PONA11, WCD99, XCHY11, AD04, EKA06, GW01, GCJP03, GZ06, HMF93, ND07, PN09, PVRM11, SHGA13]. Modelling [Bun98, BWB^{+}01, Dow94, Fin90, HS85, Kam93, PBS86, PM94, YDF87, AHBU07, BNSM01, HK98, Kat92, Kof05, KRSS16, MVA09, RJS06, Sha09]. Models [MH10, Amm90, CGR^{+}09, CO91, FCD03, FCD04, HCK90, JTN95, RCNL15, TD92, XS02, You01, vdA01]. Modified [KBB09, MVA09]. Modular [KBG^{+}94, MHH93, Joh96, LC98, LC00b, DHH96]. Monitor [GLK89, JOR^{+}00, XZ11, YH97]. Monitoring [HRG87, DD11, LPTH99, SDC99, TC06, WSZG15]. monotonic [NK03]. Monte [SSP03]. Moony [LYL95]. morphology [LS94]. motion [BLY10, OO12]. motion-aware [BLY10]. move [WC08b]. moving [MYYW06]. MPEG [KSH96]. MPEG-coded [KSH96]. MPLS [SYY^{+}10]. MR [HMF93]. MR-1 [HMF93]. MT [TJY^{+}11]. MT-BTRIMER [TJY^{+}11]. Multi [CMZ10, GL06, GKK03, ZKWM14, And98, HC08, HBL16, IMA14, IT07, KHA06b, LJL^{+}12, LTLH08, LP03, LC00b, MG02, NCB06, PVRM11, RMR00, SW12, SYY^{+}10, SA03, TL95, TJY^{+}11, WGK^{+}11, XZ11, YOM^{+}12]. Multi-Agent [GL06, GKK03, HC08, KHA06b, NCB06]. Multi-category [CMZ10]. multi-channel [SA03]. multi-code [MG02]. multi-core [IMA14, WGK^{+}11]. multi-domain [LTLH08]. multi-exponentiation [LC00b]. multi-join [TL95]. multi-level [IMA14, LP03]. multi-objective [And98]. multi-OS [PVRM11]. multi-party [SW12]. multi-path [LJL^{+}12, SYY^{+}10]. multi-platform [XZ11]. multi-source [IT07]. Multi-stage [ZKWM14]. multi-threaded [TJY^{+}11]. multi-workflow [RMR00]. multiaccess [Bed88]. multiagents [ACB09]. multibus [Kam93]. Multicast [CL96, CSXC11, CS15, CL07, XM97]. multicasting [DG01a, RM98]. multiclass [GMF96, SDN14]. multicomputer [KH96, WMFK95]. multidatabase [YMST98]. multidimensional [CY94, Cho11, KW01, KW02]. Multigauge [Yan93]. multikey [SK92]. Multimedia [GET10, LQ14b, CKJ10, CHKW99, FAL^{+}01, FCGC13, IT07, KHMK98, KJKK07, KFW04, KFW06, LPY03, LNS^{+}07, Pap98, SSB^{+}07]. multimodal [MCSV14]. Multiple [JWZ14, ADSP12, CC95b, GS01, KNEDK92, KC94, LY12, Li98, MCMM95, PS95, QYCG10, TX16, TH97, YLZ11]. multiplexer [CH97, SB96b]. multiplexing [KSH96]. multiplication [LC98, SBK^{+}92]. Multiport [WK85]. Multiprocessor [DHH96, Gai87, IYD87, MHPS96, Ric90, SE85, Bat00, CSQW14, CPS85, GP90, HMF93, LM91, Li99, LC96b, LCH92, LJ97, SG98, WBA94, WF92]. multiprocessor-based [CSQW14]. Multiprocessors [SK86, Agu03, Kat92, Zhu96, ZS06]. Multiprogramming [DG87]. multiqueues [RS88]. multisend [CS15]. multiservice [ELG00]. multisignature [HCC98, Wu01]. multisignatures [CL00]. multistage [PLL00, XM97, YS90]. Multitasking [Kar86, Kar87, Kar91, KH96]. multivalued [SHGA13]. multivariate [MK11]. Multiversion [Son87]. multiwavelet [CZY^{+}03]. multiword [Met16]. mutation [BXST12]. MuTEAM [CPS85]. mutual [CG00, HDPC13, Hof93, JM99, MS96, SM90, VK08]. MVS [EGH^{+}86].

n [ASDOK03]. n-cube [ASDOK03]. natural [GLT05, Kof05, Lam09, LB03]. nearest [ZHZ^{+}15]. Neat [You01]. needs [The94]. Negative [TRLD10]. Negotiating [DSS00]. neighbor [ZHZ^{+}15]. Neighbour [Met16]. Net [KRSS16, VB87, BDPV90, HCK90, MYYW06, Sto90, You01, ZD07]. Net-Based [VB87]. Nets [CF86, WDF87, BT98, CJV01, DJ02, HMF93, HF94, LJ96, ND07, PM94, RMR00]. Network [HJS87, LSG09, LWD04, ON12, ZXG08, Ang16, Apo00, BBMC98, CO91, CP90, GKK03, Hsu99, HL99, Ito94, JSS^{+}98, KL00, KBB09, LCWC14, LTKK09, Lee04, LH98, LM92, MRDF91, MC95, MG95, PLL00, RS88, SL97, SVN96, SS13, TEQ11, Wei95, WCJ09, XLW11, YLS12, ZL15]. network-based [YLS12]. networking [JIB03]. Networks [BI86, BG87, GET10, Kar86, Kar87, LL10, MA97, PL88, ASDOK03, AMuRKK14, And98, BXST12, CW91, CWY98, CH03, CH04, CC97, CCDL09, Cho11, CT93, CL96, DG01a, DS97, ELG00, GQW^{+}14, GS10, GFH^{+}10, GGDG01, HF08, Hof93, HLLC96, IT07, Kam93, KS97, KTV14, KC10, KKHN12, KKHN13, KA97, LMI90, LT97, LC00a, LJL^{+}12, LHC08, LSBW14, LY95, LC96a, LH02, LJ97, MLK15, MOK03, MCMM95, ON12, PYHO04, PM94, QYCG10, QKSN09, SLPK02, SYY^{+}10, SF02, SSB^{+}07, SM90, SBK^{+}92, TDGNH97, TJS15, TZL96, Vla07, WL01, WC08a, WC08b, WH11, WWS16, WW01, XM97, XLSM10, YS90, Yan00, Yan02, YOM^{+}12, YDR^{+}09, ZYM96, ZZZZ15]. neural [PN09, TEQ11, Wei95, ZL15, LWD04]. neural-genetic [PN09]. neuron [Wei95]. nine [HABJ05]. NN [KC10, YLK11]. node [HL16, PTX^{+}09, TK00, WCLC04]. nodes [KKEG^{+}09, Red91, YOM^{+}12]. noise [RES^{+}16b]. Non [BT98, LT97, Ray00, ERY^{+}16, FSA97, ON12, RES^{+}16a, RES^{+}16b, RS14, RS93]. Non-blocking [LT97, Ray00]. non-classical [RS14]. non-cooperative [ON12]. Non-exponential [BT98]. non-program [RS93]. non-uniform [FSA97]. non-uniformly [ERY^{+}16, RES^{+}16a, RES^{+}16b]. nonhomogeneous [Gut01]. nonnegative [JWZ14]. nonsubsampled [ZHZ^{+}15]. norm [Meg90]. norm-reducing [Meg90]. normal [LCWC12]. Notes [You01]. nothing [TL95, Tan97]. novel [HK98, LQLL11, LWL07, MNK14, WCJ09, WWDL09, WWD04]. Novell [SL97]. NP [Kap92]. NP-complete [Kap92]. number [CS97, Kra97, SVL02, SL02, TDGNH97]. numerical [DJ96, LJ96].

O [Ano87, SIGC15]. OASIS [HSS^{+}11]. Object [HKL99, KHL00, Pap99, SK98, THT98, TP09, dA99, AR92, BGNP01, BT92, CJV01, CD94, CMZ10, CL92, DKP95, Eli13, FAL^{+}01, FCD03, FCD04, GVCV13, GMN98, GW01, HPF02, HHD98, Hin99, JPMAB00, KHR^{+}09, KLMS02, KJKK07, LSR^{+}92, LYW93, LNS^{+}07, LWL07, ND07, NHK14, RXR99, SZM^{+}01, SSH99, SED^{+}99, SS99, TD92, TR02b, TKCR04, XS02, YMST98]. Object-based [THT98, NHK14]. Object-Oriented [HKL99, SK98, KHL00, Pap99, dA99, AR92, BGNP01, BT92, CJV01, CL92, Eli13, FCD03, FCD04, GMN98, GW01, HHD98, Hin99, LSR^{+}92, ND07, SSH99, SED^{+}99, SS99, TR02b]. object-relational [TKCR04, YMST98]. objectbase [Mul92]. objective [And98]. Objects [GET10, BT92, KLK01, KC10, KSK99, MMO00, MYYW06, RS93, TRAR02, ZJ02]. Obligation [ND07]. oblivious [Kha06a]. observation [KS08]. Observed [LBLB13]. Occam [CMM91, CPV90]. ocean [LCWC12]. OCoN [GW01]. ODMG [SD99]. off [LH02]. office [MPC91]. offline [KTK14]. offloading [CS15]. offs [DYB91, LSBW14]. offshore [Kat14]. Okinawa [Kat14]. OMT [MN98]. on-board [VG00]. on-demand [WCLC04]. One [WLC93, WBY^{+}13]. One-dimensional [WBY^{+}13]. One-shot [WLC93]. Online [VM98, CZT^{+}16, HBL16, Mor07, SYY^{+}10]. only [Apo92]. Onto [HC08]. Onto-agent [HC08]. Ontological [MCD^{+}15, RCNL15]. Ontology [Hin99, AKT12, BWG06, HC08, KSL12, SL12, TLB06, WWDL09]. ontology-based [HC08]. ontology-driven [AKT12]. OO [SD99]. OORHS [YW98b]. Open [DBBA10, Kar86, BCPS10, FM10, NPR10, PC13, TJS15]. open-source [BCPS10, PC13]. OpenHMI [NPR10]. OpenHMI-tester [NPR10]. Operating [Gai87, GVCV13, JC97, KS99, KKK^{+}02, MW90, SZ03]. Operational [MB88]. Operations [EGH^{+}86, CS15, Kha06a, LS97]. Optimal [AL03, Car15, DDL^{+}90, LM91, SKR08, TR06, Cha02, HGR12, HA08, OSZZ97, Pap94, QKSN09]. optimally [LY95]. optimisation [GH11]. optimization [CHCL90, CY90, HM04, KS08, LWD04, MMH16, Tan97, TKCR04, VKAR15, ZLZL16, ZZQ95]. Optimizing [SZ03, GH11]. Optmized [AKT12]. Or-BAC [CM04]. ordering [CL96]. organic [vRB11]. organisational [CH06]. organising [HPTC11]. organization [CC14, KW01]. organizational [DSS00, GAHL00]. organizations [KTJ14, LL16]. Oriented [HKL99, SKV12, SK98, AR92, BGNP01, BT92, CJV01, CD94, CL92, CFLZ06, CH06, Eli13, FCD03, FCD04, GMN98, GW01, GD06, HPF02, HHD98, Hin99, JPMAB00, KHL00, LSR^{+}92, ND07, Pap99, RJS06, SSH99, SED^{+}99, SS99, TD92, TR02b, TC06, TLB06, TJ01, WLXD15, dA99]. original [GHW15]. OSGi [RW11]. outlier [LJW^{+}15, MKK15]. output [CH97]. outsourced [Kha06a]. outsourcing [CF06, GCM14, Kat14]. overcorrection [YET94]. overflow [LWD04]. overhead [Hsu94]. Overlay [GET10, ON12]. Overview [SD99, BT98, MMH16].

P2P [HCD04a, ASH12, CZ09, GET10, Ham08, HF08, HCD04b, LM09, SL09]. Packed [UPSL^{+}13]. packet [LJL^{+}12, XLSM10]. pages [ZLL^{+}15]. pair [CCC05, GS10]. pair-wise [GS10]. Pairing [HCD08a, HCD08b]. Pairing-based [HCD08a, HCD08b]. pairings [VK08]. PAKE [HTJ08]. paper [RHS^{+}14]. paradigm [CD94, Sha96]. Parallel [BSK85, CP90, DD90, GE97, GZG95, HM04, IT07, LSG09, LM92, LD96, MRSW98, RSP91, Rav92, Rav95, SZ91, SG98, TR02a, TR02b, TRT04, YZP97, Agu03, AL03, BDMV97, BPX06, BYX16, CSQW14, DJ96, DMP98, DZG92, EG97, HK98, JTL01, JSS^{+}98, LP93, LH98, LOZ01, LL94, LC00b, LS97, MK90, Meg90, Nem96, QAS91, QA93, QZK09, QY15, QAY09, RSG96, SY90, Sha96, SIGC15, Sta93, TZC09, TH94, WYT01, Zom98, FC99]. Parallelism [ZZQ95, Pos98, WY96]. Parallelization [LCWC12]. parallelized [WGK^{+}11]. Parameter [ADMB15, LJW^{+}15, RES^{+}16a]. parameters [BBMC98]. ParCE [AR92]. PARIS [CPT08]. parity [YWD00]. parsing [AR92, LM92]. Part [MRDF91, RMDF91, SVL02, SL02]. partial [GCG^{+}13]. particularly [You01]. partition [LHW04, TY97, XLH16]. partition-based [TY97]. partitionable [Li99]. partitioning [GZG95, KR95, TZC09]. party [SW12]. PASSI [CCSS06]. Passing [HCY87, DAFG95, RS88, TTD10]. password [HTJ08, LYL12a, NHK14]. password-authenticated [HTJ08]. Path [HL99, BGM98, HGR12, KHL00, LY92, LJL^{+}12, RM02, RKZ99, SYY^{+}10, TZC09]. path-tables [RM02]. pathfinding [LY12]. paths [RM02, WC08b]. pattern [BYX16, HK90, MK90]. Patterns [OSN02, SSH99, SD99, XCHY11]. patterns-based [SD99]. PC [SL97]. PEARL [HPF02]. Peer [BBO08, CPT08, KKY04, MVA09, YDR^{+}09, HCD04a]. peer-assisted [YDR^{+}09]. Peer-to-peer [BBO08, CPT08, KKY04, HCD04a]. people [CP15, ZKWM14]. per-VC [ACA99]. perception [BqQHj^{+}15]. Performability [IDY88]. Performance [Apo92, BVCW96, DG87, DD98, FG97, HMF93, KNP16, Kat92, KKEG^{+}09, KK96, LMI90, LYW93, LH98, LH97, LCH92, MN14, MZ03, MOK03, PBS86, RM98, RP90, SK86, SE85, SM90, Sta93, SPB11, TTD10, TZL96, WD90, XM97, AK92, CS15, CHCL90, CCdF^{+}07, CQ96, FSA97, Hof93, JTL01, KW02, Kim03, KS08, KS04, KSH96, Li99, LNR04, MCMM95, OSN02, Omo91, Red91, RD92, SB96b, Tho13, WW01, YS90, ZLZL16, SBK^{+}92]. performance/reliability [YS90]. period [BLY10]. period-based [BLY10]. periodic [DD98, OESHK07, OESHK08]. Permission [JM99]. Permission-based [JM99]. permutation [AGP16, MS97, MC95]. permutations [HL99]. person [PN10, PN07]. person-centric [PN10, PN07]. personal [CF06]. personalization [ABP09]. Personalized [KSL12]. perspective [DE00, LCWY11, TJS15]. Pesa [SZ91]. Pesa-I [SZ91]. Petri [BDPV90, BT98, CJV01, CF86, DJ02, HCK90, HMF93, HF94, KRSS16, LJ96, MYYW06, PM94, RMR00, Sto90, WDF87, You01, ZD07]. Petri-net [MYYW06]. Phase [HK90, BSMB14, CD93, Tan97]. phases [BCL^{+}01]. Phishing [KJI13]. phones [WYO12]. Physical [XSZ16, VG00]. Physically [Wal90]. picture [WYO12]. piecewise [MK10]. pipeline [Omo91]. Pipelined [SBK^{+}92, CC98b, zLTC95, SSC97, TL95]. Pipelining [DGGE88]. placement [BXST12, HA08, LS04, RSP91, Rav92, Rav95, SYNB99, TF06]. planning [ACB09]. platform [GMC^{+}13, NPR10, NCB06, TCD07, WLG^{+}11, XZ11]. platforms [AL03, PVRM11, SC10, SC11, Tam08]. Point [RMDF91]. points [EG97, LKK14]. Poisson [SE07]. policy [LWL07, WSL93]. polygonal [MN14]. polynomials [CW91, LC96a]. Port [SL88]. portable [FH02a, FH02b, MCSV10]. portion [HL16]. positioning [IST^{+}13]. possible [AGP16]. postures [ZHZ^{+}15]. power [Lee09, QAS91, ZZZZ15]. Practical [WBA94, Hui00, Kha06a, WWS16]. practices [CL16]. Pre [LPY03, CCZ03, GFZ^{+}16, YLK11]. pre-computed [YLK11]. Pre-emptive [LPY03, CCZ03]. pre-split [GFZ^{+}16]. Precedence [WL91, Hsu94]. Precedence-agreement [WL91]. precision [KLMS02]. predicates [HF08, SF11]. Predictable [Rom98, ABRW93, PRSM01]. Predicting [LNR04]. prediction [CQ96, Dow94, Li98, SDN14, ZLZL16]. predictions [HOGS11]. preemption [IST^{+}13]. Preface [GATT12, KN12, LQ14a]. preferences [HCL^{+}06]. prefetching [WYT01]. prefix [CC98b]. Preparations [FJ04]. presence [LNR04, TK00]. preserving [YWS06]. prevent [KKP11]. prevention [AR04]. pricing [HWH^{+}16]. prime [RS14, PBS86]. primitives [DAFG95]. principles [EKA06, KS99]. Prioritized [MCMM95, ELG00, Wan97].

Priority [BG87, Hel13, HYY88, KHL^{+}08, SE85, CCZ03, LH97, SED^{+}99, SB96b]. Priority-based [Hel13]. Priority-Driven [SE85]. Privacy [DM06, LBLB13, OO06, UI06, CL16, LYA06, YWS06]. privacy-preserving [YWS06]. private [XLH16]. probabilistic [Amm90, CQN^{+}16]. problem [CL07, GZ06, LS04, LWL93, You01]. problem-centered [You01]. Problems [BD96, Dow87, DDL^{+}90, Kap92, Yen08, You01]. procedure [Sto90]. process [CCSS06, FM10, GH11, KiMKT94, LY96, MVA09, PS10, TM94, TMNL08, TP09, ZLZL16]. processes [Hel13, RFMP07]. Processing [LQ14b, AD13, AD14, BLY10, CHE12, CMM91, DGGE88, DYB91, GKS06, JTL01, Kap92, KW01, KC10, Kof05, MLK15, MLGC94, QAS91, QAY09, SY90, SSC97, TL95, Tan97, TR02a, TRT04, WYO12, WLC93, WBY^{+}13, YWL^{+}03, ZXTS10]. processor [BSS90, HWL97, LP93, MHHP92, OSZ95, WGK^{+}11, YCFX09, Zhu96, CLO92]. processors [Hsu94, IMA14, KP90, MHPS96, PS95]. product [LP03, YWS06]. Production [SZ91, ACB09]. Professionals [Lee85]. Program [Sha88, CWS^{+}00, RS93]. Programmable [BSK85, Ang16, XFA99]. Programme [Lee85]. Programming [MÁGA00, YUM10, AdPT06, HPF02, KLMS02, LH98, MMD99, Pap98, YW98a, YW98b]. programs [Agu03, BST98, BA12, Duj91, EG97, HCK90]. Progressive [LBLB13]. project [SD99, GPR10]. projective [AM00]. projects [ÇG11, FM10]. Propagation [LDF86, Mae10, XFZ09, DSS00]. properties [GQW^{+}14, LB96, LT97, LMPJ12, SS93]. proposal [Jen01]. Prosumer [RPD^{+}13, RPDH15]. Protecting [GMP13]. protection [OO06]. Protocol [ET10, HCY87, HYY88, BWB^{+}01, CDFP90, CCZ03, GS10, HTJ08, JC02, LJ96, LHC08, Ray00, RW11, SC10, SC11, VT95, WP95, WCLC04, WH11, YWS06, ZJ02]. protocols [AMuRKK14, GS01, HDPC13, Koc94, LMI90, LH98, MG02, NC03, NC04, SED^{+}99, WD90, WCD99]. Prototyping [Abi00]. Provably [Pap94]. provide [CL96]. provides [LWD04, YW98b]. providing [CCdF^{+}07]. provisioning [WC08a]. provisions [WCL15]. Proxy [LS04]. PSC [YSS^{+}16]. Pseudo [ML86]. Pseudo-Exhaustive [ML86]. PSQL [XFA99]. public [FBZS12, HCD08a, HCD08b, LL16, Wu01]. public-key [HCD08a, HCD08b]. publication [XLH16]. purely [YSAMB96]. Purpose [ET10, KKK^{+}02, War07]. Purpose-based [ET10]. PVM [WLC07]. PWSMS [BBO08]. pyramid [CH03, CH04, MG95].

QoS [ADSP12, CL07, GVBVME13, GVBVME14, JIB03, PYHO04, SSB^{+}07, WC08a, WLXD15]. QoS-aware [PYHO04]. QR [WYO12]. quad [KBB09]. Quality [Lee85, SV13, BWG06, HL16, SZM^{+}01, TYW09, ZL16]. quality-aware [BWG06]. quantitative [BSMB14]. quantization [CZY^{+}03]. queries [AKT12, CCC05, Dan07, HF08, SF11, SK92, TL95, WST05, WLC93]. Query [KHL00, AKT12, BKBM10, BKBM13, BST98, CU07, CZT^{+}16, CHE12, CY90, GKS06, HM04, JTL01, KC10, LPTH99, MLK15, MMH16, Tan97, TR02a, TRT04, TKCR04, YWL^{+}03, ZXTS10, ZZQ95]. querying [FJ04, Kat05, LCWY11, VSZMCV08]. queue [KC94, SZ03]. Queueing [BI86, LP93, CT93, CO91]. queues [CO91]. queuing [ACA99, KL00]. quickest [RM02]. quorums [CG00].

radial [HA08]. radio [GQW^{+}14]. RAID [TX16, YWD00]. RAID-5 [YWD00]. RAID5 [Tho97]. railway [BGFL08, BNSM01]. random [Hel13, Wal90]. ranking [NHK14]. rapid [JTN95]. Rate [RSBW14, CZY^{+}03, LJL^{+}12, NK03, TR06, TZL96, ZYM96]. rate-allocation [LJL^{+}12]. rate-based [TZL96, ZYM96]. rate-monotonic [NK03]. rates [LJL^{+}12]. RBAC [CHL^{+}12]. RDCT [WWD04]. RDF [CH11, KFW04]. re [DR09, ZKWM14]. re-identification [ZKWM14]. re-weighting [DR09]. Reactor [DKO86]. read [Apo92]. read-only [Apo92]. readers [HDPC13]. readiness [ÇG11]. readings [MLK15]. ready [SZ03]. Real [DKO86, DZG92, HKL99, HA08, KG07, SK98, SK86, TRAR02, WDF87, ABRW93, BGNP01, CZT^{+}16, CPV90, FAL^{+}01, GVBVME13, GVBVME14, GH11, GMC^{+}13, HPF02, HHD98, HSS^{+}11, IT07, Jen01, KBG^{+}94, KLK01, KLMS02, KB02, KKK^{+}02, KJKK07, KSK99, LPY03, OESHK07, OESHK08, PONA11, RW11, Rom98, RXR99, Sel02, SZ03, SSC97, SED^{+}99, SS99, VM98, WH11, WGK^{+}11, WWDL09, Wu16, ZZWD15]. Real-coded [HA08]. Real-Time [DKO86, HKL99, SK98, SK86, WDF87, TRAR02, ABRW93, BGNP01, CZT^{+}16, CPV90, FAL^{+}01, GVBVME13, GVBVME14, GH11, GMC^{+}13, HHD98, HSS^{+}11, IT07, Jen01, KBG^{+}94, KLK01, KLMS02, KB02, KKK^{+}02, KJKK07, LPY03, OESHK07, OESHK08, PONA11, RW11, RXR99, Sel02, SZ03, SSC97, SED^{+}99, SS99, VM98, WH11, WGK^{+}11, WWDL09, Wu16, ZZWD15]. realistic [KHA06b]. Realization [SL88, HBL16]. realizing [KS99]. rearrangeable [HL99]. reasoning [FCHJ05, KNEDK92]. received [EQ11]. receives [KT95]. receiving [HCC98]. Recognition [ZHZ^{+}15, LKK14]. recognizer [MCSV10]. recommendation [CZT^{+}16, KSL12, Mor07]. Reconfigurable [DDL91, CCDL09, KR92, LS97, OSZ95, SB96a, TCD07, WMFK95, WWD04]. reconfiguration [BSS90, BDMV97, Joh96]. Reconstruction [San90, YUM10]. Recovery [CKT94, Rom94, Rom98, AJTT15, BDMV97, DG01b, Hsu99, KB02, LYL12a, NMMS02, RW00, RS93, Rom02, Wu01]. recruitment [RPD^{+}13, RPDH15]. recurrence [Che93]. recurrent [Sha96]. recyclable [RHS^{+}14]. Redistribution [LSD95]. reduce [AKT12, TF06]. reduced [Buc00, RSP91]. reducing [Meg90]. reduction [JKC16, KKHN12, KKHN13]. redundant [Gut01, MP88]. referrers [QL13, QL15]. refinement [CU07]. reflection [TV08]. reflective [PXY93]. region [ZEZF11]. regularization [JWZ14]. rekeying [GS10]. related [LL16, TZL96, XSZ16]. relation [CMZ10, CT93, SVL02, SL02]. relational [LY92, TKCR04, YMST98, ZT03]. relationship [SHGA13]. relationships [TD92]. relay [RSBW14, ZZZZ15]. release [GHW15]. Relevance [BCLG98]. Reliability [Dow87, YQ95, CD94, Dow94, Gut01, Li98, Sah95, Sha09, Wan03, YS90]. Reliable [LLHG06, JSM96, SN09, Wei95, WWS16]. remainder [WWH95]. remanufacturing [Zha16]. remote [LYW93, WSZG15, YQ95]. removal [MKK15]. renewal [MVA09]. rental [FBZS12]. repairable [Bat00]. replacement [LNS^{+}07, ZS06]. Replicated [SA86, AM00, JSM96, MÁGA00, ZJ02]. Replication [Son87, GCG^{+}13, Koc94, KDP05, LYW93, NMMS02, The94, WY02]. Representation [VB87, DR09, LQX^{+}14, MCD^{+}15, QLL14, TN93]. representations [Buc00]. reputation [PC13, WCL15, WWS16]. reputation-based [WCL15]. Request [LTLH08, SSP03]. Request-driven [LTLH08]. requirement [HWL97, TR06]. requirements [AST99, ALP99, CSB16, FCHJ05, GLT05, HABJ05, Lam09, TC05, TMNL08]. Rescheduling [GHW15]. Research [GFZ^{+}16, SWJF11, WST05, CKJ10, You01]. reservation [CCZ03]. resilience [SF02]. resilient [GS10]. resistant [BNP05]. resolution [LL94, MCMM95]. Resource [CHE12, GVCV13, ASH12, Che07, Hel13, HOGS11, JDG92, JWH^{+}15, KC94, LM91, LYL^{+}12b, PM94, Red88, SPB11]. resource-sharing [Red88]. resources [Wu16]. Respect [DG87]. Response [Bat00, LSD95]. Responsive [PM99]. Restoration [CCDL09]. Restricted [Yen08, AHBU07]. restructuring [KP90]. resulting [dA99]. results [KS08, LL94, Lov96]. RETE [MK90]. RETRAN [Sha96]. Retrieval [LQ14b, DR09, FXZ16, LQX^{+}14, MM15, OO12, QZK09, WWD04]. retrieving [HHCASEQ11]. reuse [SVL02, SL02, TMNL08]. reuse-based [TMNL08]. reutilization [MOEMK^{+}15]. reveal [SSH99]. revenue [HWH^{+}16]. reversible [MNK14, YW98b]. Review [Ano01a, Dow87, RG88, Toh94, Ano02, GCM14]. revisited [MOZR03]. rework [GHW15]. rewriting [Pap99, ZT03]. RFID [HDPC13, YHLC13]. RHS [WY96]. rights [AHBU07]. rigorous [ND07]. Ring [BG87, Bow86, Din04, PLL00, SBK^{+}92]. ring-banyan [PLL00]. RIO [SLPK02]. risk [BSMB14, FCHJ05, HCHD06, HDCH10, SMSJ10, YSS^{+}16]. risk-based [HDCH10]. risk-informed [FCHJ05]. road [RW11]. robot [TEQ11]. Robust [LJW^{+}15, CPS85, FM14]. role [BCPS10, Cer04, Cra06, LTLH08]. role-based [Cra06]. rotation [HJXY12]. rough [FL16]. roundtrip [LWL07]. route [LS04, SSS01]. routed [ACR00]. router [BXST12]. Routing [AMuRKK14, BI86, DS97, ASDOK03, ACR00, BD96, BGM98, BDMV97, CH03, CH04, CL07, CLO92, JIB03, LT97, LHC08, MS97, MC95, Pap94, PYHO04, QKZ^{+}07, SYY^{+}10, WP95, WCLC04, WH11, WYLW16, WY02, XQZ11, YS99]. routing-based [WY02]. row [CLO92]. RPS [HS85]. RSA [WLC07]. RTL [ZCL03]. RTOS [WGK^{+}11]. Rules [TRLD10, BBMC98, PV96, SL14]. run [DE00]. run-time [DE00]. running [WW01]. runtime [HOGS11].

sabotage [KHL^{+}08]. sabotage-tolerance [KHL^{+}08]. Safe [HPF02, BLY10, LPY03, MK10]. safely [Toh94]. Safety [DKO86, GMC^{+}13, Hal09, SSC91]. Same [IB10]. sampled [LQLL11]. sampling [ERY^{+}16, RES^{+}16a, RES^{+}16b, ZCG^{+}15]. saving [YOM^{+}12]. SBASS [PKC07]. SBDA [KA97]. Scalability [MD03, MD04]. Scalable [ABP09, CSQW14, ASH12, FXZ16, IT07, Koc94, NKWK07, SSP03, YDR^{+}09]. scalar [CZY^{+}03, YWS06]. scale [CPT08, KTJ14, OSN02, ZLL^{+}15]. scaled [ZEZF11]. scaling [CQ96, QYCG10]. scenario [AST99]. scenario-based [AST99]. scenarios [CSB16]. scheduler [SG98]. Scheduling [GET10, Nem96, AL03, BqQHj^{+}15, CSQW14, CH98, Hel13, KKK^{+}02, KHL^{+}08, KKHN12, KKHN13, LY96, LC96b, NK03, SZ03, TY97, Tho13, VM98, WLXD15, WMFK95, Wu16, ZL15, Zhu96]. Schema [GKS06, AP16, CMZ10, MFU02]. Schema-aware [GKS06]. schemas [PHPd98]. Schemata [FCD03, FCD04]. scheme [BDX93, CC95a, CL00, DS97, HJXY12, Hsu08, HWH^{+}16, HCC98, Joh96, Kat14, KBG^{+}94, KB02, KC10, LJL^{+}12, LWL93, LH02, MS96, MHH93, PYHO04, RZ01, SSP03, SLPK02, SYY^{+}10, SSS01, The94, TK00, VK08, Wan03, WLY^{+}14, WCJ09, WWH95, WC01, XLSM10]. schemes [HCD08a, HCD08b, KK96, KA93, RM98, WC08a, Wu01]. science [SWJF11, ZCK11, LL15]. scientific [KKEG^{+}09]. SCR [HABJ05]. SCSI [CHKW99]. SDD [DZ15]. SDD-matcher [DZ15]. SDITPM [LWL07]. search [BPX06, CL07, HF08, Kha06a, LP03, VKAR15, YLK11]. searches [PKC07]. searching [JKC16, Lee09, SK92]. SEC [PA10]. Secondary [Red88]. secret [WLY^{+}14]. secrets [CWY98, WY93]. SECTET [AHBU07]. Secure [BV13, CC95a, ET10, KVR03, KVR04, LWL93, Gab06, JC02, Jon03, KFW06, LTKK09, LYL12a, LHC08, LTLH08, RC98, RFMP07, VK08, WH97, YDR^{+}09, YLS12]. securities [EKA06]. Security [Ye15, BSMB14, CL16, Dam04, DGM06, GKK03, HCD08a, HCD08b, HLLC96, LHC08, NC04, UI06]. secutity [NC03]. seek [TF06]. Segment [PKC07]. segmentation [HHCASEQ11, NXG^{+}16]. Selected [LL10]. selecting [GCM14, YSS^{+}16]. Selection [KJI13, Cha08, CL07, Duj91, Hel13, KC94, LY92, TRT04, VKAR15]. selectively [WY93]. Self [LMPJ12, MC95, SS93, DFT97, FG97, HPTC11, KS00, LT97, Nan94]. self-checking [Nan94]. Self-diagnosability [SS93]. Self-routing [MC95, LT97]. Self-similar [LMPJ12, FG97]. self-stabilizing [DFT97, KS00]. sellers [Mor07]. semantic [CC15, CC16, CPT08, DZ15, MLGC94, PV96, RSG96, RSDD10, WZWD15, XC15, YLQW13, TBEH04]. semantic-driven [DZ15]. Semantics [Dam04, TS99, KFW06, LYA06, ND07, PHPd98, TD92, ZCL03]. Semantics-aware [Dam04]. SemCrypt [GKS06]. semi [CC14, WLC93, YWW96]. semi-hypermedia [YWW96]. semi-join [WLC93]. semi-structured [CC14]. sender [TJ01]. Sensitive [GMP13]. Sensitivity [DG87, MB88]. Sensor [LL10, AMuRKK14, Cho11, GS10, KKHN12, KKHN13, LJL^{+}12, LSBW14, MLK15, QYCG10, QKSN09, WH11, WWS16, WCJ09, XLSM10, YOM^{+}12]. separate [WY93]. separation [FL16]. sequence [PKC07]. sequences [MN14]. sequencing [SF11]. sequential [HN94, MNK14, Ray02, SW12]. Serc [PBS86]. serial [SVN96]. serialization [TN93]. series [KS08, ML09a, ML09b, MK10]. serious [MOEMK^{+}15]. server [AHH^{+}16, HDPC13, HFS^{+}08, TH94]. servers [Bul11]. Service [CMMA03, CMMA04, EKA06, HCL^{+}06, HRG87, LBLB13, SKV12, AdPT06, BBO08, BWG06, BVCW96, CFLZ06, CH06, DGM06, GD06, KKY04, KC94, KLMS02, KI09, LCWC14, LV08, LWL07, LLHG06, OSN02, ON12, RDRLG^{+}15, SL12, SZM^{+}01, SLPK02, SF02, SPBK12, TZL96, WCL15, XCHY11, Ye15, ZMMS04, ZL16]. service-oriented [CFLZ06, CH06, GD06]. Services [LdOS13, MH10, CCdF^{+}07, EKA06, Kat05, KHMK98, KKLK02, KFW06, KVR03, KVR04, PM99, PRSM01, SLPK02, SC09, SW06, WOE06, YLQW13]. Session [QL13, QL15]. Set [AJTT15, Cha02, SED^{+}99]. sets [YLK11]. setup [WCLC04]. SGHC [LY95]. shape [LQX^{+}14]. shaping [CH97]. Shared [DKP95, Gai87, SV13, Che07, DAFG95, DMP98, FSA97, KH96, KA97, LYL95, SG98, TL95, Tan97, Wu16]. shared-buffer [KA97]. shared-bus [KH96]. shared-memory [DMP98]. shared-nothing [TL95, Tan97]. sharing [BBO08, CDY92, CPT08, Fin90, FMP94, RSBW14, LCH92, PM94, Red88, WLY^{+}14]. SHDS [YWW96]. shedding [Cha08]. shield [UI06]. ship [YSS^{+}16]. ship-selecting [YSS^{+}16]. shorten [LWD04, WWD04]. shortens [LWL07]. shortest [BGM98, TZC09]. shot [WLC93]. should [BKBM10, BKBM13]. shuffle [KR95]. shuffle-exchange [KR95]. Sicily [GPR10]. side [FH02a, FH02b]. signal [Pap94]. signalling [MRDF91, RMDF91]. signature [Wu01, YLS12]. signature/multisignature [Wu01]. signatures [JLL01, LDZP08, LW10]. significance [FMS^{+}11]. Significant [VT95]. SIM [AAKD09]. SIM-based [AAKD09]. SIMD [RSP91]. similar [FG97, LMPJ12]. similarity [Dan07]. Simple [Bow86]. Simulated [BBKT87, Din04, RP90]. Simulating [DAFG95]. Simulation [DFT97, Kar86, Kar87, Kar91, Mae10, WMFK95, Zom98, BCLG98, DMP98, GKK03, HK98, KL00, KHA06b, LD96, NCMH00, PVRM11, QY15, ZL16]. simulations [CWBZ11, Men99, OSN02]. simulator [SCL97]. simulators [JSS^{+}98]. simultaneously [Wu01]. Single [CLO92, MG02, ACA99, GHW15, Kra97, YOM^{+}12]. Single-code [MG02]. single-machine [GHW15]. site [LSD95]. situation [CP15]. size [WWD04]. sizeable [WW01]. Skew [TL95]. skewed [ZS06]. SLA [QAY09]. Slab [Mae10]. slave [TJY^{+}11]. sleep [KKHN12, KKHN13]. sliding [BYX16]. slot [SVL02, SL02]. small [KTJ14]. smart [Hsu08, KHE01, MCD^{+}15, RCNL15, RPD^{+}13, RPDH15, VK08]. smart-grid [RPD^{+}13, RPDH15]. smiling [CCL11]. smooth [HFS^{+}08]. smoothing [FCGC13]. SMTP [Bul11]. SMX [RKZ99]. snooping [WCD99]. SOA [GVBVME13, GVBVME14, LY12, ZCK11]. SOAP [KS04, KVR03, KVR04]. SOAP-based [KVR03, KVR04]. Social [MGOB15, LTKK09, RJS06, WYLW16]. SOFIT [AGT98]. Soft [KKK^{+}02, FL16, SMSJ10]. Software [Dow87, GL06, Kat14, MMI85, NTI^{+}07, SSC91, Abi00, Amm90, AGT98, BCPS10, BDX93, CD94, CWS^{+}00, Dow94, Eli13, GW01, Gut01, Hog04, KTJ14, KG07, KiMKT94, Lam09, Li98, LSD95, LP03, MD03, MD04, QAS91, RW00, Sha09, SCKN11, SS13, TLB06, XS02, YHLC13, You01, Zhe03, dLR01]. software-based [AGT98]. software-implemented [MD03, MD04]. Solid [YUM10]. Soliton [Mae10]. solution [Kha06a, LYL12a, MOZR03, Meg90, SS13, YCFX09, You01]. solution- [You01]. solutions [Car15, DGL00, vdAJ00a]. Solving [Dan07, GZ06, PV96]. some [XQZ11]. SoPC [ZZWD15]. sort [LT97, TR02b]. sort-Clos [LT97]. sort-merge [TR02b]. Sorting [BDG^{+}94, PS95, RHS^{+}14]. Source [DBBA10, BCPS10, FM10, FMS^{+}11, IT07, PC13]. sources [BGRV99]. space [AM00, CC15, CC16, JKC16, WZWD15, Zha16]. spaces [CJV01]. spam [QWK14]. spambots [HPTC11]. spanning [AA99, PTX^{+}09]. sparse [GZG95]. spatial [BLY10, Cho11, CL92, KC10, KDP05, LHW04, MCSV14, WBY^{+}13]. spatio [CCC05]. spatio-temporal [CCC05]. Special [CFLZ06, DGM06, DBBA10, DM06, FHFL07, GL06, GMPR08, Hal09, Ham08, HKL99, KG07, LWHS06, LM09, LSG09, LQ14b, LK16, LL10, MH10, MA97, SKV12, SK98, SWJF11, TS99, XSZ16, ZXG08, Zom98]. specification [Jen01, MLLS94, RDRLG^{+}15, VT95]. specifications [DE00, HABJ05, KSK99, LJ96, LB03, SEuH96]. specified [HCC98]. specify [BGRV99]. Specifying [KFW04]. spectral [Car14]. spectrum [RSBW14]. speculation [FC99]. speech [CHZ11]. Speed [Sch88, Sah95, TTD10]. split [GFZ^{+}16]. splitting [HZS12, SSC97]. sporadic [VM98]. stability [CSXC11]. stabilizing [DFT97, KS00]. stage [CT93, WC08a, ZKWM14]. stages [Li98]. staging [ABNY12]. stamp [CL00]. stamping [CCdF^{+}07]. standard [Duj91, Sel02, YS99]. standard-cell [YS99]. standards [Tam08]. star [GGDG01, Hsu99, IST^{+}13, LB96, SS93]. Start [MYU10]. State [BI86, THDC08, TN93, CJV01, Kam93, OESHK07, OESHK08, Rom02, Zha16]. State-Dependent [BI86]. static [CDY92, DS97]. statistical [LJW^{+}15, MK11]. statistics [LQLL11]. status [SWJF11]. steganographic [LYC02]. stereo [ZZWD15]. stigmergic [CU07]. Stochastic [BDPV90, ERY^{+}16, BT98, Buc00, GQW^{+}14, HMF93, KS97, LNR04, RES^{+}16b]. Storage [MLGC94, PRPS13, CHKW99, Tho13]. Strategies [JSS^{+}98, AD04, SZ03, TL95, WLC93]. Strategy [ML86, RKZ99]. stream [LKK14, ML09a, ML09b]. Streaming [GET10, IT07, KKY04, KJKK07, ML09a, ML09b, SL09, TR06]. streams [Amm90, Cha08, FCGC13, SSP03, SL14]. string [KWL07]. strong [TRAR02]. Strongly [NMMS02, YUM10, RC98]. structural [CH11, KNP16, PV96]. Structure [HS88, MMI85, KHR^{+}09, KWL07, Lee09, SZ03, Sta93, SK92, WBY^{+}13, YA97]. structured [CC14, HF08, LSD95, NKWK07, THDC08]. Structures [Gai87, HOGS11, SB96a]. studies [Kof05]. Study [Kar86, Kar87, KSH96, AMuRKK14, BXST12, CDFP90, EKA06, KKEG^{+}09, RD92, The94]. style [GD06]. sub [FXZ16]. sub-image [FXZ16]. Subjective [Kri01]. Subsequence [KS08, ML09a, ML09b, MK10, PKC07]. substring [Hui00]. subtuplespace [FC99]. subtype [Car14]. summarized [HC13]. Superprocesses [Sha88]. supervision [JC02]. supervisor [KBG^{+}94]. supervisory [TC06]. supplier [GCM14]. supply [SS13]. support [AdPT06, BKBM10, BKBM13, FCHJ05, GMC^{+}13, HFS^{+}08, LP03, MdlFD03, WGK^{+}11, YMST98].

Supporting [HM87, Sha88, Kha06a, KDP05, RW00, SZM^{+}01, TH94]. supports [FMP94]. surveillance [CL16, DD11, XC15, ZEZF11]. survey [AK92, CHE12, HM04, RS14, XFZ09]. switch [And98, CC97, KS97, KA97]. switches [DD98, FG97, FSA97, LY96, LH97, LD96, RM98, TH97]. switching [BGM98, LL96, XM97, ZYM96]. SYN [TV08]. Synchronization [ET10, EG97, KJKK07, Tho97, YET94]. Synchronizing [Gai87]. synchronous [Sto90]. Synthesis [BI86, SEuH96, YSAMB96, ZCL03]. synthesising [Che07]. synthetic [Duj91, SSP03]. System [CKT94, CF86, DKO86, Gai87, HRG87, Kar86, Kar87, MMI85, MYU10, PBS86, SL88, VB87, XSZ16, ACB09, AR04, Apo92, AAKD09, BZLX16, BBO08, BGNP01, Bed88, BCL^{+}01, CF06, CZ09, CCL11, CL92, CI94, CP15, EG97, Fin90, FMP94, Gun93, GP90, HDCH10, IST^{+}13, IMA14, JC97, Joh96, Kar91, KH96, KTV14, KS99, KKK^{+}02, LTKK09, LSD95, LYL95, LPTH99, MRDF91, MLLS94, Mul92, MGOB15, MW90, OO12, OSZ95, SL97, SARAL05, SG98, SDC99, SSH99, Sri91, Tho13, WSZG15, WH97, XFA99, YSS^{+}16, YHLC13, YC13, YMST98, YDR^{+}09, YZP97, YWW96, ZD07, DHH96, HS88]. system-level [SSH99]. System-related [XSZ16]. System/88 [HS88]. Systematic [Che93, WDF87, GCM14, Lov96]. systemC [PONA11, PVRM11]. systemC-AMS [PONA11]. Systems [DGHE88, Elm87, EGH^{+}86, GL06, GMPR08, Ham08, HS85, LDF86, PRPS13, SZ91, SK98, Son87, WDF87, YDF87, ZXG08, Abi00, Amm90, AJTT15, ABRW93, BGFL08, Bat00, BDMV97, BNSM01, BM97, Bun98, CSQW14, CHKW99, CCHL03, Che07, CPV90, CDY92, CHE12, CT93, CH06, CPS85, CMRR02, DAFG95, DYB91, DJ02, DG01b, ERY^{+}16, FC99, GVCV13, GMN98, GW01, GH11, GST03, HC08, HPF02, Hal09, HHD98, HJXY12, HOGS11, HSS^{+}11, HS96, HA08, HBL16, IS97, IDY88, JDG92, JSM96, JC97, KNEDK92, KBG^{+}94, KI09, KRSS16, Kri08, LM91, LY92, LPY03, LH98, Li98, Li99, LCH92, LCWY11, MP88, MJRIV14, MCSV14, MÁGA00, NKWK07, Nan94, Nem96, ND07, NCB06, OESHK07, OESHK08, PONA11, PHPd98, PV96, QAS91, RES^{+}16a, RES^{+}16b, Ray00, Red88, Red91, RS88, Rom94, Rom98, RXR99, SY90, SZ03]. systems [SSC91, SCKN11, SKR08, TL95, Tan97, THT98, TRT04, TN93, TC06, UI06, VM98, WL91, WSL93, WBA94, Wan97, Wan03, Wei95, WF92, ZMMS04, ZL16, dA99]. Systolic [Sch88, Che93, KR92].

T [GMN98]. T-CORBA [GMN98]. tables [RM02]. Tabu [BPX06]. tags [HDPC13]. taking [CZT^{+}16]. Tamper [BNP05]. Tamper-resistant [BNP05]. target [HL16]. task [IS97, LC96b, LSM96, QY15, SN09, SL97, VM98, WLXD15, WMFK95]. tasks [Hsu94, KHL^{+}08, WGK^{+}11, Wu16]. Taxi [CZT^{+}16]. Taxi-taking [CZT^{+}16]. TCM [WWDL09]. TCP [Apo00, KW10, TV08]. TCP/IP [Apo00]. technique [CL00, CO91, KL00, KK96, LWD04, LC00b, MCSV14, MRSW98, ND07]. techniques [BT98, Che93, Dam04, ELG00, FCGC13, KiMKT94, Lam09]. Technologies [XSZ16, GKK03]. technology [BZLX16, CMRR02, DD11, HBL16, JOR^{+}00, LB03, MFU02, QAY09]. tele [KKLK02]. tele-audio [KKLK02]. telemedicine [WWDL09]. teletraffic [LMPJ12]. Temporal [DJ02, CCC05, GE97, RSBW14, HN94, SD99, VT95]. Term [HC13, HWH^{+}16, Pap99]. test [AGP16, CI94, PRB13]. Testability [RG88, Lam09]. testable [KR92]. tester [NPR10]. Testing [BSK85, ML86, Dow94, HCK90, NPR10, SLA^{+}11]. text [ABP09, WZWD15]. texture [KNU12]. their [Joe00, Yan93, dA99]. theorem [WWH95]. theoretical [LP93, XQZ11]. theory [CS97, LYL^{+}12b, SF11, YLQW13]. Things [BqQHj^{+}15]. Thrashing [WYT01]. threaded [TJY^{+}11]. threshold [Hsu08, JLL01]. Throughput [SE07, SVL02, SL02]. throughputs [SL02]. tight [KI09]. tightly [SG98]. tightly-coupled [SG98]. Time [DKO86, HKL99, HJS87, KJKK07, SK98, SK86, WDF87, Apo00, ABRW93, Bat00, BGNP01, CL00, CHL^{+}12, CZT^{+}16, CPV90, CCdF^{+}07, DE00, DG01b, FAL^{+}01, GVBVME13, GVBVME14, GMN98, GH11, GMC^{+}13, HPF02, HHD98, HJXY12, HSS^{+}11, Hui00, IT07, Jen01, KC94, KBG^{+}94, KLK01, KLMS02, KB02, KKK^{+}02, KG07, KS08, KI09, KSK99, LPY03, LWD04, LWL07, MMO00, ML09a, ML09b, MK10, MYYW06, OESHK07, OESHK08, OSZZ97, PONA11, PVRM11, QAS91, QA93, RW11, Rom98, RXR99, Sel02, SZ03, Sha09, SSC97, SED^{+}99, SS99, TF06, TRAR02, VM98, WH11, WGK^{+}11, WWDL09, WWD04, Wu16, ZZWD15, ZMMS04]. time- [OSZZ97]. time-approximate [PVRM11]. time-aware [GMN98]. time-based [DG01b, KJKK07]. time-bounded [KB02]. time-constraint [CHL^{+}12]. time-series [KS08, ML09a, ML09b, MK10]. Time-triggered [KJKK07, PONA11]. Timed [WDF87, Bun98, Pap98, PM94]. timeliness [KS99]. timeliness-guaranteed [KS99]. timely [SN09]. Times [BG87, GHW15, HWL97, Sto90]. Timing [SYNB99, YS99, KKP11, MN98, Nem96]. Timing-driven [YS99]. TLB [Lee09]. TMO [KHMK98, KKLK02, NKWK07]. TMO-based [KKLK02]. TMO-structured [NKWK07]. Token [HCY87, Ray02, MS96, RS88, ZJ02]. Token-based [Ray02, MS96, ZJ02]. token-passing [RS88]. Tolerance [PL88, BDX93, CD94, FMP94, KKLK02, KHL^{+}08, MG95, NCMH00, RW00, TH94, VG00, YA97]. Tolerant [Elm87, GMPR08, IYD87, YDF87, ASDOK03, ACR00, BSS90, BDMV97, CC98a, CH03, CH04, CC95b, CMRR02, DG01a, Gun93, HS88, HS96, JM99, KKY04, KBG^{+}94, KiMKT94, LV08, LY95, LYL95, MRDF91, NMMS02, PLL00, PHPd98, RMDF91, Toh94, WF93, YS90, YET94, ZMMS04]. toleration [Rom98]. TOOBIS [SD99]. tool [ABP09, AGT98, FCHJ05, GLT05, KNEDK92]. toolkit [BGRV99, CMMA03, CMMA04]. Tools [HABJ05, LH98, SLA^{+}11]. Toolset [HABJ05, LSR^{+}92]. Top [CPV90]. Top-down [CPV90]. topic [HHCASEQ11]. Topics [LM09, LL10]. Topological [FL16]. topologies [BGM98, DDL^{+}90]. topology [Ric90, Sah95]. tori [MS97]. trace [FM14]. trace-fuzzy [FM14]. traceback [KTK14]. traces [BA12]. tracking [CCL11, ZEZF11]. trade [DYB91, LSBW14]. trade-offs [DYB91, LSBW14]. Tradeoff [Sch88]. tradeoffs [TN93, YS90]. trading [EKA06, HBL16]. traditional [WWDL09]. traffic [BBMC98, Car15, DD98, FG97, FSA97, GMF96, Kat92, KKHN12, KKHN13, KSH96, KW10, MOK03, MVA09, QKZ^{+}07, RM98, SE07, SYY^{+}10, SB96b, Vla07, XC15, ZS06]. traffic-based [QKZ^{+}07]. trajectories [CZT^{+}16]. transaction [AAKD09, DYB91, KDP05, SARAL05, TYY^{+}12]. Transactional [HDCH10, JPMAB00, KHR^{+}09]. Transactions [BV13, Apo92, HCHD06]. transcoding [LS04]. transducer [KHE01]. Transfer [RMDF91, BVCW96, LWL07]. transfers [Bed88]. transform [CZY^{+}03, TDGNH97, ZHZ^{+}15]. Transformation [LHW04, FCD03, FCD04, GH11, PHPd98].

Transformation-based [LHW04]. transformations [KKP11]. transforms [DZG92]. Transient [YDF87, AHH^{+}16]. transitions [HK90]. translator [TJY^{+}11]. TransLib [JPMAB00]. Transmission [HJS87, GDK88, GFH^{+}10, IT07, KKLK02, WCJ09]. transputer [CPV90, DJ96, DDL91, GP90, LM92, SCL97, SBK^{+}92]. transputer-based [CPV90, DJ96, DDL91, GP90, SCL97]. transputers [CHCL90, CP90, DDL^{+}90, MW90]. Transversal [MOEMK^{+}15]. tree [CSXC11, FXZ16, IMA14, KBB09, MC95, MG95, Sah95, SG90, SB96a, THDC08, Vla07, Yan00]. trees [AA99, CC98a, CC95b, CL96, Kha06a, PTX^{+}09, PRB13, SDN14]. Triangular [CG00, Sta93]. triggered [KJKK07, MNK14, PONA11]. trunking [KW10]. Trust [KTV14, SW06, CL16, CPCB11, HCD04a, PCB10]. Trustable [TBEH04]. TrustBus [FHFL07]. trusted [WCL15]. Trustworthiness [HCD04b, PS10]. TTL [WK85]. TTL-compatible [WK85]. TTM [PONA11]. TTP [KHE01]. TTP/A [KHE01]. tuner [WWD04]. tuning [ADMB15, KKEG^{+}09, LWD04]. tuple [CQN^{+}16]. tuples [Cha08]. tutorial [Ray00]. Two [CD93, IB10, KR95, SCL97, ACR00, BSMB14, JIB03, KR92, KWL07, Kof05, OSZ95, PRB13, Tan97, WC08a, WF92, ZZZZ15]. two-dimensional [ACR00, KR92, OSZ95]. two-fault [WF92]. Two-layer [SCL97]. two-level [JIB03, KWL07]. two-phase [BSMB14, Tan97]. two-stage [WC08a]. Two-way [KR95, ZZZZ15]. type [Zhe03]. Typed [YUM10].

U-Science [LL15]. ubiquitous [BSR11, BNP05, CMMA03, CMMA04, KSL12, MGOB15, NTI^{+}07, TYW09]. Ultrasonic [YUM10]. UMIC [CKJ10].

UML [BCL^{+}01, RFMP07, Sel02]. UML-based [BCL^{+}01]. unbalanced [HA08]. unbounded [ABRW93]. unbreakable [KFW06]. uncertain [CQN^{+}16]. uncertainty [CQN^{+}16]. Unclear [DKO86]. undeniable [LW10]. underlying [FBZS12]. underwater [GFH^{+}10]. unification [Rom94]. unified [Sha09, YW98a]. Uniform [MMD99, FSA97]. uniformly [ERY^{+}16, RES^{+}16a, RES^{+}16b]. units [EG97]. universal [KHE01]. Unix [Ano87]. Unplanned [RS93]. unpredictability [Met16]. update [GKS06, LPTH99, ZJ02]. updating [GE97, MYYW06]. URLs [QWK14]. usability [ÇG11, MJRIV14]. USB [LYL12a]. Use [PDK95, Kri08, MJRIV14, SA03, TKCR04, WWDL09]. User [Kat05, LC00a, TCD07, BBMC98, FMS^{+}11, LWL93, LWD04, OO12, Red91, ZLL^{+}15]. User-centric [Kat05]. user-defined [FMS^{+}11]. user-network [BBMC98]. users [AST99]. Using [CL92, CKT94, GD06, NXG^{+}16, SSH99, WDF87, AKT12, AP16, BGFL08, BZLX16, CJV01, CW91, CD94, CGR^{+}09, CC95a, CH03, CH04, CZT^{+}16, CMRR02, DKP95, DD11, FCGC13, FH02a, FH02b, GCG^{+}13, GH11, GDK88, HJXY12, HPTC11, Hel13, HN94, HCK90, HMF93, Ito94, JWH^{+}15, JIB03, Kam93, KS97, KTV14, KHMK98, KW01, KSL12, KFW04, KiMKT94, LJ96, Lee09, LKK14, LYL^{+}12b, LC96a, MK11, MHPS96, MN14, MFU02, MN98, NHK14, NCB06, OESHK07, OESHK08, PONA11, PM94, PS95, QKZ^{+}07, RAF15, SKR08, SIGC15, VK08, Wan03, Wei95, WBY^{+}13, WYT01, YHLC13, YLK11, ZL16]. uSMIL [KFW06]. utilising [SL97]. Utility [LYL^{+}12b, JWH^{+}15]. Utility-based [LYL^{+}12b]. utilization [LP93].

validate [NCMH00]. Validation

[CO91, HF94, TC05]. valued [MLLS94].

Variable [LDZP08, FCGC13, Ric90, TR06]. variable-bandwidth [FCGC13]. Variable-length [LDZP08]. variance [LKK14]. variances [RAF15]. VBR [KSH96]. VC [ACA99]. vector [HFS^{+}08, Rav92, Rav95]. vehicle [IST^{+}13, SCL97]. vehicular [JOR^{+}00, WC08b]. ventilator [GA14]. Verification [HF94, BWB^{+}01, CDFP90, HN94, MLLS94, NC03, NC04, OESHK07, OESHK08, VT95, Wu01]. verify [SW06]. Verifying [LJ96]. version [Li98, Rom02]. versions [Gut01]. very [CZY^{+}03, Kra97]. VHDL [ZCL03]. via [CU07, Car14, CF86, HF94]. vibration [GFZ^{+}16]. video [CZY^{+}03, KSH96, LKK14, MG02, SL09, TR06, XC15, YDR^{+}09, ZEZF11]. video/voice [MG02]. Vietnamese [NXG^{+}16]. view [CH11, RKZ99, YMST98]. views [CGR^{+}09]. violations [XC15]. Virtual [SV13, FBZS12, GAHL00, JWH^{+}15, WLG^{+}11, CKJ10]. virtualization [XZ11]. visual [OO12, PMO16, KHL00]. visualization [SC09]. visualizing [KHL00]. VLSI [Pos98, RMB91, RC98, SEuH96, YS99]. vocabulary [FXZ16]. Voice [HJS87, MG02, TCD07]. Voice/Data [HJS87]. Volume [Ano97, Ano98, Ano01b, Ano03, Ano06, Ano07]. VOQL [KHL00]. voting [CY94, JC02, TK00]. vs [JC97]. VSM [HHCASEQ11]. VUI [TCD07].

Waiting [BG87]. WAN [WLG^{+}11]. Ward [PHPd98]. warping [KS08]. waste [RHS^{+}14]. watchdog [MHPS96]. water [BqQHj^{+}15, LL16]. water-related [LL16]. waveform [Car14]. Waveguide [Mae10]. wavelength [Din04, jKsJdChK12]. way [KR95, Spi99, ZZZZ15]. WDM [CCDL09, Din04, jKsJdChK12]. weather [DJ96]. Web [DGM06, BZLX16, CSB16, QL15, WCL15, ZLL^{+}15, BBO08, BGRV99, CZ09, CMZ10, FAL^{+}01, HPTC11, HFS^{+}08, KFW06, KVR03, KVR04, LV08, MMD99, QL13, SSP03, SC09, SW06, Spi99, TJS15, TS99, TBEH04, WOE06, WY02, YLQW13]. Web-based [FAL^{+}01, TJS15]. WebSources [BGRV99]. weighted [PTX^{+}09, TK00, Yan00, Yan02]. weighting [DR09]. Wide [MMD99, TS99]. Wiener [RES^{+}16a]. WiFi [Car15]. WiMax [Car15]. window [BYX16, YWW96]. Wireless [LL10, BXST12, CCZ03, Cho11, GS10, KTV14, KKHN12, KKHN13, LJL^{+}12, LYL^{+}12b, LSBW14, LH02, QYCG10, QKSN09, SP16, WH11, WWS16, XLSM10, YOM^{+}12]. Wisdom [MW90]. wise [GS10, LASS00]. within [MCD^{+}15, RCNL15]. Without [LBLB13, HDPC13]. WKBZ [LCWC12]. word [NXG^{+}16]. workflow [DGL00, GAHL00, GCJP03, QAY09, RMR00, vdAJ00a, vdA01]. workflows [DE00, DSS00, Joe00, KVG08]. workload [SY04]. workstation [YQ95]. workstation-based [YQ95]. workstations [Men99]. world [Jon03, MMD99, TS99]. worlds [TC05]. wormhole [ACR00]. worms [XFZ09]. wrapper [BGRV99]. wrappers [BGRV99]. write [WCD99]. write-invalidate [WCD99]. WSI [Ito94, KP90]. WSNs [HL16]. Wu [HLLC96]. WWW [LPTH99].

X [YWW96]. X-Window [YWW96].

XDRC [PXY93]. XML

[AP16, AD04, BKBM10, BKBM13, Cra06, FCD03, FJ04, FCD04, Gab06, GKS06, Jon03, LB03, LWHS06, MdlFD03, MM15, SY04, SA03, YWL^{+}03, Zhe03]. XML-based [AD04]. XPath [HF08, SF11]. XQuery [ZT03].

yielding [NK03]. yielding-first [NK03]. yields [Ito94].

ZigBee [RSDD10]. ZigBee/Bluetooth [RSDD10].

References

Avresky:1999:EGS

[AA99] D. R. Avresky and K. M.

Altawil. Embedding general spanning trees in faulty hypercubes. International Journal of Computer Systems Science and Engineering, 14(3): 167–174, May 1999. CODEN CSSEEI. ISSN 0267-6192.

Ashraf:2009:SBE

[AAKD09] Manzur Ashraf, Syed Mahfuzul Aziz, M. Lutful Kabir, and Biswajit K. Dey. A SIMbased electronic transaction authentication system. International Journal of Computer Systems Science and Engineering, 24(4):??, July 2009. CODEN CSSEEI. ISSN 0267-6192.

Abid:2000:PED

[Abi00] M. Abid. Prototyping environment for the design of hardware/software systems. International Journal of Computer Systems Science and Engineering, 15(3): 137–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Ahmed:2012:DIM

[ABNY12] Ejaz Ahmed, Nik Bessis, Peter Norrington, and Yong Yue. Data integration: the method of data mapping, matching and loading grid staging catalogue. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Antonellis:2009:STC

[ABP09] Ioannis Antonellis, Christos Bouras, and Vassilis Poulopoulos. Scalable text classification as a tool for personalization. International Journal of Computer Systems Science and Engineering, 24 (6):??, November 2009. CO-DEN CSSEEI. ISSN 0267-6192.

Audsley:1993:IUA

[ABRW93] N. C. Audsley, A. Burns, M. F. Richardson, and A. J. Wellings. Incorporating unbounded algorithms into predictable real-time systems. Computer Systems Science and Engineering, 8(2):80–89, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Arulambalam:1999:IPV

[ACA99] A. Arulambalam, X. Chen, and N. Ansari. Impact of per-VC and single FIFO queuing on ABR congestion control. International Journal of Computer Systems Science and Engineering, 14(5):275– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Aguilar:2009:MSP

[ACB09] José Aguilar, Jorge Chacal, and César Bravo. A multiagents system for planning and management of the production factors. International Journal of Computer Systems Science and Engineering, 24 (2):??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Avresky:2000:FTR

[ACR00] D. R. Avresky, C. Cunningham, and H. Ravichandran. Fault-tolerant routing for wormhole routed twodimensional meshes. International Journal of Computer Systems Science and Engineering, 15(6):??, December 2000. CODEN CSSEEI. ISSN 0267-6192.

Ardagna:2004:CMS

[AD04] C. Ardagna and S. De capitani di Vimercati. A comparison of modeling strategies in defining XML-based access control languages. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Anitha:2013:ADI

[AD13] R. Anitha and K. Duraiswamy. Application of digital image processing in the determination of mass of the diamonds. International Journal of Computer Systems Science and Engineering, 28(5): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Anitha:2014:ADI

[AD14] R. Anitha and K. Duraiswamy. Application of digital image processing in the determination of mass of the diamonds. International Journal of Computer Systems Science and Engineering, 29(5): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Al-Dabbagh:2015:PFT

[ADMB15] Rawaa Dawoud Al-Dabbagh, Saad Mekhilef, and Mohd Sapiyan Baba. Parameters’ fine tuning of differential evolution algorithm. International Journal of Computer Systems Science and Engineering, 30(2): ??, March 2015. CODEN CSSEEI. ISSN 0267-6192.

Aversano:2006:GPA

[AdPT06] L. Aversano, M. di Penta, and K. Taneja. A genetiv programming approach to support the design of service compositions. International Journal of Computer Systems Science and Engineering, 21 (4):??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Al-Dossari:2012:IQA

[ADSP12] Hmood Al-Dossari, Jianhua Shao, and Alun Preece. Improving QoS assessment over multiple attributes with asynchronous data. International Journal of Computer Systems Science and Engineering, 27 (2):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Alishvandi:2016:EDD

[AGP16] Hamid Alishvandi, Gholam Hosein Gouraki, and Hamid Parvin. An enhanced dynamic detection of possible invariants based on best permutation of test cases. International Journal of Computer Systems Science and Engineering, 31(1):??, January 2016. CODEN CSSEEI. ISSN 0267-6192.

Avresky:1998:SBF

[AGT98] D. R. Avresky, S. J. Geoghegan, and P. K. Tapadiya. A software-based fault-injection tool (SOFIT). International Journal of Computer Systems Science and Engineering, 13 (6):327–337, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Aguilar:2003:HAB

[Agu03] J. Aguilar. Heuristic algorithm based on a genetic algorithm for mapping parallel programs on hypercube multiprocessors. International Journal of Computer Systems Science and Engineering, 18 (4):??, July 2003. CODEN CSSEEI. ISSN 0267-6192.

Alam:2007:FMR

[AHBU07] Muhammad Alam, Michael Hafner, Ruth Breu, and Stefan Untherthiner. A framework for modelling restricted delegation of rights in the SECTET. International Journal of Computer Systems Science and Engineering, 22(5): ??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Angius:2016:EFE

[AHH^{+}16] Alessio Angius, András Horváth, Sami M Halawani, Omar Barukab, Ab Rahman Ahmad, and Gianfranco Balbo. Exploiting flow equivalent server in transient analysis. International Journal of Computer Systems Science and Engineering, 31(4):??, July 2016. CODEN CSSEEI. ISSN 0267-6192.

Arvalo:2015:SAL

[AJTT15] Sergio Arévalo, Ernesto Jiménez, Jian Tang, and Rommel Torres. Set agreement and the loneliness failure detector in crash-recovery systems. International Journal of Computer Systems Science and Engineering, 30(3):??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Ammar:1992:CAP

[AK92] Reda A. Ammar and Matthew J. Krzych. Computer aided performance engineering. A survey. Computer Systems Science and Engineering, 7(3): 170–189, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Alipanah:2012:OOD

[AKT12] Neda Alipanah, Latifur Khan, and Bjavani Thurisingham. Optmized ontology-driven query expansion using mapreduce framework to facilitate federated queries. International Journal of Computer Systems Science and Engineering, 27(2):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Aguilar:2003:OSL

[AL03] J. L. Aguilar and E. L. Leiss.

Optimal scheduling of loops with dependence cycles on parallel platforms. International Journal of Computer Systems Science and Engineering, 18(4):??, July 2003. CODEN CSSEEI. ISSN 0267-6192.

Ade:1999:BMR

[ALP99] Marleen Adé, Rudy Lauwereins, and J. A. Peperstraete. Buffer memory requirements in DSP applications. International Journal of Computer Systems Science and Engineering, 14(3):155–165, May 1999. CODEN CSSEEI. ISSN 0267-6192.

Ashihara:2000:AFP

[AM00] H. Ashihara and M. Matsumoto. Application of finite projective space to replicated data management. International Journal of Computer Systems Science and Engineering, 15(2):87–??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Ammar:1990:AAC

[Amm90] Reda A. Ammar. Algorithmic approach to construct probabilistic grammatical models of command streams in interactive software systems. Computer Systems Science and Engineering, 5(4):195– 201, October 1990. CODEN CSSEEI. ISSN 0267-6192.

Ali:2014:RPM

[AMuRKK14] Shahzad Ali, Sajjad A. Madani, Atta ur Rehman Khan, and Imran Ali Khan. Routing protocols for mobile sensor networks: a comparative study. International Journal of Computer Systems Science and Engineering, 29(2): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Andijani:1998:BMS

[And98] A. A. Andijani. Buffer management of a switch for ATM networks: a multi-objective approach. International Journal of Computer Systems Science and Engineering, 13 (6):379–386, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Anghelescu:2016:FIP

[Ang16] Petre Anghelescu. FPGA implementation of programmable cellular automata encryption algorithm for network communications. International Journal of Computer Systems Science and Engineering, 31 (5):??, September 2016. CO-DEN CSSEEI. ISSN 0267-6192.

Anon:1987:DDH

[Ano87] Anon. Design of I/O drivers for a high-availability Unix. Computer Systems Science and Engineering, 2(4):193– 196, October 1987. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:1990:E

[Ano90] Anonymous. Editorial. Computer Systems Science and Engineering, 5(1):3–??, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:1992:E

[Ano92]  Anonymous.  Editorial.

Computer Systems Science and Engineering, 7(2):67–??, April 1992. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:1997:IV

[Ano97] Anonymous. Index to volume 12, 1997. International Journal of Computer Systems Science and Engineering, 12(6): ??, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:1998:IV

[Ano98] Anonymous. Index to volume 13, 1998. International Journal of Computer Systems Science and Engineering, 13(6): ??, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:2001:BR

[Ano01a]  Anonymous.  Book re-

view. International Journal of Computer Systems Science and Engineering, 16(6):

??, November 2001. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:2001:IV

[Ano01b] Anonymous. Index to volume 16, 2001. International Journal of Computer Systems Science and Engineering, 16(6): ??, November 2001. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:2002:BR

[Ano02]  Anonymous.  Book re-

view. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Anonymous:2003:IV

[Ano03] Anonymous. Index to volume 18, 2003. International Journal of Computer Systems Science and Engineering, 18(6): ??, December 2003. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:2006:IV

[Ano06] Anonymous. Index to volume 21, 2006. International Journal of Computer Systems Science and Engineering, 21(6): ??, November 2006. CODEN CSSEEI. ISSN 0267-6192.

Anonymous:2007:IV

[Ano07] Anonymous. Index to volume 22, 2007. International Journal of Computer Systems Science and Engineering, 22(6): ??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Alqarni:2016:XSM

[AP16] Ahmad Abdullah Alqarni and Eric Pardede. XML schema matching using early dissimilarity detection approaches. International Journal of Computer Systems Science and Engineering, 31(5): ??, September 2016. CODEN CSSEEI. ISSN 0267-6192.

Apostolopoulos:1992:PRO

[Apo92] Theodore K. Apostolopoulos. Performance of readonly transactions in a distributed system. Computer Systems Science and Engineering, 7(4):236–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Apostolopoulos:2000:TDN

[Apo00] T. Apostolopoulos. A time dependent network fault management model in a TCP/ IP environment. International Journal of Computer Systems Science and Engineering, 15 (3):165–??, May 2000. CO-DEN CSSEEI. ISSN 0267-6192.

Alpert:1992:POO

[AR92] Sherman R. Alpert and Mary Beth Rosson. ParCE: an object-oriented approach to context-free parsing. Computer Systems Science and Engineering, 7(2):136–144, April 1992. CODEN CSSEEI. ISSN 0267-6192.

Anand:2004:GMA

[AR04] Sheila Anand and V. Ramachandran. A generic model for an application based intrusion prevention detection system. International Journal of Computer Systems Science and Engineering, 19(4): ??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Al-Sadi:2003:AFT

[ASDOK03] J. Al-Sadi, K. Day, and M. Ould-Khaoua. Analysis of fault-tolerant routing algorithms in k-ary n-cube networks. International Journal of Computer Systems Science and Engineering, 18(2): ??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Ali:2012:FSA

[ASH12] Hesham A. Ali, Mofreh M. Salem, and Ahmed A. Hamza. A framework for scalable autonomous P2P resource discovery for the grid implementation. International Journal of Computer Systems Science and Engineering, 27(4): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Achour:1999:BGB

[AST99] C. Ben Achour, C. Souveyet, and M. Tawbi. Bridging the gap between users and requirements engineering: the scenario-based approach. International Journal of Computer Systems Science and Engineering, 14(6):379– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Bradel:2012:ITJ

[BA12] Bory J. Bradel and Tarek S.

Abdelrahman. Inlining with traces in Java programs. International Journal of Computer Systems Science and Engineering, 27(4):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Bataineh:2000:RTA

[Bat00] S. Bataineh. Response time analysis of repairable multiprocessor systems. International Journal of Computer Systems Science and Engineering, 15(4):201–??, July 2000. CODEN CSSEEI. ISSN 0267-6192.

Barra:1987:ADA

[BBKT87] J. R. Barra, M. Becker, E. F.

Kouka, and M. Tricot. Application of data analysis methods and of simulated annealing for the automatic layout of circuits. Computer Systems Science and Engineering, 2(1):3–15, January 1987. CODEN CSSEEI. ISSN 0267-6192.

Baiocchi:1998:DCT

[BBMC98] Andrea Baiocchi, Nicola Blefari-Melazzi, and Francesca Cuomo. Dimensioning criteria of traffic contract parameters and connection admission control rules in an ATM user-network interface. International Journal of Computer Systems Science and Engineering, 13(4):207–217, July 1998. CODEN CSSEEI. ISSN 0267-6192.

Barhamgi:2008:PPP

[BBO08] M. Barhamgi, D. Benslimane, and A. M. Ouksel. PWSMS: A peer-to-peer Web service management system for data sharing in collaborative environments. International Journal of Computer Systems Science and Engineering, 23(2): ??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Baldoni:2001:DDB

[BCHR01] R. Baldoni, G. Cioffi, J.-M. Helary, and M. Raynal. Direct dependency-based determination of consistent global checkpoints. International Journal of Computer Systems Science and Engineering, 16 (1):??, January 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Bondavalli:2001:DAE

[BCL^{+}01] A. Bondavalli, M. Dal Cin, D. Latella, I. Majzik, A. Pataricza, and G. Savoia. Dependability analysis in the early phases of UML-based system design. International Journal of Computer Systems Science and Engineering, 16(5): ??, September 2001. CODEN CSSEEI. ISSN 0267-6192.

Bassiouni:1998:RFD

[BCLG98] Mostafa Bassiouni, Ming-Hsing Chiu, Margaret Loper, and Michael Garnsey. Relevance filtering for distributed interactive simulation. International Journal of Computer Systems Science and Engineering, 13(1):39–47, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Barbosa:2010:COS

[BCPS10] Luis S. Barbosa, Antonio Cerone, Alexander K. Petrenko, and Siraj A. Shaikh. Certification of open-source software: A role for formal methods? International Journal of Computer Systems Science and Engineering, 25 (4):??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Bagherzadeh:1996:PRB

[BD96] Nader Bagherzadeh and Martin Dowd. Problems on routing bounded distance assignments in hypercubes. International Journal of Computer Systems Science and Engineering, 11(4):221–226, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Bhagavathi:1994:SMD

[BDG^{+}94] D. Bhagavathi, W. M. Denny, C. E. Grosch, P. J. Looges, and S. Olariu. Sorting and merging on the DAP. International Journal of Computer Systems Science and Engineering, 9(3):175–183, July 1994. CODEN CSSEEI. ISSN 0267-6192.

Byrski:2012:ABC

[BDKD12] Aleksander Byrski, Roman Debski, and Marek Kisiel-Dorohinicki. Agent-based computing in an augmented cloud environment. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Bieker:1997:FTR

[BDMV97] Bernd Bieker, Geert Deconinck, Erik Maehle, and Johan Vounckx. Fault-tolerant routing, reconfiguration and backward error recovery for parallel systems. International Journal of Computer Systems Science and Engineering, 12 (4):245–253, July 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Becker:1990:SPN

[BDPV90] M. Becker, C. Dekoninck, J. P. Prost, and B. Verrier. Stochastic Petri net model for the FPS/264. Computer Systems Science and Engineering, 5(2):65–72, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Bondavalli:1993:CEF

[BDX93] Andrea Bondavalli, Felicita Di Giandomenico, and Jie Xu. Cost-effective and flexible scheme for software fault tolerance. Computer Systems Science and Engineering, 8 (4):234–244, October 1993.

CODEN CSSEEI. ISSN 0267-6192.

Bedi:1988:FSA

[Bed88] S. Bedi. File system activity and the distribution of disc transfers on a multiaccess system. Computer Systems Science and Engineering, 3(4): 199–200, October 1988. CO-DEN CSSEEI. ISSN 0267-6192.

Bassiouni:1987:CEW

[BG87] M. A. Bassiouni and A. Gupta.

Computing estimates of waiting times in ring local area networks with priority classes. Computer Systems Science and Engineering, 2(4):155– 160, October 1987. CODEN CSSEEI. ISSN 0267-6192.

Banci:2008:MDD

[BGFL08] Michele Banci, Stefania Gnesi, Alessandro Fantechi, and Giovanni Lombardi. Model driven development of railway systems using diversity. International Journal of Computer Systems Science and Engineering, 23(5):??, September 2008. CODEN CSSEEI. ISSN 0267-6192.

Beraldi:1998:ETA

[BGM98] Roberto Beraldi, Paolo Giacomazzi, and Riccardo Melen. Efficient topologies for ATM switching fabrics with shortest path and deflection routing. International Journal of Computer Systems Science and Engineering, 13(5):271– 281, September 1998. CO-DEN CSSEEI. ISSN 0267-6192.

Becker:2001:IEC

[BGNP01] L. B. Becker, M. Gergeleit, E. Nett, and C. E. Pereira. An integrated environment for the complete development cycle of an object-oriented distributed real-time system. International Journal of Computer Systems Science and Engineering, 16(2):??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Bright:1999:WGT

[BGRV99] Laura Bright, Jean-Robert Gruser, Louiqa Raschid, and Maria Esther Vidal. A wrapper generation toolkit to specify and construct wrappers for Web accessible data sources (WebSources). International Journal of Computer Systems Science and Engineering, 14 (2):83–97, March 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Balsamo:1986:SQN

[BI86] S. Balsamo and G. Iazeolla. Synthesis of queueing networks with block and statedependent routing. Computer Systems Science and Engineering, 1(2):99–108, January 1986. CODEN CSSEEI. ISSN 0267-6192.

Bekiropoulos:2010:LFG

[BKBM10] K. Bekiropoulos, E. Keramopoulos, O. Beza, and P. Mouratidis. A list of features that a graphical XML query language should support. International Journal of Computer Systems Science and Engineering, 25(5): ??, September 2010. CODEN CSSEEI. ISSN 0267-6192.

Bekiropoulos:2013:LFG

[BKBM13] K. Bekiropoulos, E. Keramopoulos, O. Beza, and P. Mouratidis. A list of features that a graphical XML query language should support. International Journal of Computer Systems Science and Engineering, 28(2): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Bamba:2010:MAS

[BLY10] Bhuvan Bamba, Ling Liu, and Philip S. Yu. A motion-aware safe period-based framework for spatial alarm processing. International Journal of Computer Systems Science and Engineering, 25(3): ??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Bozyigit:1997:LBF

[BM97] M. Bozyigit and M. Melhi.

Load balancing framework for distributed systems. International Journal of Computer Systems Science and Engineering, 12(5):287–293, September 1997. CODEN CSSEEI. ISSN 0267-6192.

Bouganim:2005:TRU

[BNP05] Luc Bouganim, François Dang Ngoc, and Philippe Pucharel. Tamper-resistant ubiquitous data management. International Journal of Computer Systems Science and Engineering, 20(2):??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Bondavalli:2001:HMC

[BNSM01] A. Bondavalli, M. Nelli, L. Simoncini, and G. Mongardi. Hierarchical modelling of complex control systems: dependability analysis of a railway interlocking. International Journal of Computer Systems Science and Engineering, 16(4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Bowen:1986:DSC

[Bow86] Jonathan Bowen. Design of a simple Cambridge ring interface adapter. Computer Systems Science and Engineering, 1(2):93–98, January 1986. CODEN CSSEEI. ISSN 0267-6192.

Baldoni:1998:ECB

[BPRS98] Roberto Baldoni, Ravi Prakash, Michel Raynal, and Mukesh Singhal. Efficient Δ-causal broadcasting. International Journal of Computer Systems Science and Engineering, 13 (5):263–269, September 1998.

CODEN CSSEEI. ISSN 0267-6192.

Blesa:2006:GPI

[BPX06] Maria J. Blesa, Jordi Petit, and Fatos Xhafa. Generic parallel implementations for tabu search. International Journal of Computer Systems Science and Engineering, 21(6): ??, November 2006. CODEN CSSEEI. ISSN 0267-6192.

Bao-quan:2015:DAC

[BqQHj^{+}15] Jin Bao-quan, Bai Qing, Zhang Hong-juan, Wang dong, and Gao Yan. Design and analysis of communication scheduling applied in water inrush perception layer of mine Internet of Things. International Journal of Computer Systems Science and Engineering, 30(5): ??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Boswell:1985:DPL

[BSK85] C. Boswell, K. K. Saluja, and K. Kinoshita. Design of programmable logic arrays for parallel testing. Computer Systems Science and Engineering, 1(1):5–16, October 1985. CODEN CSSEEI. ISSN 0267-6192.

Bhattacharjee:2014:TPQ

[BSMB14] Jaya Bhattacharjee, Anirban Sengupta, Chandan Mazumdar, and Mridul Sankar Barik. A two-phase quantitative methodology for enterprise information security risk analysis. International Journal of Computer Systems Science and Engineering, 29(1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Bai:2011:AUL

[BSR11] Xin Bai, Michael B. Smith, and Rathika Rajaravivarma. Affordance of ubiquitous learning through cloud computing. International Journal of Computer Systems Science and Engineering, 26(5): ??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Beresford-Smith:1990:ERA

[BSS90] B. Beresford-Smith and H. Schroder. Effective reconfiguration algorithms in fault-tolerant processor arrays. Computer Systems Science and Engineering, 5(3):169–177, July 1990. CODEN CSSEEI. ISSN 0267-6192.

Black:1998:QLA

[BST98] C. Black, S. M. Shatz, and S. Tu. A query language for automated general analysis of Concurrent Ada programs. International Journal of Computer Systems Science and Engineering, 13(2): 83–95, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Burneau:1992:CEO

[BT92] Jean-Christophe Burneau and Odile Thiery. Classifying evolving objects in objectoriented knowledge bases. Computer Systems Science and Engineering, 7(2):86–90, April 1992. CODEN CSSEEI. ISSN 0267-6192.

Bobbio:1998:NES

[BT98] Andrea Bobbio and Miklos Telek. Non-exponential stochastic Petri nets: an overview of methods and techniques. International Journal of Computer Systems Science and Engineering, 13(6):339– 351, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Buchholz:2000:ECE

[Buc00] P. Buchholz. Efficient computation of equivalent and reduced representations for stochastic automata. International Journal of Computer Systems Science and Engineering, 15(2):93–??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Bulus:2011:DAS

[Bul11] Ercan Bulus. Designing attacks for SMTP servers. International Journal of Computer Systems Science and Engineering, 26(1):??, January 2011. CODEN CSSEEI. ISSN 0267-6192.

Bundell:1998:MDS

[Bun98] Gary A. Bundell. Modelling of distributed systems with timed event graphs. International Journal of Computer Systems Science and Engineering, 13(1):49–54, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Biswas:2013:SCT

[BV13] Debmalya Biswas and Krishnamurthy Vidyasankar. Secure cloud transactions. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Budrikis:1996:PEC

[BVCW96] Zigmantas Budrikis, Steven Van Luinen, Antonio Cantoni, and Vaughan Wittorff. Performance evaluation of controlled cell transfer service in ATM LAN. International Journal of Computer Systems Science and Engineering, 11 (6):353–360, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Burns:2001:MVA

[BWB^{+}01] A. Burns, A. J. Wellings, F. Burns, A. M. Koelmans, M. Koutny, A. Romanovsky, and A. Yakovlev. Modelling and verification of an atomic action protocol implemented in Ada. International Journal of Computer Systems Science and Engineering, 16(3): ??, May 2001. CODEN CSSEEI. ISSN 0267-6192.

Bleul:2006:OQA

[BWG06] S. Bleul, T. Weise, and K. Geihs. An ontology for quality-aware service discovery. International Journal of Computer Systems Science and Engineering, 21(4):

??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Barolli:2012:SEM

[BXST12] Admir Barolli, Fatos Xhafa, Christian Sánchez, and Makoto Takizawa. A study on the effect of mutation in genetic algorithms for mesh router placement in wireless mesh networks. International Journal of Computer Systems Science and Engineering, 27(1): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Bo:2016:FPP

[BYX16] Chen Bo, Duan Cheng YOng, and Gao Xiue. A frequent pattern parallel mining algorithm based on distributed sliding window. International Journal of Computer Systems Science and Engineering, 31 (2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Bao:2016:SKM

[BZLX16] Wenshu Bao, Haiying Zhou, Wang Lu, and Feng Xie. The system of knowledge management using web based learning and education technology. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Carmichael:2014:DDS

[Car14] James Carmichael. Diagnosis of dysarthria subtype via spectral and waveform analysis. International Journal of Computer Systems Science and Engineering, 29(1): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Carvalho:2015:OWB

[Car15] Glaucio H. S. Carvalho. Optimal WiMax backhauling solutions for WiFi traffic. International Journal of Computer Systems Science and Engineering, 30(2):??, March 2015. CODEN CSSEEI. ISSN 0267-6192.

Chen:1995:SIB

[CC95a] Jonathan Jen-Rong Chen and Henry Ker-Chang Chang. Secure information broadcasting scheme using embedded locks. International Journal of Computer Systems Science and Engineering, 10(2):67–74, April 1995. CODEN CSSEEI. ISSN 0267-6192.

Chung:1995:FTE

[CC95b] K.-L. Chung and Yu-W. Chen. Fault-tolerant embedding multiple complete binary trees into hypercubes. International Journal of Computer Systems Science and Engineering, 10(3):187–191, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Chiang:1997:BSN

[CC97] Wei-Kuo Chiang and Rong-Jaye Chen. Block-switch networks: a cost-effective class of interconnection networks. International Journal of Computer Systems Science and Engineering, 12(3): 175–185, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Chen:1998:FTE

[CC98a] Yu-Wei Chen and Kuo-Liang Chung. On fault-tolerant embedding complete binary trees into hypercubes. International Journal of Computer Systems Science and Engineering, 13(2):97–100, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Chung:1998:FPP

[CC98b] Kuo-Liang Chung and Yu-Wei Chen. Fast pipelined prefix computations on MIMD hypercubes. International Journal of Computer Systems Science and Engineering, 13 (2):125–130, March 1998. CO-DEN CSSEEI. ISSN 0267-6192.

Chakraborty:2014:GOS

[CC14] Supriya Chakraborty and Nabendu Chaki. Generic organization of semi-structured data. International Journal of Computer Systems Science and Engineering, 29(1): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Chahooki:2015:BSG

[CC15] Mohammad Ali Zare Chahooki and Nasrollah Moghaddam Charkari. Bridging the semantic gap for automatic image annotation by learning the manifold space. International Journal of Computer Systems Science and Engineering, 30(4):??, July 2015. CODEN CSSEEI. ISSN 0267-6192.

Chahooki:2016:BSG

[CC16] Mohammad Ali Zare Chahooki and Nasrollah Moghaddam Charkari. Bridging the semantic gap for automatic image annotation by learning the manifold space. International Journal of Computer Systems Science and Engineering, 31(1):??, January 2016. CODEN CSSEEI. ISSN 0267-6192.

Chung:2005:CPQ

[CCC05] Chin-Wan Chung, Sunghee Choi, and Yong-Jin Choi. Closest pair queries in spatiotemporal databases. International Journal of Computer Systems Science and Engineering, 20(2):??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Cilardo:2007:DIH

[CCdF^{+}07] A. Cilardo, D. Cotroneo, C. di Flora, A. Mazzeo, L. Romano, and S. Russo. Design and implementation of a high performance architecture for providing digital time stamping services to mobile devices. International Journal of Computer Systems Science and Engineering, 22(3): ??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Chiu:2009:RDI

[CCDL09] Angela L. Chiu, Gagan Choudhury, Robert Doverspike, and Guangzhi Li. Restoration design in IP over reconfigurable WDM networks. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Chen:2003:FEA

[CCHL03] L.-C. Chen, D.-R. Chen, C.- C. Hsu, and H.-R. Lin. Fast and efficient allocation for mesh systems. International Journal of Computer Systems Science and Engineering, 18 (3):??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Chen:2011:FBD

[CCL11] Jungying Chen, Dan Chen, and Oliver Lemon. A featurebased detection and tracking system for gaze and smiling behaviours. International Journal of Computer Systems Science and Engineering, 26 (3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Chella:2006:APA

[CCSS06] Antonio Chella, Massimo Cossentino, Luca Sabatucci, and Valeria Seidita. Agile PASSI: An agile process for designing agents. International Journal of Computer Systems Science and Engineering, 21(2):??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Chan:2003:ARP

[CCZ03] H. C. B. Chan, H. Chen, and J. Zhang. An adaptive reservation protocol with a preemptive priority mechanism for wireless ATM LANs. International Journal of Computer Systems Science and Engineering, 18(2):??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Chang:1993:TPM

[CD93] Elizabeth Chang and T. S. Dillon. Two phase methodology for the design of the human computer interface. Computer Systems Science and Engineering, 8(3):131– 143, July 1993. CODEN CSSEEI. ISSN 0267-6192.

Chang:1994:ASR

[CD94] Elizabeth Chang and Tharam S. Dillon. Achieving software reliability and fault tolerance using the object oriented paradigm. Computer Systems Science and Engineering, 9 (2):118–121, April 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Chang:2003:E

[CDF03] E. Chang, E. Damiani, and L. Feng. Editorial. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Carchiolo:1990:CSP

[CDFP90] Vincenza Carchiolo, Antonella Di Stefano, Alberto Faro, and Giuseppe Pappalardo. Case study in protocol verification. Computer Systems Science and Engineering, 5(3):178–186, July 1990. CODEN CSSEEI. ISSN 0267-6192.

Ciciani:1992:DSL

[CDY92] Bruno Ciciani, Daniel M.

Dias, and Philip S. Yu. Dynamic and static load sharing in hybrid distributedcentralized database systems. Computer Systems Science and Engineering, 7(1):25– 41, January 1992. CODEN CSSEEI. ISSN 0267-6192.

Ceravolo:2004:ERH

[Cer04] P. Ceravolo. Extracting role hierarchies from authentication data flows. International Journal of Computer Systems Science and Engineering, 19 (3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Corsini:1986:MCS

[CF86] P. Corsini and G. Frosini.

Methodology for complex system description via extended Petri nets. Computer Systems Science and Engineering, 1 (4):205–212, October 1986. CODEN CSSEEI. ISSN 0267-6192.

Carminati:2006:SCO

[CF06] B. Carminati and E. Ferraru.

A system for controlled outsourcing of personal data. International Journal of Computer Systems Science and Engineering, 21(1):??, January 2006. CODEN CSSEEI. ISSN 0267-6192.

Chung:2006:SIE

[CFLZ06] J-Y Chung, G. Feuerlicht, W. Lamersdorf, and C. Zirpins. Special issue: Engineering design and composition of service-oriented applications. International Journal of Computer Systems Science and Engineering, 21(4): ??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Chu:2000:TLQ

[CG00] J.-L. Chu and R. K. Guha. Triangular level quorums for distributed mutual exclusion. International Journal of Computer Systems Science and Engineering, 15(6): ??, December 2000. CODEN CSSEEI. ISSN 0267-6192.

Cetyn:2011:AUR

[ÇG11] Görkem Çetyn and Mehmet Göktürk. Assessing usability readiness of collaborative projects. International Journal of Computer Systems Science and Engineering, 26(4): ??, July 2011. CODEN CSSEEI. ISSN 0267-6192.

Chatterjee:2009:FFE

[CGR^{+}09] A. Chatterjee, M. Goel, N. Raghuvanshi, P. Gupta, and S. G. Dhande. Facial feature extraction in frontal views using activeshape models. International Journal of Computer Systems Science and Engineering, 24(6): ??, November 2009. CODEN CSSEEI. ISSN 0267-6192.

Chao:1997:DAS

[CH97] H. Jonathan Chao and Jun S.

Hong. Design of an ATM shaping multiplexer with guaranteed output burstiness. International Journal of Computer Systems Science and Engineering, 12(2):131– 141, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Cheng:1998:ICA

[CH98] Lily Cheng and Herman Hughes. Integrating call admission and cell scheduling. International Journal of Computer Systems Science and Engineering, 13(4):219– 226, July 1998. CODEN CSSEEI. ISSN 0267-6192.

Chen:2003:FTR

[CH03] D-R Chen and C-C Hsu.

Fault-tolerant routing for pyramid networks using the least level minimal routing method. International Journal of Computer Systems Science and Engineering, 18(1): ??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Chen:2004:FTR

[CH04] D.-R. Chen and C.-C. Hsu.

Fault-tolerant routing for pyramid networks using the least level minimal routing method. International Journal of Computer Systems Science and Engineering, 19(1): ??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Colman:2006:ASO

[CH06] A. Colman and J. Han. Adaptive service-oriented systems: an organisational approach. International Journal of Computer Systems Science and Engineering, 21(4):??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Cho:2011:SVD

[CH11] Eun-Sun Cho and Manpyo Hong. A structural view definition for RDF data. International Journal of Computer Systems Science and Engineering, 26(1):??, January 2011. CODEN CSSEEI. ISSN 0267-6192.

Chaudhuri:2002:ODA

[Cha02] P. Chaudhuri. An optimal distributed algorithm for finding a set of fundamental cycles in a graph. International Journal of Computer Systems Science and Engineering, 17(1): ??, January 2002. CODEN CSSEEI. ISSN 0267-6192.

Chang:2008:AST

[Cha08] Joong Hyuk Chang. Adaptive selection of tuples over data streams for efficient load shedding. International Journal of Computer Systems Science and Engineering, 23(4):

??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Chalmers:1990:IPE

[CHCL90] A. G. Chalmers, J. W.

Hearne, K. J. Cameron, and N. E. Ligeti. Implementation and performance evaluation of an optimization algorithm on transputers. Computer Systems Science and Engineering, 5(1):42–46, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Chen:1993:SDS

[Che93] Ling Chen. Systematic design of systolic arrays from affine recurrence techniques. Computer Systems Science and Engineering, 8(3):144– 153, July 1993. CODEN CSSEEI. ISSN 0267-6192.

Cheung:2007:FMS

[Che07] K. S. Cheung. A formal method for synthesising components of shared resource systems. International Journal of Computer Systems Science and Engineering, 22(6): ??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Cokuslu:2012:RAQ

[CHE12] Deniz Cokuslu, Abdelkader Hemeurlain, and Kayhan Erciyes. Resource allocation for query processing in grid systems: a survey. International Journal of Computer Systems Science and Engineering, 27 (4):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Chen:1999:SDM

[CHKW99] Meng Chang Chen, Jan-Ming Ho, Ming-Tat Ko, and Shie-Yuan Wang. A SCSI disk model for multimedia storage systems. International Journal of Computer Systems Science and Engineering, 14(3): 147–154, May 1999. CODEN CSSEEI. ISSN 0267-6192.

Chen:2012:MLB

[CHL^{+}12] Hsing-Chung Chen, Yung-Fa Huang, Shu-Hong Lee, Chi-Tung Chen, and Ho-Lung Hung. A mobile locationbased with time-constraint RBAC associated database management model. International Journal of Computer Systems Science and Engineering, 27(6):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Cho:2011:DMC

[Cho11] Haengrae Cho. Distributed multidimensional clustering based on spatial correlation in wireless sensor networks. International Journal of Computer Systems Science and Engineering, 26(4):??, July 2011. CODEN CSSEEI. ISSN 0267-6192.

Chen:2011:CSE

[CHZ11] Xiangxian Chen, Hai Huang, and Jiafang Zhang. In-car speech enhancement based on ensemble empirical mode decomposition. International Journal of Computer Systems Science and Engineering, 26 (1):??, January 2011. CO-DEN CSSEEI. ISSN 0267-6192.

Clapp:1994:LST

[CI94] Kent C. Clapp and Ravishankar K. Iyer. Large system test data analysis. Computer Systems Science and Engineering, 9(2):122–133, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Ceska:2001:GUS

[CJV01] M. Ceska, V. Janousek, and T. Vojnar. Generating and using state spaces of objectoriented Petri nets. International Journal of Computer Systems Science and Engineering, 16(3):??, May 2001. CODEN CSSEEI. ISSN 0267-6192.

Cao:2010:MMM

[CKJ10] Yiwei Cao, Ralf Klamma, and Matthias Jarke. Mobile multimedia management for Virtual Campfire — the German excellence research cluster UMIC. International Journal of Computer Systems Science and Engineering, 25 (3):??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Christmansson:1994:DFC

[CKT94] Jörgen Christmansson, Zbigniew Kalbarczyk, and Jan Torin. Dependable flight control system using data diversity with error recovery. Computer Systems Science and Engineering, 9(2):142–150, April 1994. CODEN CSSEEI. ISSN 0267-6192. URL

                                        http://www.ce.chalmers.
                                        se/Documents/com_sys.ps.
                                        Z.

Choi:1992:UOO

[CL92] Amelia Choi and W. S. Luk. Using an object-oriented database system to construct a spatial database kernel for GIS applications. Computer Systems Science and Engineering, 7(2):100–121, April 1992. CODEN CSSEEI. ISSN 0267-6192.

Cordova:1996:MTP

[CL96] Javier Cordova and Yann-Hang Lee. Multicast trees to provide message ordering in mesh networks. International Journal of Computer Systems Science and Engineering, 11 (1):3–13, January 1996. CO-DEN CSSEEI. ISSN 0267-6192.

Chen:2000:IBD

[CL00] J. J.-R. Chen and Y. Liu. An ID-based digital multisignatures scheme with time stamp technique. International Journal of Computer Systems Science and Engineering, 15(2):105–??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Chen:2007:GAF

[CL07] Rung-Ching Chen and Cheun-Chieh Liao. A genetic algorithm with fuzzy selection and local search for multicast routing problem on the QoS constraint. International Journal of Computer Systems Science and Engineering, 22 (4):??, July 2007. CODEN CSSEEI. ISSN 0267-6192.

Chang:2016:ISP

[CL16] Shuchih Ernest Chang and Anne Yenching Liu. Information security in practices: Exploring privacy and trust in computer and Internet surveillance. International Journal of Computer Systems Science and Engineering, 31 (2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Chennapragada:1992:SRR

[CLO92] D. Chennapragada, P. Looges, and S. Olariu. Single row routing with indifference graphs on the Distributed Array Processor. Computer Systems Science and Engineering, 7(4):249–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Cupens:2004:AAM

[CM04] C. F. Cupens and A. Miége.

AdOrBAC: an administration model for Or-BAC. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Crookes:1991:OIA

[CMM91] D. Crookes, P. J. Morrow, and P. J. McParland. Occam implementation of an algebrabased language for low-level image processing. Computer Systems Science and Engineering, 6(1):30–??, January 1991. CODEN CSSEEI. ISSN 0267-6192.

Chen:2003:SDM

[CMMA03] X-X Chen, M. Minami, H. Morikawa, and T. Aoyama. Service description model and toolkit for ubiquitous computing. International Journal of Computer Systems Science and Engineering, 18(1): ??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Chen:2004:SDM

[CMMA04] X.-X. Chen, M. Minami, H. Morikawa, and T. Aoyama. Service description model and toolkit for ubiquitous computing. International Journal of Computer Systems Science and Engineering, 19(1): ??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Cotroneo:2002:FTA

[CMRR02] D. Cotroneo, N. Mazzocca, L. Romano, and S. Russo. A fault tolerant access to legacy database systems using CORBA technology. International Journal of Computer Systems Science and Engineering, 17(6):??, November 2002. CODEN CSSEEI. ISSN 0267-6192.

Chen:2010:MCW

[CMZ10] Xiaowu Chen, Yongtao Ma, and Qinping Zhao. Multicategory Web object extraction based on relation schema. International Journal of Computer Systems Science and Engineering, 25(6): ??, November 2010. CODEN CSSEEI. ISSN 0267-6192.

Conway:1991:VAT

[CO91] A. E. Conway and D. O’Brien.

Validation of an approximation technique for queueing network models with chaindependent FCFS queues. Computer Systems Science and Engineering, 6(2):117–??, April 1991. CODEN CSSEEI. ISSN 0267-6192.

Cooper:1990:PAD

[CP90] R. K. Cooper and D. A.

Peshkin. Parallel alternating direction implicit method on a network of transputers. Computer Systems Science and Engineering, 5(1): 47–52, January 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Coronato:2015:ASD

[CP15] Antonio Coronato and Giovanni Paragliola. An anomalous situation detection system for cognitive impaired people. International Journal of Computer Systems Science and Engineering, 30(1): ??, January 2015. CODEN CSSEEI. ISSN 0267-6192.

Costa:2011:ELD

[CPCB11] Danielle Costa, Luci Pirmez, Luiz F. R. C. Carmo, and Luiz F. H. Bacellar. Enhancing Levenshtein distance algorithm for assessing behavioral trust. International Journal of Computer Systems Science and Engineering, 26(2): ??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Corsini:1985:MED

[CPS85] P. Corsini, C. A. Prete, and L. Simoncini. MuTEAM: An experience in the design of robust multiprocessor systems. Computer Systems Science and Engineering, 1(1): 23–35, October 1985. CO-DEN CSSEEI. ISSN 0267-6192.

Comito:2008:PPP

[CPT08] C. Comito, S. Patarin, and D. Talia. PARIS: A peerto-peer architecture for largescale semantic data sharing. International Journal of Computer Systems Science and Engineering, 23(2): ??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Ciccarella:1990:TAD

[CPV90] G. Ciccarella, G. Paris, and G. Valent. Top-down approach to the design of OC-CAM and transputer-based real-time systems. Computer Systems Science and Engineering, 5(1):5–12, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Clement:1996:ASA

[CQ96] Mark J. Clement and Michael J.

Quinn. Architectural scaling and analytical performance prediction. International Journal of Computer Systems Science and Engineering, 11(3):159–167, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Chen:2016:DPF

[CQN^{+}16] Fengjuan Chen, Wenyu Qu, Lihai Nie, Junfeng Wu, and Yuanyuan Li. Discovering probabilistic frequent closed itemsets in uncertain database with tuple uncertainty. International Journal of Computer Systems Science and Engineering, 31(2): ??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Crampton:2006:AHR

[Cra06] Jason Crampton. Applying hierarchical and role-based access control to XML documents. International Journal of Computer Systems Science and Engineering, 21(5): ??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

Chen:1997:NIB

[CS97] Jonathan J-R Chen and Ping-Tai Sun. New ID-based cryptosystem based on number theory. International Journal of Computer Systems Science and Engineering, 12(1):

37–41, January 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Cao:2015:IAL

[CS15] Jijun Cao and Jinshu Su. Improving application layer multicast forwarding performance by offloading multisend operations. International Journal of Computer Systems Science and Engineering, 30(3): ??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Chawla:2016:GSB

[CSB16] Shailey Chawla, Sangeeta Srivastava, and Punam Bedi. Goals and scenarios based web requirements engineering. International Journal of Computer Systems Science and Engineering, 31(1): ??, January 2016. CODEN CSSEEI. ISSN 0267-6192.

Cao:2014:SHS

[CSQW14] Yangjie Cao, Hongyang Sun, Depei Qian, and Weiguo Wu. Scalable hierarchical scheduling for malleable parallel jobs on multiprocessor-based systems. International Journal of Computer Systems Science and Engineering, 29(2): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Cao:2011:DDD

[CSXC11] Jijun Cao, Jinshu Su, Jing Xie, and Feng Chen. DBMISD-D: A distributed algorithm for constructing and maintaining high stability applicationlayer multicast tree. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Colajanni:1993:ERB

[CT93] M. Colajanni and S. Tucci.

Equivalence relation between Fork-Join systems and Nstage queueing networks. Computer Systems Science and Engineering, 8(1):3–12, January 1993. CODEN CSSEEI. ISSN 0267-6192.

Camp:2007:IEG

[CU07] Garrett Camp and Mihaela Ulieru. InOrder: enhancing Google via stigmergic query refinement. International Journal of Computer Systems Science and Engineering, 22(4):??, July 2007. CODEN CSSEEI. ISSN 0267-6192.

Chang:1991:BCC

[CW91] Chin-Chen Chang and Tzong-Chen Wu. Broadcasting cryptosystem in computer networks using interpolating polynomials. Computer Systems Science and Engineering, 6(3):185–??, July 1991. CODEN CSSEEI. ISSN 0267-6192. See comments [HLLC96, LC96a].

Chen:2011:GIH

[CWBZ11] Dan Chen, Lizhe Wang, Congcong Bian, and Xuguang Zhang. A grid infrastructure for hybrid simulations. International Journal of Computer Systems Science and Engineering, 26(3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Chung:2000:CMS

[CWS^{+}00] C.-M. Chung, Y.-H. Wang, T. K. Shih, H.-C. Keh, and J.-F. Chen. The complexity measurement of software through program decomposition. International Journal of Computer Systems Science and Engineering, 15(2):127– ??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Chang:1998:BSC

[CWY98] Chin-Chen Chang, Tzong-Chen Wu, and Yi-Shiung Yeh. Broadcasting secrets in communication networks. International Journal of Computer Systems Science and Engineering, 13(2):121– 124, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Cornell:1990:IAB

[CY90] Douglas W. Cornell and Philip S. Yu. Integrated approach to buffer management and query optimization. Computer Systems Science and Engineering, 5(4): 243–251, October 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Chang:1994:EAE

[CY94] Her-Kun Chang and Shyan-Ming Yuan. Efficient algorithm for evaluating the availability of multidimensional voting. Computer Systems Science and Engineering, 9 (1):21–24, January 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Casey:2009:ACE

[CZ09] John Casey and Wanlei Zhou.

Adjacency cache: an efficient P2P Web cache indexing and lookup system. International Journal of Computer Systems Science and Engineering, 24 (5):??, September 2009. CO-DEN CSSEEI. ISSN 0267-6192.

Chen:2016:TTR

[CZT^{+}16] Yifan Chen, Xiang Zhao, Jiuyang Tang, Weiming Zhang, and Haichuan Shang. Taxitaking recommendation using real-time trajectories: an online query based approach. International Journal of Computer Systems Science and Engineering, 31(2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Chen:2003:FSQ

[CZY^{+}03] J. Chen, J. Zhou, S. Yu, J. Zheng, C. Zhang, and L. Zhong. A framework of scalar quantization multiwavelet transform for very low bit rate video coding. International Journal of Computer Systems Science and Engineering, 18(4):??, July 2003. CODEN CSSEEI. ISSN 0267-

deAraujo:1999:OOH

[dA99] T. Vianna de Araujo. Objectoriented hierarchies and their resulting complex systems. International Journal of Computer Systems Science and Engineering, 14(6):353–??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Datta:1995:SSM

[DAFG95] Ajoy Kumar Datta, Oscar R. Aguilar, Mitchell Flatebo, and Sukumar Ghosh. Simulating shared memory primitives in asynchronous message passing systems. International Journal of Computer Systems Science and Engineering, 10(2):75–83, April 1995. CODEN CSSEEI. ISSN 0267-6192.

Damiani:2004:SAT

[Dam04] E. Damiani. Semantics-aware techniques for security. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Dang:2007:SAS

[Dan07] Tran Khanh Dang. Solving approximate similarity queries. International Journal of Computer Systems Science and Engineering, 22(1– 2):??, January/March 2007. CODEN CSSEEI. ISSN 0267-6192.

Damiani:2010:SIO

[DBBA10] Ernesto Damiani, Luis Barbosa, Peter T. Breuer, and Claudio A. Ardagna. Special issue: Open source certification. International Journal of Computer Systems Science and Engineering, 25(4): ??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Das:1990:PHA

[DD90] Sajal K. Das and Narsingh Deo. Parallel Hungarian algorithm. Computer Systems Science and Engineering, 5 (3):131–136, July 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Dimakopoulos:1997:BH

[DD97] Vassilios V. Dimakopoulos and Nikitas J. Dimopoulos. Broadcasting in hypercycles. International Journal of Computer Systems Science and Engineering, 12(3): 187–192, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Dowlatshahi:1998:PAA

[DD98] Mehran Dowlatshahi and S. K. De. Performance analysis of ATM switches fed with periodic and correlated traffic. International Journal of Computer Systems Science and Engineering, 13(5):303– 310, September 1998. CO-DEN CSSEEI. ISSN 0267-

Dolia:2011:NAT

[DD11] Prashant M. Dolia and Ashwin R. Dobariya. New approach towards event monitoring through IP surveillance technology using 4G mobile devices. International Journal of Computer Systems Science and Engineering, 26(1): ??, January 2011. CODEN CSSEEI. ISSN 0267-6192.

Deng:1990:OTT

[DDL^{+}90] X. L. Deng, T. Dillon, K. Lew, J. Rankin, E. Smith, and D. Suter. Optimal topologies of transputers for different classes of problems. Computer Systems Science and Engineering, 5(1):36–41, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Deng:1991:RTB

[DDL91] X.-L. Deng, T. Dillon, and K. Lew. Reconfigurable transputer-based architecture. Computer Systems Science and Engineering, 6(3): 152–??, July 1991. CODEN CSSEEI. ISSN 0267-6192.

Desel:2000:HSL

[DE00] J. Desel and T. Erwin. Hybrid specifications: looking at workflows from a run-time perspective. International Journal of Computer Systems Science and Engineering, 15 (5):??, September 2000. CO-DEN CSSEEI. ISSN 0267-6192.

Datta:1997:SSS

[DFT97] Ajoy Kumar Datta, Mitchell Flatebo, and Visalakshi Thiagarajan. Simulation of selfstabilizing algorithms. International Journal of Computer Systems Science and Engineering, 12(5):295–306, September 1997. CODEN CSSEEI. ISSN 0267-6192.

Dowdy:1987:PSR

[DG87] Lawrence W. Dowdy and Karen D. Gordon. Performance sensitivity with respect to the distribution of the multiprogramming level. Computer Systems Science and Engineering, 2(2):92–98, April 1987. CODEN CSSEEI. ISSN 0267-6192.

De:2001:EFT

[DG01a] I. De and B. Gupta. Efficient fault tolerant multicasting in hybrid incomplete hypercube interconnection networks. International Journal of Computer Systems Science and Engineering, 16(1): ??, January 2001. CODEN CSSEEI. ISSN 0267-6192.

Dutta:2001:ETB

[DG01b] D. Dutta and B. Gupta.

An efficient time-based checkpoint and recovery algorithm for distributed systems involving mobile computers. International Journal of Computer Systems Science and Engineering, 16(4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Datta:1988:PCG

[DGGE88] Ajoy K. Datta, Sukumar Ghosh, Muhundan Gopalan, and Ahmed K. Elmagarmid. Pipelining and concurrency in graphics and imageprocessing applications. Computer Systems Science and Engineering, 3(3):107–116, July 1988. CODEN CSSEEI. ISSN 0267-6192.

DiGiandomenico:1988:GDA

[DGGS88] F. Di Giandomenico, M. L. Guidotti, F. Grandoni, and L. Simoncini. Gracefully degradable algorithm for Byzantine agreement. Computer Systems Science and Engineering, 3(1):32–40, January 1988. CODEN CSSEEI. ISSN 0267-6192.

DiGiandomenico:1989:EEB

[DGGS89] F. Di Giandomenico, M. L. Guidotti, F. Grandoni, and L. Simoncini. Evaluating the efficiency of Byzantine agreement algorithms. Computer Systems Science and Engineering, 4(1):3–11, 1989. CO-DEN CSSEEI. ISSN 0267-6192.

Datta:1988:DAM

[DGHE88] Ajoy K. Datta, Sukumar Ghosh, Douglas Harms, and Ahmed K. Elmagarmid. Deadlock-avoidance mechanisms in distributed systems. Computer Systems Science and Engineering, 3(2):67–82, April 1988. CODEN CSSEEI. ISSN 0267-6192.

Deiters:2000:FWM

[DGL00] W. Deiters, T. Goesmann, and T. Löffeler. Flexibility in workflow management: dimensions and solutions. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Damiani:2006:SIW

[DGM06] Ernesto Damiani, Gabriele Gianini, and Hiroshi Maruyama. Special issue: Web service security. International Journal of Computer Systems Science and Engineering, 21(5): ??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

DalCin:1996:MEM

[DHH96] M. Dal Cin, H. Hessenauer, and W. Hohl. The Modular Expandable Multiprocessor System, MEMSY. International Journal of Computer Systems Science and Engineering, 11(4):211–219, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Din:2004:SAA

[Din04] Der-Rong Din. A simulated annealing algorithm for finding minimal wavelength on a WDM ring. International Journal of Computer Systems Science and Engineering, 19 (4):??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Dash:1996:GNW

[DJ96] S. K. Dash and B. Jha. Global numerical weather model integrated on a transputer-based parallel computer. International Journal of Computer Systems Science and Engineering, 11(2):93–98, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Ding:2002:TPN

[DJ02] Z. Ding and C. Jiang. Temporal Petri nets model of concurrent systems. International Journal of Computer Systems Science and Engineering, 17 (6):??, November 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Didic:1986:DCD

[DKO86] Milena M. Didic, Peter Kohlhepp, and Reinhold Oberle. Design considerations for a distributed realtime unclear reactor safety system. Computer Systems Science and Engineering, 1 (2):82–92, January 1986. CO-DEN CSSEEI. ISSN 0267-6192.

Damodaran-Kamal:1995:SAC

[DKP95] Suresh K. Damodaran-Kamal and Niki Pissinou. Shared access and communication in distributed databases using an object environment. International Journal of Computer Systems Science and Engineering, 10(4):214–222, October 1995. CODEN CSSEEI. ISSN 0267-6192.

deLemos:2001:EHS

[dLR01] R. de Lemos and A. Romanovsky. Exception handling in the software lifecycle. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Dillon:2006:SIP

[DM06] Tharam Dillon and Mukesh Mohania. Special issue: Privacy data management. International Journal of Computer Systems Science and Engineering, 21(1):??, January 2006. CODEN CSSEEI. ISSN 0267-6192.

Deo:1998:LBP

[DMP98] Narsingh Deo, Muralidhar Medidi, and Sushil K. Prasad. Load balancing in parallel battlefield management simulation on local- and sharedmemory architectures. International Journal of Computer Systems Science and Engineering, 13(1):55–65, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Downs:1987:RPS

[Dow87] T. Downs. Reliability problems in software engineering — a review. Computer Systems Science and Engineering, 2(3):131–147, July 1987.

CODEN CSSEEI. ISSN 0267-6192.

Downs:1994:MST

[Dow94] Tom Downs. Modelling software testing for reliability prediction. Computer Systems Science and Engineering, 9(2):104–111, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Das:2009:CFR

[DR09] Gita Das and Sid Ray. A compact feature representation and feature re-weighting in content-based image retrieval. International Journal of Computer Systems Science and Engineering, 24(2): ??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Demaine:1997:RAS

[DS97] Erik Demaine and Sampalli Srinivas. Routing algorithms on static interconnection networks: a classification scheme. International Journal of Computer Systems Science and Engineering, 12(6):359– 367, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Donatelli:2000:NPC

[DSS00] S. Donatelli, M. Sarini, and C. Simone. Negotiating propagation of changes in inter-organizational workflows. International Journal of Computer Systems Science and Engineering, 15(5): ??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Dujmovic:1991:CCS

[Duj91] Jozo J. Dujmovic. Clustering, comparison and selection of standard synthetic benchmark programs. Computer Systems Science and Engineering, 6(4):195–210, October 1991. CODEN CSSEEI. ISSN 0267-6192.

Dias:1991:ATO

[DYB91] Daniel M. Dias, Philip S. Yu, and Brian T. Bennett. Analysis of trade-offs between centralized and geographically distributed transaction processing systems. Computer Systems Science and Engineering, 6(1):45–??, January 1991. CODEN CSSEEI. ISSN 0267-6192.

Demey:2015:SMS

[DZ15] Yan Tang Demey and Gang Zhao. SDD-matcher: a semantic-driven data matching framework. International Journal of Computer Systems Science and Engineering, 30 (3):??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Dubey:1992:RFT

[DZG92] A. Dubey, M. Zubair, and C. E. Grosch. Real Fourier transforms on a massively parallel machine. Computer Systems Science and Engineering, 7(4):243–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Evans:1997:AIP

[EG97] D. J. V. Evans and A. M. Goscinski. Automatic identification of parallel units and synchronization points in programs for parallel execution on a distributed system. International Journal of Computer Systems Science and Engineering, 12(5):307– 315, September 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Ennis:1986:AMO

[EGH^{+}86] R. L. Ennis, J. H. Griesmer, S. J. Hong, M. Karnaugh, J. K. Kastner, D. A. Klein, K. R. Milliken, M. I. Schor, and H. M. Van Woerkom. Automation of MVS operations; an expert-systems approach. Computer Systems Science and Engineering, 1 (2):119–124, January 1986. CODEN CSSEEI. ISSN 0267-6192.

Erradi:2006:SDP

[EKA06] A. Erradi, N. Kulkarni, and S. Anand. Service design principles: a case study in modeling services for the securities trading domain. International Journal of Computer Systems Science and Engineering, 21 (4):??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Ersoy:2000:AIT

[ELG00] C. Ersoy, A. Levi, and O. Gumrah. Artificial intelligence techniques for discrete link capacity assignment in prioritized multiservice networks. International Journal of Computer Systems Science and Engineering, 15(3): 191–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Elish:2013:OOS

[Eli13] Mahmoud O. Elish. objectoriented software. International Journal of Computer Systems Science and Engineering, 28(4):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Elmagarmid:1987:FTD

[Elm87] Ahmed K. Elmagarmid.

Fault-tolerant deadlock detection in distributed database systems. Computer Systems Science and Engineering, 2 (1):16–28, January 1987. CO-DEN CSSEEI. ISSN 0267-6192.

El-Qawasmeh:2011:CRE

[EQ11] Eyas El-Qawasmeh. Categorizing received email to improve delivery. International Journal of Computer Systems Science and Engineering, 26 (2):??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Enxing:2016:SGI

[ERY^{+}16] Zheng Enxing, Liu Ranran, Jiang Yifeng, Sun Yuxia, and Ni FuYin. Stochastic gradient identification for hammerstein systems with non-uniformly sampling. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Enokido:2010:PBS

[ET10] Tomoya Enokido and Makoto Takizawa. A purpose-based synchronization protocol for secure information flow control. International Journal of Computer Systems Science and Engineering, 25(2): ??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Fahmi:2001:DWB

[FAL^{+}01] H. Fahmi, W. G. Aref, M. Latif, B. Shafiq, A. Ghafoor, P. Liu, and L. Hsu. Distributed Web-based framework for real-time multimedia object communication. International Journal of Computer Systems Science and Engineering, 16(2):??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Feng:2012:IDU

[FBZS12] Zhenqian Feng, Bing Bai, Baokang Zhao, and Jinshu Su. ICTree: discovering the underlying connections of your rental virtual machines in the public clouds. International Journal of Computer Systems Science and Engineering, 27 (6):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Feng:1999:PLS

[FC99] M. D. Feng and C. K. Y. Chen. Parallel Lisp with speculation and subtuplespace on distributed systems. International Journal of Computer Systems Science and Engineering, 14(5):283–??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Feng:2003:STO

[FCD03] L. Feng, E. Chang, and T. S.

Dillon. Schemata transformation of object-oriented conceptual models to XML. International Journal of Computer Systems Science and Engineering, 18(1):??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Feng:2004:STO

[FCD04] L. Feng, E. Chang, and T. S.

Dillon. Schemata transformation of object-oriented conceptual models to XML. International Journal of Computer Systems Science and Engineering, 19(1):??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Fernandez:2013:EMC

[FCGC13] Javier Fernández, Alejandro Calderón, Félix García, and Jesús Carretero. Enhancing multimedia caching algorithms for variable-bandwidth streams using adaptive and smoothing techniques. International Journal of Computer Systems Science and Engineering, 28(5):??, ???? 2013. CODEN CSSEEI. ISSN 0267-

Feather:2005:ATS

[FCHJ05] Martin S. Feather, Steven L. Cornford, Kenneth A. Hicks, and Kenneth R. Johnson. Applications of tool support for risk-informed requirements reasoning. International Journal of Computer Systems Science and Engineering, 20(1):??, January 2005. CODEN CSSEEI. ISSN 0267-6192.

Fan:1997:PAA

[FG97] Yanhe Fan and Nicolas Georganas. Performance analysis of ATM switches with selfsimilar input traffic. International Journal of Computer Systems Science and Engineering, 12(2):95–102, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Friedman:2002:CSEa

[FH02a] R. Friedman and E. Hadad. Client-side enhancements using portable interceptors. International Journal of Computer Systems Science and Engineering, 17(2):??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Friedman:2002:CSEb

[FH02b] R. Friedman and E. Hadad. Client-side enhancements using portable interceptors. International Journal of Computer Systems Science and Engineering, 17(3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Fachrunnisa:2012:DE

[FH12] Olivia Fachrunnisa and Farookh Khadeer Hussain. digital enterprise. International Journal of Computer Systems Science and Engineering, 27 (3):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Fang:2013:ABC

[FH13] Jywe-Fei Fang and Chien-Hung Huang. Algorithmic bipanpath connectivity of the hypercube. International Journal of Computer Systems Science and Engineering, 28 (3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Fischer-Hubner:2007:SIT

[FHFL07] Simone Fischer-Hübner, Steven Furnell, and Costas Lambrinoudakis. Special issue: TrustBus 2006. International Journal of Computer Systems Science and Engineering, 22 (5):??, September 2007. CO-DEN CSSEEI. ISSN 0267-6192.

Finkel:1990:MDL

[Fin90]  David Finkel.  Modelling

dynamic load-sharing in a distributed computing system. Computer Systems Science and Engineering, 5(2): 89–94, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Feng:2004:PEX

[FJ04] L. Feng and W. Jonker.

Preparations for encrypted XML metadata querying. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Fu:2016:TSA

[FL16] Li Fu and Zhen Liu. Topological separation axioms of soft rough formal context. International Journal of Computer Systems Science and Engineering, 31(2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Francalanci:2010:EAB

[FM10] Chiara Francalanci and Francesco Merlo. Empirical analysis of the bug fixing process in open source projects. International Journal of Computer Systems Science and Engineering, 25 (4):??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Fooprateepsiri:2014:HRA

[FM14] Rerkchai Fooprateepsiri and Pichet Mareo. A highly robust approach to face authentication based on trace-fuzzy ART combination. International Journal of Computer Systems Science and Engineering, 29(3):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Finkel:1994:LSS

[FMP94] David Finkel, Xiannong Meng, and Sanjay Parikh. Load sharing that supports fault tolerance in a distributed computing system. International Journal of Computer Systems Science and Engineering, 9(4):246– 254, October 1994. CODEN CSSEEI. ISSN 0267-6192.

Frantzeskou:2011:SUD

[FMS^{+}11] Georgia Frantzeskou, Stephen G.

MacDonell, Efstathios Stamatatos, Stelios Georgiou, and Stefanos Gritzalis. The significance of user-defined identifiers in Java source code authorship identification. International Journal of Computer Systems Science and Engineering, 26(2):??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Ferscha:2013:E

[FP13] Alois Ferscha and Siani Pearson. Editorial. International Journal of Computer Systems Science and Engineering, 28 (6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Fong:1997:AMP

[FSA97] Simon Fong, Samar Singh, and Mohammed Atiquzzaman. Analytical model and performance analysis of shared buffer ATM switches under non-uniform traffic. International Journal of Computer Systems Science and Engineering, 12(2):81–93, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Ficco:2013:AID

[FVD13] Massimo Ficco, Salvatore Venticinque, and Beniamino Di Martino. An advanced intrusion detection framework for cloud computing. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Feng:2016:SVT

[FXZ16] Quan-Dong Feng, Miao Xu, and Xin Zhang. The scalable vocabulary tree based model for sub-image retrieval. International Journal of Computer Systems Science and Engineering, 31(6):??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Guler:2014:DFL

[GA14] Hasan Guler and Fikret Ata. Design of a fuzzy-labview-based mechanical ventilator. International Journal of Computer Systems Science and Engineering, 29(3):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Gabillon:2006:LFS

[Gab06] Alban Gabillon. A logical formalization of secure XML database. International Journal of Computer Systems Science and Engineering, 21(5): ??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

Grefen:2000:CCO

[GAHL00] P. Grefen, K. Aberer, Y. Hoffner, and H. Ludwig. CrossFlow: Cross-organizational workflow management in dynamic virtual enterprises. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Gait:1987:SMA

[Gai87]  Jason Gait.  Synchroniz-

ing multiprocessor access to shared operating system data structures. Computer Systems Science and Engineering, 2 (4):186–191, October 1987. CODEN CSSEEI. ISSN 0267-6192.

Gervasi:2012:P

[GATT12] Osvaldo Gervasi, Bernady O.

Apduhan, C. J. Kenneth Tan, and David Taniar. Preface. International Journal of Computer Systems Science and Engineering, 27(3): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Garcia:2013:UPR

[GCG^{+}13] J. Daniel Garcia, Jesus Carretero, Javier Garcia, Luis Miguel Sanchez, and Felix Garcia. by using partial replication. International Journal of Computer Systems Science and Engineering, 28 (3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Guntama:2003:EAD

[GCJP03] E. Guntama, E. Chang, N. Jayaratna, and L. Pudhota. Extension of activity diagrams for flexible business workflow modeling. International Journal of Computer Systems Science and Engineering, 18(3): ??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Grossi:2014:SRD

[GCM14] Lucas Grossi and Jose A. Calvo-Manzano. A systematic review in the decision analysis of selecting an IT outsourcing supplier. International Journal of Computer Systems Science and Engineering, 29 (3):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Guennoun:2006:UGG

[GD06] K. Guennoun and K. Drira. Using graph grammars for interaction style description: applications for serviceoriented architectures. International Journal of Computer Systems Science and Engineering, 21(4):??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Guha:1988:ACD

[GDK88] Ratan K. Guha, Arthur F. Dickinson, and Allan Kikawa. Analysis of CMH deadlockdetection algorithm using transmission delays. Computer Systems Science and Engineering, 3(4):181–188, October 1988. CODEN CSSEEI. ISSN 0267-6192.

Gal:1997:PEM

[GE97] Avigdor Gal and Opher Etzion. Parallel execution model for updating temporal databases. International Journal of Computer Systems Science and Engineering, 12 (5):317–327, September 1997. CODEN CSSEEI. ISSN 0267-6192.

Goudarzinemati:2010:SAC

[GET10] Alireza Goudarzinemati, Tomoya Enokido, and Makoto Takizawa. Scheduling algorithms for concurrently streaming multimedia objects in P2P overlay networks. International Journal of Computer Systems Science and Engineering, 25(2):??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Guo:2010:EAT

[GFH^{+}10] Zhongwen Guo, Yuan Feng, Lu Hong, Ying Guo, and Meng Yang. Efficient and adaptive transmission algorithms for underwater acoustic networks. International Journal of Computer Systems Science and Engineering, 25 (6):??, November 2010. CO-DEN CSSEEI. ISSN 0267-6192.

Guo:2016:REB

[GFZ^{+}16] Tao Guo, Xiang Fang, Mingshou Zhong, Huayuan Ma, and Yiming Mao. Research on the effects of blasting vibration attenuation by pre-split crack with HHT method. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Gupta:2001:BCI

[GGDG01] B. Gupta, K. Ghosh, D. Dutta, and A. Gupta. Broadcasting in complete and incomplete star interconnection networks. International Journal of Computer Systems Science and Engineering, 16(4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Gilles:2011:MBO

[GH11] Olivier Gilles and Jerome Hugues. A MDE-based optimisation process for real-time systems: optimizing systems at the architecture-level using the real DSL and library of transformation and heuristics. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Guo:2015:RRJ

[GHW15] Yandong Guo, Min Huang, and Qing Wang. Rescheduling rework jobs on single-machine of original jobs with release times. International Journal of Computer Systems Science and Engineering, 30(6):

??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Gorodetski:2003:MAT

[GKK03] V. Gorodetski, I. Kotenko, and O. Karsaev. Multiagent technologies for computer network security: attack simulation, intrusion detection and intrusion detection learning. International Journal of Computer Systems Science and Engineering, 18 (4):??, July 2003. CODEN CSSEEI. ISSN 0267-6192.

Grun:2006:SAL

[GKS06] Katharina Grün, Michael Karlinger, and Michael Schrefl. Schema-aware labelling of XML documents for efficient query and update processing in SemCrypt. International Journal of Computer Systems Science and Engineering, 21 (1):??, January 2006. CO-DEN CSSEEI. ISSN 0267-6192.

Garcia:2006:SIS

[GL06] Alessandro Garcia and Carlos Lucena. Special issue: Software engineering for multiagent systems. International Journal of Computer Systems Science and Engineering, 21 (2):??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Gait:1989:IMC

[GLK89] Jason Gait, Terry Laskodi, and Rick Krull. Interactive monitor for a communication kernel. Computer Systems Science and Engineering, 4 (3):154–160, July 1989. CO-DEN CSSEEI. ISSN 0267-6192.

Gnesi:2005:ATA

[GLT05] Stefania Gnesi, Giuseppe Lami, and Gianluca Trentanni. An automatic tool for the analysis of natural language requirements. International Journal of Computer Systems Science and Engineering, 20 (1):??, January 2005. CO-DEN CSSEEI. ISSN 0267-6192.

Gonzalez:2013:HBP

[GMC^{+}13] Apolinar Gonzalez, Walter Mata, Alfons Crespo, Miguel Masmano, José Félix, and Alvaro Aburto. A hypervisor based platform to support real-time safety critical embedded Java applications. International Journal of Computer Systems Science and Engineering, 28(3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Gelenbe:1996:DCL

[GMF96] Erol Gelenbe, Xiaowen Mang, and Yutao Feng. Diffusion cell loss estimate for ATM with multiclass bursty traffic. International Journal of Computer Systems Science and Engineering, 11(6):325– 333, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Gergeleit:1998:CMO

[GMN98] Martin Gergeleit, Michael Mock, and Edgar Nett. T-CORBA: Making objectoriented systems time-aware. International Journal of Computer Systems Science and Engineering, 13(3):151–160, May 1998. CODEN CSSEEI. ISSN 0267-6192.

Gershoni:2013:MPS

[GMP13] Tomer Gershoni, Miranda Mowbray, and Siani Pearson. Mechanisms for protecting sensitive information in cloud computing. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Guelfi:2008:SIE

[GMPR08] Nicolas Guelfi, Henry Muccini, Patrizio Pelliccione, and Alexander Romanovsky. Special issue: Engineering fault tolerant systems. International Journal of Computer Systems Science and Engineering, 23(5):??, September 2008. CODEN CSSEEI. ISSN 0267-6192.

Guo:1990:CCG

[GP90] Qingping Guo and Yakup Paker. Concurrent communication and granularity assessment for a transputer-based multiprocessor system. Computer Systems Science and Engineering, 5(1):21–28, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Gentile:2010:KDL

[GPR10] Antonio Gentile, Roberto Pirrone, and Giuseppe Russo. Knowledge discovery for the Linguistic Atlas of Sicily Project. International Journal of Computer Systems Science and Engineering, 25(2): ??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Guan:2014:MAM

[GQW^{+}14] Jianfeng Guan, Wei Quan, Lili Wang, Changqiao Xu, Feilong Tang, and Hongke Zhang. Modeling and analysis of mobility stochastic properties in cognitive radio networks. International Journal of Computer Systems Science and Engineering, 29(6): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Gini:1994:CCB

[GR94] Giuseppina C. Gini and C. Rogialli. CONSTRIC-TOR: a constraint-based language. International Journal of Computer Systems Science and Engineering, 9(4): 255–261, October 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Gunavathi:2001:CAN

[GS01] K. Gunavathi and A. Shanmugam. Comparative analysis of new multiple access protocols for bus LANs. International Journal of Computer Systems Science and Engineering, 16(6):??, November 2001. CODEN CSSEEI. ISSN 0267-6192.

Guo:2010:CRP

[GS10] Song Guo and An-Ni Shen.

A compromise-resilient pairwise rekeying protocol in hierarchical wireless sensor networks. International Journal of Computer Systems Science and Engineering, 25(6): ??, November 2010. CODEN CSSEEI. ISSN 0267-6192.

Goel:2003:MDS

[GST03] S. Goel, H. Sharda, and D. Taniar. Messaging in distributed systems. International Journal of Computer Systems Science and Engineering, 18(6):??, December 2003. CODEN CSSEEI. ISSN 0267-6192.

Gunter:1993:DIA

[Gun93] Willi Gunter. Design and implementation of the Attempto fault-tolerant system. Computer Systems Science and Engineering, 8(2):101– 108, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Gutjahr:2001:RMN

[Gut01] W. J. Gutjahr. A reliability model for nonhomogeneous redundant software versions with correlated failures. International Journal of Computer Systems Science and Engineering, 16(6):??, November 2001. CODEN CSSEEI. ISSN 0267-6192.

Garcia-Valls:2013:BDQ

[GVBVME13] Marisol García-Valls, Pablo Basanta-Val, Marga Marcos, and Elisabet Estévez. A bidimensional QoS model for SOA and real-time middleware. International Journal of Computer Systems Science and Engineering, 28(5): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Garcia-Valls:2014:BDQ

[GVBVME14] Marisol García-Valls, Pablo Basanta-Val, Marga Marcos, and Elisabet Estévez. A bidimensional QoS model for SOA and real-time middleware. International Journal of Computer Systems Science and Engineering, 29(5): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Garcia-Valls:2013:RMM

[GVCV13] Marisol García-Valls, Alfons Crespo, and Joan Vila. Resource management for mobile operating systems based on the active object model. International Journal of Computer Systems Science and Engineering, 28(4): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Giese:2001:OAO

[GW01] H. Giese and G. Wirtz. The OCoN approach for objectoriented distributed software systems modeling. International Journal of Computer Systems Science and Engineering, 16(3):??, May 2001. CODEN CSSEEI. ISSN 0267-6192.

Guttmann:2006:ALM

[GZ06] Christian Guttmann and Ingrid Zukerman. Agents with limited modeling abilities: Implications on collaborative problem solving. International Journal of Computer Systems Science and Engineering, 21(3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Ghose:1995:PPS

[GZG95] M. Ghose, M. Zubair, and C. E. Grosch. Parallel partitioning of sparse matrices. International Journal of Computer Systems Science and Engineering, 10(1):33– 40, January 1995. CODEN CSSEEI. ISSN 0267-6192.

Hooshmand:2008:RCG

[HA08] R. Hooshmand and M. Ataei.

Real-coded genetic algorithm applied to optimal placement of capacitor banks for unbalanced distribution systems with meshed/ radial configurations. International Journal of Computer Systems Science and Engineering, 23(4): ??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Heitmeyer:2005:TCR

[HABJ05] C. Heitmeyer, M. Archer, R. Bharadwaj, and R. Jeffords. Tools for constructing requirements specifications: the SCR toolset at the age of nine. International Journal of Computer Systems Science and Engineering, 20(1): ??, January 2005. CODEN CSSEEI. ISSN 0267-6192.

Halang:2009:SIA

[Hal09] Wolfgang A. Halang. Special issue: Automated control systems for the safety integrity levels 3 and 4. International Journal of Computer Systems Science and Engineering, 24 (1):??, January 2009. CO-DEN CSSEEI. ISSN 0267-6192.

Hameurlain:2005:E

[Ham05] Abdelkader Hameurlain. Editorial. International Journal of Computer Systems Science and Engineering, 20(2): ??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Hameurlain:2008:SID

[Ham08] Abdelkader Hameurlain. Special issue: Data management in grid and P2P systems: Editorial. International Journal of Computer Systems Science and Engineering, 23(2): ??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Hu:2016:RMA

[HBL16] Tian Run Hu, Zhang Yan Bin, and Jia Lixin. The realization of multi agent based online trading systems based on J2EE technology. International Journal of Computer Systems Science and Engineering, 31(6):??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Hulina:1990:CAE

[HC90] Paul T. Hulina and Lee D.

Coraor. Coprocessor architectures for efficient address computation and memory accessing. Computer Systems Science and Engineering, 5 (3):137–146, July 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Hadzic:2008:AMD

[HC08] Maja Hadzic and Elizabeth Chang. Onto-agent methodology for design of ontology-based multi-agent systems. International Journal of Computer Systems Science and Engineering, 23(1): ??, January 2008. CODEN CSSEEI. ISSN 0267-6192.

Hsiao:2013:TEC

[HC13] Wen-Feng Hsiao and Te-Min Chang. Term expansion on the categorization of summarized documents. International Journal of Computer Systems Science and Engineering, 28(4):??, ???? 2013. CODEN CSSEEI. ISSN 0267-

Hwang:1998:EMS

[HCC98] Shin-Jia Hwang, Chien-Yuang Chen, and Chin-Chen Chang. An encryption/ multisignature scheme with specified receiving groups. International Journal of Computer Systems Science and Engineering, 13(2):109–112, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Hussain:2004:CTP

[HCD04a] F. K. Hussain, E. Chang, and T. S. Dillon. Classification of trust in Peer-to-Peer (P2P) communication. International Journal of Computer Systems Science and Engineering, 19 (2):??, March 2004. CODEN CSSEEI. ISSN 0267-6192.

Hussain:2004:TCM

[HCD04b] F. K. Hussain, E. Chang, and T. S. Dillon. Trustworthiness and CCCI metrics in P2P communication. International Journal of Computer Systems Science and Engineering, 19 (3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Han:2008:PBPa

[HCD08a] Song Han, Elizabeth Chang, and Tharam Dillon. Pairingbased public-key encryption schemes with backward-and-forward security. International Journal of Computer Systems Science and Engineering, 23(1):??, January 2008. CODEN CSSEEI. ISSN 0267-6192.

Han:2008:PBPb

[HCD08b] Song Han, Elizabeth Chang, and Tharam Dillon. Pairingbased public-key encryption schemes with backward-and-forward security. International Journal of Computer Systems Science and Engineering, 23(4):??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Hussain:2006:MRM

[HCHD06] Omar Khadeer Hussain, Elizabeth Chang, Farookh Khadeer Hussain, and Tharam S. Dillon. A methodology for risk measurement in etransactions. International Journal of Computer Systems Science and Engineering, 21 (1):??, January 2006. CO-DEN CSSEEI. ISSN 0267-6192.

Ho:1990:BTC

[HCK90] Hong-Fa Ho, Gen-Heuy Chen, and Te-Son Kuo. Branch testing of concurrent programs using Petri net models. Computer Systems Science and Engineering, 5(2): 116–125, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Huang:2006:SDT

[HCL^{+}06] C-L Huang, K-M Chao, C-C Lo, P. Wang, and J-Y Chung. Service discovery through consensus based preferences. International Journal of Computer Systems Science and Engineering, 21(4):

??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Hughes:1987:GBP

[HCY87] Herman Hughes, Ke Ren Chuang, and Parviz Yegani. Give-up-based protocol for token passing. Computer Systems Science and Engineering, 2(2):85–91, April 1987. CODEN CSSEEI. ISSN 0267-6192.

Hussain:2010:TRB

[HDCH10] Pmar Khadeer Hussain, Tharam Dillon, Elizabeth Chang, and Farookh Hussain. Transactional risk-based decision making system in e-business interactions. International Journal of Computer Systems Science and Engineering, 25 (1):??, January 2010. CO-DEN CSSEEI. ISSN 0267-6192.

Han:2013:RMA

[HDPC13] Song Han, Tharam Dillon, Vidy Potdar, and Elizabeth Chang. RFID mutual authentication protocols for tags and readers with and without a server. International Journal of Computer Systems Science and Engineering, 28 (2):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Helmy:2013:PBS

[Hel13] Tarek Helmy. Priority-based scheduling for dependent processes using random selection of resource holders. International Journal of Computer Systems Science and Engineering, 28(3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Ho:1994:VVH

[HF94] Carlson Ho and Kevin E. Forward. Verification and validation of hardware designs via hardware Petri nets. Computer Systems Science and Engineering, 9(1):65–72, January 1994. CODEN CSSEEI. ISSN 0267-6192.

He:2008:AXQ

[HF08] W. He and L. Fegaras. Answering XPath queries with search predicates in structured P2P networks. International Journal of Computer Systems Science and Engineering, 23(2):??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Horng:2008:ADW

[HFS^{+}08] Shi-Jinn Horng, Pingzhi Fan, Ming-Yang Su, Yuan-Hsin Chen, Cheng-Ling Lee, and Shao-Wei Lan. Anomaly detection for Web server based on smooth support vector machine. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Hasan:2012:CBH

[HGR12] Mahmudul Hasan, Marina L.

Gavrilova, and Jon G. Rokne. Clearance-based homotopic optimal path computation. International Journal of Computer Systems Science and Engineering, 27(3): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Harrag:2011:EEV

[HHCASEQ11] Fouzi Harrag, Aboubekeur Hamdi-Cherif, Abdul Malik S. Al-Salman, and Eyas El-Qawasmeh. Evaluating the effectiveness of VSM model and topic segmentation in retrieving Arabic documents. International Journal of Computer Systems Science and Engineering, 26(1):??, January 2011. CODEN CSSEEI. ISSN 0267-6192.

Hammer:1998:MBD

[HHD98] Dieter K. Hammer, Andrew A. Hanish, and Tharam S. Dillon. Modeling behavior and dependability of object-oriented real-time systems. International Journal of Computer Systems Science and Engineering, 13(3): 139–150, May 1998. CODEN CSSEEI. ISSN 0267-6192.

Hines:1999:OOO

[Hin99] M. L. Hines. Ontology of object-oriented database evolution. International Journal of Computer Systems Science and Engineering, 14(6):371– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Hughes:1987:VDI

[HJS87] Herman Hughes, Horng Jiang, and Susanne Smith.

Voice/data integrated transmission on a time division CSMA/CD network. Computer Systems Science and Engineering, 2(3):125–130, July 1987. CODEN CSSEEI. ISSN 0267-6192.

Han:2012:NIL

[HJXY12] Yu Han, Naigao Jin, Guangwei Xu, and Kangbin Yim. A new iterative learning control scheme using the geometric rotation for discrete time systems. International Journal of Computer Systems Science and Engineering, 27(5): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Hogg:1990:PTH

[HK90] Tad Hogg and J. O. Kephart.

Phase transitions in highdimensional pattern classification. Computer Systems Science and Engineering, 5 (4):223–232, October 1990. CODEN CSSEEI. ISSN 0267-6192.

Hirata:1998:NMA

[HK98] Celso Hirata and Jeff Kramer.

A novel modelling approach for parallel simulation. International Journal of Computer Systems Science and Engineering, 13(1):27–37, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Hammer:1999:SIO

[HKL99] Dieter K. Hammer, Tohru Kikuno, and Insup Lee. Special issue: Object-oriented real-time distributed computing. International Journal of Computer Systems Science and Engineering, 14(4): ??, July 1999. CODEN CSSEEI. ISSN 0267-6192.

Hsu:1999:PEP

[HL99] C.-C. Hsu and S.-S. Lin. Path establishment of permutations in a rearrangeable interconnection network. International Journal of Computer Systems Science and Engineering, 14(5):267–??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Hu:2016:IQM

[HL16] Yanling Hu and Anfeng Liu. Improving the quality of mobile target detection through portion of node with full duty cycle in WSNs. International Journal of Computer Systems Science and Engineering, 31 (1):??, January 2016. CO-DEN CSSEEI. ISSN 0267-6192.

Hwang:1996:SCW

[HLLC96] Tzonelih Hwang, Narn-Yih Lee, Chuan-Ming Li, and Chin-Chen Chang. On the security of Chang and Wu’s broadcasting cryptosystem for computer networks. International Journal of Computer Systems Science and Engineering, 11(5):311–314, September 1996. CODEN CSSEEI. ISSN 0267-6192. See [CW91, LC96a].

Hurson:1987:DMA

[HM87] A. R. Hurson and L. L. Miller.

Database machine architecture for supporting incomplete information. Computer Systems Science and Engineering, 2(3):107–116, July 1987. CODEN CSSEEI. ISSN 0267-6192.

Hameurlain:2004:PQO

[HM04] Abdelkader Hameurlain and Franck Morvan. Parallel query optimization methods and approaches: a survey. International Journal of Computer Systems Science and Engineering, 19(5): ??, September 2004. CODEN CSSEEI. ISSN 0267-6192.

Ho:1993:PMM

[HMF93] C. Ho, S. A. Mabbs, and K. E.

Forward. Performance modeling of the MR-1 multiprocessor using extended deterministic and stochastic Petri nets. Computer Systems Science and Engineering, 8(4): 195–209, October 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Hurson:1991:FDE

[HMPF91] A. R. Hurson, L. L. Miller, S. H. Pakzad, and C. Fan. Functional dependencies to enforce integrity constraints in database machine environments. Computer Systems Science and Engineering, 6 (2):91–??, April 1991. CO-DEN CSSEEI. ISSN 0267-6192.

Hiraishi:1994:EII

[HN94] Hiromi Hiraishi and Tatsuya Nakae. Efficient inverse image computation algorithm for sequential machine verification using temporal logics. Computer Systems Science and Engineering, 9(2): 112–117, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Hoffman:1993:IPE

[Hof93] Lorrie L. Hoffman. Improved performance evaluation of distributed mutual exclusion algorithms for computer networks. Computer Systems Science and Engineering, 8 (2):67–71, April 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Hogben:2004:FSA

[Hog04] G. Hogben. The formation of software agreements between autonomous agents. International Journal of Computer Systems Science and Engineering, 19(3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Henkler:2011:MDR

[HOGS11] Stefan Henkler, Simon Oberthür, Holger Giese, and Andreas Seibel. Model-driven runtime resource predictions for advanced mechatronic systems with dynamic data structures. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Hong:1987:ECD

[HP87] Yang-Chang Hong and Thomas H.

Payne. Efficient computation of dataflow graphs in a hypercube architecture. Computer Systems Science and Engineering, 2(1):29–41, January 1987. CODEN CSSEEI. ISSN 0267-6192.

Halang:2002:SOO

[HPF02] W. A. Halang, C. E. Pereira, and A. H. Frigeri. Safe object oriented programming of distributed real time systems in PEARL. International Journal of Computer Systems Science and Engineering, 17(2): ??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Hayati:2011:CWS

[HPTC11] Pedram Hayati, Vidyasagar Potdar, Alex Talevski, and Kevin Chai. Characterisation of Web spambots using self organising maps. International Journal of Computer Systems Science and Engineering, 26 (2):??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Hull:1987:ICS

[HRG87] M. E. C. Hull, R. G. Roulston, and T. R. S. Gregg. Interactive computing service monitoring system. Computer Systems Science and Engineering, 2(4):179–185, October 1987. CODEN CSSEEI. ISSN 0267-6192.

Haring:1985:MRD

[HS85] G. Haring and H. Schelch. Modelling of RPS disc systems. Computer Systems Science and Engineering, 1(1): 53–64, October 1985. CO-DEN CSSEEI. ISSN 0267-6192.

Harrison:1988:SSF

[HS88] E. S. Harrison and E. J. Schmitt. Structure of System/88, a fault-tolerant computer. Computer Systems Science and Engineering, 3(3): 140–162, July 1988. CODEN CSSEEI. ISSN 0267-6192.

Hiltunen:1996:ADF

[HS96] Matti A. Hiltunen and Richard D. Schlichting. Adaptive distributed and faulttolerant systems. International Journal of Computer Systems Science and Engineering, 11(5):275–285, September 1996. CODEN CSSEEI. ISSN 0267-6192.

Hill:2011:OAD

[HSS^{+}11] James Hill, Hunt Sutherlandy, Paul Staudingery, Thomas Silveriaz, Douglas C. Schmidt, John Slabyz, and Nikita Visnevskiy. OASIS: an architecture for dynamic instrumentation of enterprise distributed real-time and embedded systems. International Journal of Computer Systems Science and Engineering, 26(6):??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Hsu:1994:APT

[Hsu94] Chiun-Chieh Hsu. On the allocation of processors to tasks with precedence constraints and communication overhead. International Journal of Computer Systems Science and Engineering, 9(3): 164–174, July 1994. CODEN CSSEEI. ISSN 0267-6192.

Hsu:1999:ERB

[Hsu99] C.-C. Hsu. Embedability and recovery of a bubblesort star graph interconnection network. International Journal of Computer Systems Science and Engineering, 14 (5):307–??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Hsu:2008:CEC

[Hsu08] Chien-Lung Hsu. Cryptanalysis of elliptic curve cryptosystem on smart card access control with threshold scheme. International Journal of Computer Systems Science and Engineering, 23(6): ??, November 2008. CODEN CSSEEI. ISSN 0267-6192.

Hongfeng:2008:ECP

[HTJ08] Zhu Hongfeng, Liu Tianhua, and Liu Jie. EV-C2C-PAKE: An improved client-to-client password-authenticated key exchange protocol. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Hui:2000:PAF

[Hui00]  L. C. Hui.  A practical

algorithm to find longest common substring in linear time. International Journal of Computer Systems Science and Engineering, 15(2):73– ??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Huang:2016:LTR

[HWH^{+}16] Wen-Kai Huang, Chang-Dong Wang, Shao-Shu Huang, Zheng Li, Jian-Huang Lai, and Ling Huang. Long-term revenue maximization pricing scheme for cloud. International Journal of Computer Systems Science and Engineering, 31(2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Hsu:1997:ELA

[HWL97] Chiun-Chieh Hsu, Chien-Min Wang, and Jeen-Fong Lin. Efficient loop allocation based on the analysis of processor requirement of loops with equivalent completion times. International Journal of Computer Systems Science and Engineering, 12 (1):53–60, January 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Hughes:1988:CCP

[HYY88] Herman Hughes, Parviz Yegani, and Jih-Shyr Yih. CSMA/CD protocol with message-based priority function. Computer Systems Science and Engineering, 3(1): 13–20, January 1988. CO-DEN CSSEEI. ISSN 0267-6192.

Huachun:2012:MMA

[HZS12] Huachun, Hongke Zhang, and Fei Song. Mobility management analysis of Internet based on the splitting mechanism. International Journal of Computer Systems Science and Engineering, 27(6): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Iwashige:2010:EFD

[IB10] Jiro Iwashige and Leonard Barolli. Electromagnetic fields diffracted by two horizontal edges with arbitrary angle for different and same heights. International Journal of Computer Systems Science and Engineering, 25(2): ??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Iyer:1988:PCC

[IDY88] Balakrishna R. Iyer, Daniel M.

Dias, and Philip S. Yu. Performability comparison of configurable duplex systems. Computer Systems Science and Engineering, 3(4):201– 215, October 1988. CODEN CSSEEI. ISSN 0267-6192.

Ismail:2014:LLM

[IMA14] Muhamman Ali Ismail, Shahid H. Mirza, and Talat Altaf. ‘LogN+1’ and ‘LogN’ model: a binary tree based multilevel cache system for multicore processors. International Journal of Computer Systems Science and Engineering, 29 (2):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Ignatius:1997:TAH

[IS97] Paul P. Ignatius and C. Siva Ram Murthy. On task allocation in heterogeneous distributed computing systems. International Journal of Computer Systems Science and Engineering, 12(4):231– 238, July 1997. CODEN CSSEEI. ISSN 0267-6192.

Idris:2013:EVP

[IST^{+}13] Mohd Yamani Idna Idris, Moraali Sivalingam, Emran Mohd Tamil, Zaidi Razak, and Noorzaily Mohamed Noor. Emergency vehicle preemption system based on global positioning system (GPS), A-star (A*) algorithm and FPGA. International Journal of Computer Systems Science and Engineering, 28(3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Itaya:2007:PTM

[IT07] S. Itaya and M. Takizawa. Parallel transmission model for real-time multi-source streaming of multimedia data in scalable networks. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Ito:1994:WND

[Ito94] H. Ito. WSI network design and its evaluation using yields. Computer Systems Science and Engineering, 9(1):31–37, January 1994. CODEN CSSEEI. ISSN 0267-6192.

Iyer:1987:AFT

[IYD87] Balakrishna R. Iyer, Philip S.

Yu, and Lorenzo Donatiello. Analysis of fault tolerant multiprocessor architectures for lock engine design. Computer Systems Science and Engineering, 2(2):59–75, April 1987. CODEN CSSEEI. ISSN 0267-6192.

Jia:1997:DOS

[JC97] Xiaohua Jia and Wilson Cheng. Distributed operating systems vs. distributed system environments. International Journal of Computer Systems Science and Engineering, 12(3):193–203, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Jan:2002:SAV

[JC02] J-K Jan and Y-Y Chen. A secure anonymous voting protocol with a complete supervision. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Javagal:1992:ARD

[JDG92] Ramesh Dutt Javagal, Ajoy Kumar Datta, and Sukumar Ghosh. An algorithm for resource deadlock detection in distributed systems. Computer Systems Science and Engineering, 7(4):257–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Jensen:2001:DRT

[Jen01] E. D. Jensen. The distributed real-time specification for Java: an initial proposal. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Juidette:2003:QRD

[JIB03] H. Juidette, N. Idboufker, and A. Berraissoul. QoS routing in data networking using genetic algorithms: a new twolevel approach. International Journal of Computer Systems Science and Engineering, 18 (2):??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Jeong:2016:EIS

[JKC16] Seungdo Jeong, Sang-Wook Kim, and Byung-Uk Choi. Effective indexing and searching with dimensionality reduction on high-dimensional space. International Journal of Computer Systems Science and Engineering, 31(4): ??, July 2016. CODEN CSSEEI. ISSN 0267-6192.

Kim:2012:DBA

[jKsJdChK12] Min jun Kim, Yong suk Jang, Yong do Choi, and Sung ho Kim. Differentiated bandwidth allocation and wavelength assignment method in WDM EPON. International Journal of Computer Systems Science and Engineering, 27 (5):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Juang:2001:FBT

[JLL01] W.-S. Juang, C.-L. Lei, and H.-T. Liaw. Fair blind threshold signatures based on discrete logarithm. International Journal of Computer Systems Science and Engineering, 16 (6):??, November 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Jayaprakash:1999:PBF

[JM99] S. Jayaprakash and C. R.

Muthukrishnan. Permissionbased fault-tolerant distributed mutual exclusion algorithm. International Journal of Computer Systems Science and Engineering, 14(1):51–60, January 1999. CODEN CSSEEI. ISSN 0267-6192.

Joeris:2000:MFW

[Joe00]  G. Joeris.  Modeling of

flexible workflows and their decentralized enactment in flow.net. International Journal of Computer Systems Science and Engineering, 15 (5):??, September 2000. CO-DEN CSSEEI. ISSN 0267-6192.

Johnson:1996:IMA

[Joh96] D. M. Johnson. Integrated modular avionics: a scheme for autonomous dynamic system reconfiguration. International Journal of Computer Systems Science and Engineering, 11(3):125–133, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Jonker:2003:XSD

[Jon03] W. Jonker. XML and secure data management in an ambient world. International Journal of Computer Systems Science and Engineering, 18(5): ??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Jones:2000:DIM

[JOR^{+}00] K. H. Jones, S. Olariu, L. F. Rowell, J. L. Schwing, and A. Wilhite. A database integrity monitor with applications to vehicular technology. International Journal of Computer Systems Science and Engineering, 15(6): ??, December 2000. CODEN CSSEEI. ISSN 0267-6192.

Jimenez-Peris:2000:TAO

[JPMAB00] R. Jiménez-Peris, M. Patiño Martínez, S. Arévalo, and F. J. Ballesteros. TransLib: an Ada 95 object oriented framework for building transactional applications. International Journal of Computer Systems Science and Engineering, 15(1):7–??, January 2000. CODEN CSSEEI. ISSN 0267-6192.

Jia:1996:AAR

[JSM96] Xiaohua Jia, Kentaro Shimizu, and Mamoru Maekawa. Atomic accesses to replicated files in reliable distributed file systems. International Journal of Computer Systems Science and Engineering, 11(3): 151–158, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Jurczyk:1998:SII

[JSS^{+}98] Michael Jurczyk, Thomas Schwederski, Howard Jay Siegel, Seth Abraham, and Richard Martin Born. Strategies for the implementation of interconnection network simulators on parallel computers. International Journal of Computer Systems Science and Engineering, 13(1): 5–16, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Jiang:2001:HPD

[JTL01] Y. Jiang, D. Taniar, and C. H. C. Leung. High performance distributed parallel query processing. International Journal of Computer Systems Science and Engineering, 16(5):??, September 2001. CODEN CSSEEI. ISSN 0267-6192.

Jong:1995:AGC

[JTN95] C. C. Jong, G. N. Toh, and L. S. Ng. Automatic generation of customized functional models for rapid implementation of digital integrated circuits. International Journal of Computer Systems Science and Engineering, 10 (3):156–163, July 1995. CO-DEN CSSEEI. ISSN 0267-6192.

Jia:2015:DRA

[JWH^{+}15] Xiaohua Jia, Jinhai Wang, Chuanhe Huang, Qin Liu, Kai He, Jing Wang, and Peng Li. Dynamic resource allocation based on energy utility maximization using virtual machines in cloud environment. International Journal of Computer Systems Science and Engineering, 30(6): ??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Jiang:2014:MGR

[JWZ14] Wei Jiang, Cong Wang, and Yong-Qing Zhang. Multiple graph regularization constrained nonnegative matrix factorization. International Journal of Computer Systems Science and Engineering, 29 (4):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Kim:1993:CED

[KA93] Namhyung Kim and Vinod K. Agarwal. Concurrent error detection schemes for the match function. Computer Systems Science and Engineering, 8(1):52–56, January 1993. CODEN CSSEEI. ISSN 0267-6192.

Kumar:1997:SBD

[KA97] Sanjeev Kumar and Dharma P.

Agrawal. The shared-buffer direct-access (SBDA) ATM switch architecture for broadband networks. International Journal of Computer Systems Science and Engineering, 12 (2):69–79, March 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Kamal:1993:MMI

[Kam93] A. E. Kamal. Modelling multibus interconnection networks using state aggregation. Computer Systems Science and Engineering, 8(1): 57–63, January 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Kapralski:1992:APN

[Kap92] Adam Kapralski. Algorithms for processing NP-complete and isomorphic complete problems. Computer Systems Science and Engineering, 7 (4):218–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Karatza:1986:SSM

[Kar86] Helen D. Karatza. Simulation study of multitasking of open computer system networks. Computer Systems Science and Engineering, 1 (4):193–204, October 1986.

CODEN CSSEEI. ISSN 0267-6192.

Karatza:1987:SSM

[Kar87] Helen D. Karatza. Simulation study of multitasking of closed computer system networks. Computer Systems Science and Engineering, 2 (1):42–51, January 1987. CO-DEN CSSEEI. ISSN 0267-6192.

Karatza:1991:SLB

[Kar91] Helen D. Karatza. Simulation of load balancing and multitasking in a homogeneous distributed system model. Computer Systems Science and Engineering, 6(1):37–??, January 1991. CODEN CSSEEI. ISSN 0267-6192.

Katsinis:1992:PMM

[Kat92] Constantine Katsinis. Performance modelling of messagebased multiprocessors under heavy traffic. Computer Systems Science and Engineering, 7(3):190–198, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Katasonov:2005:UCD

[Kat05]  Artem Katasonov.  User-

centric data querying for location-based services. International Journal of Computer Systems Science and Engineering, 20(2):??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Kato:2014:SIO

[Kat14] Atsushi Kato. Software industry in Okinawa: is “domestic offshore outsourcing” scheme feasible? International Journal of Computer Systems Science and Engineering, 29(1): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Kim:2002:DRT

[KB02] K. H. (Kane) Kim and L. Bacellar. The distributed realtime conversation scheme for time-bounded cooperative recovery. International Journal of Computer Systems Science and Engineering, 17(1): ??, January 2002. CODEN CSSEEI. ISSN 0267-6192.

Kumar:2009:MQT

[KBB09] Sandeep Kumar, P. K.

Bansal, and Seema Bawa. Modified quad tree network with architectural analysis (M QUAD). International Journal of Computer Systems Science and Engineering, 24 (2):??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Kim:1994:MIM

[KBG^{+}94] K. H. Kim, L. F. Bacellar, J. Goldberg, K. Masui, K. Mori, and R. Yoshizawa. Modular implementation model for real-time fault-tolerant LAN systems based on the DRB scheme with a configuration supervisor. Computer Systems Science and Engineering, 9(2):75–82, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Katsinis:1994:EQS

[KC94] Constantine Katsinis and Amelie Constant. Effect of queue selection and service time distributions in multiple resource allocation. International Journal of Computer Systems Science and Engineering, 9(3):184–194, July 1994. CODEN CSSEEI. ISSN 0267-6192.

Kim:2010:NDG

[KC10] Young-Chang Kim and Jae-Woo Chang. A new distributed grid scheme and k-NN query processing for mobile objects in spatial networks. International Journal of Computer Systems Science and Engineering, 25(3): ??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Kumar:2005:MMT

[KDP05] Vijay Kumar, Margaret H. Dunham, and Nitin Prabhu. Mobilaction: a mobile transaction framework supporting spatial replication and spatial consistency. International Journal of Computer Systems Science and Engineering, 20 (2):??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Kodali:2004:SMA

[KFW04] N. Kodali, C. Farkas, and D. Wijesekera. Specifying multimedia access control using RDF. International Journal of Computer Systems Science and Engineering, 19(3): ??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Kodali:2006:UTS

[KFW06] N. Kodali, C. Farkas, and D. Wijesekera. uSMIL: Towards secure unbreakable semantics in multimedia Web services. International Journal of Computer Systems Science and Engineering, 21(5): ??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

Kim:2007:SIR

[KG07] Doohyun Kim and Arif Ghafoor. Special issue: Real time software architecture. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Katsinis:1996:CMA

[KH96] Constantine Katsinis and Pat Hogan. Concurrent multitasking applications on a shared-bus multicomputer system. International Journal of Computer Systems Science and Engineering, 11(3): 169–176, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Khanh:2006:PSS

[Kha06a] Dang Tran Khanh. A practical solution to supporting oblivious basic operations on dynamic outsourced search trees. International Journal of Computer Systems Science and Engineering, 21(1): ??, January 2006. CODEN CSSEEI. ISSN 0267-6192.

Klugl:2006:CGM

[KHA06b] Franziska Klügl, Rainer Herrler, and Gustavo Kuhn Andriotti. Coupling GIS and multi-agent simulation towards infrastructure for realistic simulation. International Journal of Computer Systems Science and Engineering, 21 (3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Kopetz:2001:UST

[KHE01] H. Kopetz, M. Holzmann, and W. Elmenreich. A universal smart transducer interface: TTP/A. International Journal of Computer Systems Science and Engineering, 16 (2):??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Kim:2000:VVO

[KHL00] J. Kim, T. Han, and S. K. Lee. VOQL: a Visual Objectoriented database Query Language for visualizing path expressions. International Journal of Computer Systems Science and Engineering, 15(4): 215–??, July 2000. CODEN CSSEEI. ISSN 0267-6192.

Kim:2008:PBL

[KHL^{+}08] H-S Kim, C-S Hwang, S-K Lee, S-J Choi, and J-M Gil. Priority based list scheduling for sabotage-tolerance with deadline tasks in desktop grids. International Journal of Computer Systems Science and Engineering, 23(2): ??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Kim:1998:MMS

[KHMK98] Jung-Guk Kim, Jin Pyo Hong, Byoung-Joon Min, and Moon Hae Kim. Modeling of multimedia services using the TMO model. International Journal of Computer Systems Science and Engineering, 13 (3):171–178, May 1998. CO-DEN CSSEEI. ISSN 0267-6192.

Khan:2009:EOS

[KHR^{+}09] Behram Khan, Matthew Horsnell, Ian Rogers, Mikel Luján, Andrew Dinn, and Ian Watson. Exploiting object structure in hardware transactional memory. International Journal of Computer Systems Science and Engineering, 24(5):??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Kim:2009:HAD

[KI09] K. H. (Kane) Kim and Chansik Im. Hybrid approaches for derivation of tight service time bounds of distributed embedded computing systems. International Journal of Computer Systems Science and Engineering, 24 (1):??, January 2009. CO-DEN CSSEEI. ISSN 0267-

Kim:2003:DHP

[Kim03]  S-W Kim.  Development

of a high-performance index manager in a main memory DBMS. International Journal of Computer Systems Science and Engineering, 18(2): ??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Kim:2004:CCM

[Kim04] Sang-Wook Kim. Concurrency control in main-memory DBMS. International Journal of Computer Systems Science and Engineering, 19(4): ??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Kusumoto:1994:ISD

[KiMKT94] Shinji Kusumoto, Ken ichi Matsumoto, Tohru Kikuno, and Kazuhiro Tanaka. Improvement of the software development process by using fault tolerant techniques. Computer Systems Science and Engineering, 9(2):83–88, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Khonji:2013:EEF

[KJI13] Mahmoud Khonji, Andrew Jones, and Youssef Iraqi. An empirical evaluation for feature selection methods in phishing email classification. International Journal of Computer Systems Science and Engineering, 28(1):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Kim:2007:TTM

[KJKK07] M. H. Kim, E. H. Jo, D-J Kim, and J-G Kim. Time-triggered and messagetriggered object framework and global time-based synchronization for real-time multimedia streaming. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Khan:1996:PCI

[KK96] Zahira S. Khan and Eugene Kwatny. Performance comparison of interprocessor communication schemes for a hashing technique on the Connection Machine. International Journal of Computer Systems Science and Engineering, 11(4):189–194, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Kayi:2009:PAT

[KKEG^{+}09] Abdullah Kayi, Edward Kornkven, Tarek El-Ghazawi, Samy Al-Bazhra, and Gregory B. Newby. Performance analysis and tuning for clusters with ccNUMA nodes for scientific computing: a case study. International Journal of Computer Systems Science and Engineering, 24(5): ??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Kondo:2012:ITR

[KKHN12] Shinya Kondo, Akimitsu Kanzaki, Takahiro Hara, and Shojiro Nishio. Integration of traffic reduction and sleep scheduling in wireless sensor networks. International Journal of Computer Systems Science and Engineering, 27(3): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Kondo:2013:ITR

[KKHN13] Shinya Kondo, Akimitsu Kanzaki, Takahiro Hara, and Shojiro Nishio. Integration of traffic reduction and sleep scheduling for energy-efficient data gathering in wireless sensor networks. International Journal of Computer Systems Science and Engineering, 28 (4):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Kim:2002:SRT

[KKK^{+}02] J-H Kim, S-W Kim, D-Y Kim, Y-Y You, and W. Choi. Soft real-time scheduling in a general purpose operating system: scheduling daemon approach. International Journal of Computer Systems Science and Engineering, 17(1): ??, January 2002. CODEN CSSEEI. ISSN 0267-6192.

Kim:2002:TBA

[KKLK02] D-H Kim, K. H. (Kane) Kim, S. Liu, and J-H Kim. A TMObased approach to tolerance of transmission jitters in teleaudio services. International Journal of Computer Systems Science and Engineering, 17 (6):??, November 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Kadlec:2011:CTP

[KKP11] Albrecht Kadlec, Raimund Kirner, and Peter Puschner. Code transformations to prevent timing anomalies. International Journal of Computer Systems Science and Engineering, 26(6):??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Kang:2004:FFT

[KKY04] Sooyong Kang, Hyunjoo Kim, and Heon Y. Yeom. A framework for fault-tolerant peer-to-peer streaming service based on the service migration. International Journal of Computer Systems Science and Engineering, 19(4): ??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Katsaros:2000:TDQ

[KL00] T. P. Katsaros and C. Lazos. A technique for determining queuing network simulation length based on desired accuracy. International Journal of Computer Systems Science and Engineering, 15(6): ??, December 2000. CODEN CSSEEI. ISSN 0267-6192.

Kim:2001:DHR

[KLK01] K. H. (Kane) Kim, J. Liu, and M-H Kim. Deadline handling in real-time distributed objects: issues and basic approaches. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Kim:2002:CSM

[KLMS02] K. H. (Kane) Kim, J. Q. Liu, H. Miyazaki, and E. H. Shokri. A CORBA service middleware enabling highlevel high-precision real-time distributed object programming. International Journal of Computer Systems Science and Engineering, 17(2): ??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Koodziej:2012:P

[KN12] Joanna Koodziej and Hiroaki Nishino. Preface. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Kamel:1992:MMC

[KNEDK92] Ahmed Kamel, Ahmed Nazif, Ossama El-Dessouki, and Nabil Kamel. MCFS: a multiple criteria reasoning fuzzy expert systems building tool. Computer Systems Science and Engineering, 7(3):202– 208, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Kamal:2016:PEC

[KNP16] Md. Sarwar Kamal, Sonia Farhana Nimmy, and Sazia Parvin. Performance evaluation comparison for detecting dna structural break through big data analysis. International Journal of Computer Systems Science and Engineering, 31(4):??, July 2016. CODEN CSSEEI. ISSN 0267-6192.

Kagawa:2012:DEI

[KNU12] Tsuneo Kagawa, Hiroaki Nishino, and Kouichi Utsumiya. Development and evaluation of an interactive texture design method. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Koch:1994:EPS

[Koc94] Henning Koch. Embedding protocols for scalable replication management. Computer Systems Science and Engineering, 9(2):98–103, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Kof:2005:ANL

[Kof05] Leonid Kof. An application of natural language processing to domain modelling: two case studies. International Journal of Computer Systems Science and Engineering, 20 (1):??, January 2005. CO-DEN CSSEEI. ISSN 0267-6192.

Kouka:1990:ARW

[KP90] Edmond F. Kouka and Jean Luc Patry. Algorithms for restructuring WSI arrays of processors. Computer Systems Science and Engineering, 5(2):95–104, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Kim:1992:ETR

[KR92] J. H. Kim and S. M. Reddy. Easily testable and reconfigurable two-dimensional systolic arrays. Computer Systems Science and Engineering, 7(3):160–169, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Kuchlos:1995:TWP

[KR95] A. Kuchlos and C. P. Ravikumar. Two-way partitioning of shuffle-exchange and De-Bruijn graphs. International Journal of Computer Systems Science and Engineering, 10 (2):84–92, April 1995. CO-DEN CSSEEI. ISSN 0267-6192.

Krasniewski:1997:CDF

[Kra97] Andrzej Krasniewski. Coverage of delay faults: single number can be very misleading. International Journal of Computer Systems Science and Engineering, 12(1): 3–12, January 1997. CODEN CSSEEI. ISSN 0267-6192.

Kristensen:2001:SB

[Kri01] B. B. Kristensen. Subjective behaviour. International Journal of Computer Systems Science and Engineering, 16 (1):??, January 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Kristensen:2008:ICG

[Kri08] Bent Bruun Kristensen. Interaction cases: generalized use cases in ambient systems. International Journal of Computer Systems Science and Engineering, 23(4):??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Kresoja:2016:PNE

[KRSS16] Sasa Kresoja, Milos Rackovic, Srdjan Skrbic, and Bojana Dimic Surla. A Petri net extension for formal modelling of information systems. International Journal of Computer Systems Science and Engineering, 31(3):??, May 2016. CODEN CSSEEI. ISSN 0267-6192.

Kant:1997:ADC

[KS97] Latha Kant and William H.

Sanders. Analysis of the distribution of consecutive cell losses in an ATM switch using stochastic activity networks. International Journal of Computer Systems Science and Engineering, 12(2):117– 129, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Kim:1999:APR

[KS99] K. H. (Kane) Kim and Chittur Subbaraman. Architectural principles for realizing timeliness-guaranteed operating system kernels. International Journal of Computer Systems Science and Engineering, 14(4):241–249, July 1999. CODEN CSSEEI. ISSN 0267-6192.

Karaata:2000:DSS

[KS00] M. H. Karaata and K. A. Saleh. A distributed selfstabilizing algorithm for finding maximum matching. International Journal of Computer Systems Science and Engineering, 15(3):175–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Kohlhoff:2004:ESH

[KS04] Christopher Kohlhoff and Robert Steele. Evaluating SOAP for high performance applications in capital markets. International Journal of Computer Systems Science and Engineering, 19(4):??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Kim:2008:SMU

[KS08] Sang-Wook Kim and Miyoung Shin. Subsequence matching under time warping in time-series databases: observation, optimization, and performance results. International Journal of Computer Systems Science and Engineering, 23(1):??, January 2008. CODEN CSSEEI. ISSN 0267-6192.

Krunz:1996:SVM

[KSH96] Marwan Krunz, Ron Sass, and Herman Hughes. Study of VBR MPEG-coded video traffic and associated multiplexing performance. International Journal of Computer Systems Science and Engineering, 11(3):135–143, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Kurki-Suonio:1999:LLS

[KSK99] Reino Kurki-Suonio and Mika Katara. Logical layers in specifications with distributed objects and real time. International Journal of Computer Systems Science and Engineering, 14(4):217–226, July 1999. CODEN CSSEEI. ISSN 0267-6192.

Kim:2012:PRF

[KSL12] Hyunwoo Kim, Mye Sohn, and Hyun Jung Lee. Personalized recommendation framework based on CBR and CSP using ontology in a ubiquitous computing environment. International Journal of Computer Systems Science and Engineering, 27(6):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Kadamuddi:1995:MCI

[KT95] Dinesh Kadamuddi and Jeffrey J. P. Tsai. MDDT constructs for incorporating early-receives in distributed computing. International Journal of Computer Systems Science and Engineering, 10 (4):223–233, October 1995. CODEN CSSEEI. ISSN 0267-

Kalpana:2014:FLB

[KTJ14] A. M. Kalpana, K. Tamizarasu, and A. Ebenezer Jeyakumar. A fuzzy logic based framework for assessing the maturity level of Indian small scale software organizations. International Journal of Computer Systems Science and Engineering, 29(2):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Kang:2014:DEN

[KTK14] Ho-Seok Kang, Tran Tin, and Sung-Ryul Kim. Design and experiments of new IP traceback method based on offline analysis. International Journal of Computer Systems Science and Engineering, 29(6): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Khatri:2014:TEW

[KTV14] Pallavi Khatri, S. Tapaswi, and U. P. Verma. Trust evaluation in wireless ad-hoc networks using fuzzy system. International Journal of Computer Systems Science and Engineering, 29(1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Kovacs:2008:FAB

[KVG08] Máté Kovács, Dániel Varró, and László Gönczy. Formal analysis of BPEL workflows with compensation by model checking. International Journal of Computer Systems Science and Engineering, 23(5):

??, September 2008. CODEN CSSEEI. ISSN 0267-6192.

Komathy:2003:SSB

[KVR03] K. Komathy, P. Vivekanandan, and V. Ramachandran. Secure SOAP-based Web services. International Journal of Computer Systems Science and Engineering, 18(1): ??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Komathy:2004:SSB

[KVR04] K. Komathy, P. Vivekanandan, and V. Ramachandran. Secure SOAP-based Web services. International Journal of Computer Systems Science and Engineering, 19(1): ??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Kim:2001:NJP

[KW01] S.-W. Kim and K.-Y. Whang.

New join processing using a multidimensional dynamic file organization. International Journal of Computer Systems Science and Engineering, 16 (1):??, January 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Kim:2002:BLM

[KW02] S-W Kim and H-S Won. Bulkloading of a multidimensional file: algorithm and its performance evaluation. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/ September 2002. CODEN CSSEEI. ISSN 0267-6192.

Kung:2010:TTB

[KW10] H. T. Kung and S. Y. Wang. TCP trunking for bandwidth management of aggregate traffic. International Journal of Computer Systems Science and Engineering, 25(1): ??, January 2010. CODEN CSSEEI. ISSN 0267-6192.

Kim:2007:GAT

[KWL07] Min-Soo Kim, Kyu-Young Whang, and Jae-Gil Lee. n-Gram/2L-approximation: a two-level n-gram inverted index structure for approximate string matching. International Journal of Computer Systems Science and Engineering, 22(6):??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Lami:2009:AEE

[Lam09] Giuseppe Lami. Analytic effectiveness evaluation of techniques for natural language software requirements testability. International Journal of Computer Systems Science and Engineering, 24(2): ??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Lazcano:2000:WAE

[LASS00] A. Lazcano, G. Alonso, H. Schuldt, and C. Schuler. The WISE approach to electronic commerce. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Latifi:1996:CSG

[LB96] Shahram Latifi and Nader Bagherzadeh. On clusteredstar graph and its properties. International Journal of Computer Systems Science and Engineering, 11(3):145– 149, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Lee:2003:AXT

[LB03] B.-S. Lee and B. R. Bryant.

Applying XML technology for implementation of natural language specifications. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Lamanna:2013:HAB

[LBLB13] D. Davide Lamanna, Flavio Bertini, Giorgia Lodi, and Roberto Baldoni. How to act without being observed: Progressive privacy architecture in desktop-as-a-service. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Lin:1996:CBC

[LC96a] Jiann-Fu Lin and Sao-Jie Chen. Comment on ‘Broadcasting cryptosystem in computer networks using interpolating polynomials’. International Journal of Computer Systems Science and Engineering, 11(5):315–317, September 1996. CODEN CSSEEI. ISSN 0267-6192. See [CW91, HLLC96].

Linn:1996:AMT

[LC96b] Jiann-Fu Linn and Sao-Jie Chen. Analysis of multiprocessor task scheduling. International Journal of Computer Systems Science and Engineering, 11(2):117–120, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Lou:1998:FMM

[LC98] Der-Chyuan Lou and Chin-Chen Chang. A fast modular multiplication method. International Journal of Computer Systems Science and Engineering, 13(6):353–358, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Lee:2000:UIK

[LC00a] W.-B. Lee and C.-C. Chang. User identification and key distribution maintaining anonymity for distributed computer networks. International Journal of Computer Systems Science and Engineering, 15(4): 211–??, July 2000. CODEN CSSEEI. ISSN 0267-6192.

Lou:2000:EDC

[LC00b] D.-C. Lou and C.-C. Chen. An efficient divide-and-conquer technique for parallel computation of modular multiexponentiation. International Journal of Computer Systems Science and Engineering, 15 (2):111–??, March 2000. CO-DEN CSSEEI. ISSN 0267-6192.

Liu:1992:PAL

[LCH92] Jenshiuh Liu, Chi-Ming Chiang, and Herman D. Hughes. Performance analysis of loadsharing for multiprocessor systems. Computer Systems Science and Engineering, 7 (4):210–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Liu:2012:PNM

[LCWC12] Tingting Liu, Jie Cui, Bo Wang, and Gang Chen. Parallelization of normal mode-WKBZ computation for ocean acoustic field. International Journal of Computer Systems Science and Engineering, 27(5): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Lai:2014:DMM

[LCWC14] Wei Kuang Lai, Yi-Uan Chen, Tin-Yu Wu, and Ya-Yin Chen. Dual migration for mobile cloud service in 4G/ LTE network. International Journal of Computer Systems Science and Engineering, 29 (6):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Liu:2011:MCL

[LCWY11] Likun Liu, Cheng Chen, Yongwei Wu, and Guangwen Yang. Metadata changes in large file systems: a metadata querying perspective. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Li:2008:CAM

[LCX08] Zhongwen Li, Qiong Chen, and Yang Xiang. A crossauthentication model and implementation. International Journal of Computer Systems Science and Engineering, 23 (3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Liu:1996:PSA

[LD96] W. Liu and Erik Dirkx. Parallel simulation of ATM switches. International Journal of Computer Systems Science and Engineering, 11 (6):369–381, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Lew:1986:MEP

[LDF86] K. S. Lew, T. S. Dillon, and K. E. Forward. Model of error propagation in database systems. Computer Systems Science and Engineering, 1(2): 114–118, January 1986. CO-DEN CSSEEI. ISSN 0267-6192.

Lotz:2013:CMT

[LdOS13] Volkmar Lotz, Anderson Santana de Oliveira, and Jakub Sendor. Control as a means towards accountable services in the cloud. International Journal of Computer Systems Science and Engineering, 28 (6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Li:2008:VLS

[LDZP08] Zhuowei Li, Amitabha Das, Jianying Zhou, and Jagdish C. Patra. Variable-length signatures for intrusion detection. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Lee:1985:IQC

[Lee85]  Barbara K. Lee.  Imple-

menting a quality circle programme for computer professionals. Computer Systems Science and Engineering, 1 (1):65–67, October 1985. CO-DEN CSSEEI. ISSN 0267-6192.

Lee:2004:MCH

[Lee04] Yong-Jin Lee. Minimal cost heuristic algorithm for delay constrained loop network. International Journal of Computer Systems Science and Engineering, 19(4):??, July 2004. CODEN CSSEEI. ISSN 0267-6192.

Lee:2009:LPT

[Lee09] Jung-Hoon Lee. Low power TLB structure by using dynamic searching algorithm. International Journal of Computer Systems Science and Engineering, 24(4):??, July 2009. CODEN CSSEEI. ISSN 0267-6192.

Lin:1997:PEP

[LH97] Hui-Tang Lin and Herman D. Hughes. Performance evaluation of priority buffer management in ATM switches. International Journal of Computer Systems Science and Engineering, 12(6):387–393, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Leung:1998:PAN

[LH98] Ka-Cheong Leung and Mounir Hamdi. Performance assessment of network protocols and parallel programming tools for distributed computing systems. International Journal of Computer Systems Science and Engineering, 13(1):67– 80, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Lin:2002:HSW

[LH02] H-T Lin and H. D. Hughes. A hand-off scheme for wireless ATM networks. International Journal of Computer Systems Science and Engineering, 17 (3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Li:2008:ISS

[LHC08] Chua-Ta Li, Min-Shiang Hwang, and Yen-Ping Chu. Improving the security of a secure anonymous routing protocol with authenticated key exchange for ad hoc networks. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Lee:2004:TBS

[LHW04] Min-Jae Lee, Wook-Shih Han, and Kyu-Young Whang. Transformation-based spatial partition join. International Journal of Computer Systems Science and Engineering, 19 (6):??, November 2004. CO-DEN CSSEEI. ISSN 0267-6192.

Li:1998:ERP

[Li98] Keqin Li. Efficient reliability prediction for N-version software systems with multiple stages. International Journal of Computer Systems Science and Engineering, 13(2):113– 120, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Li:1999:AMA

[Li99] Keqin Li. An analytical model for approximating performance of partitionable multiprocessor systems. International Journal of Computer Systems Science and Engineering, 14(3):131–145, May 1999. CODEN CSSEEI. ISSN 0267-6192.

Lai:1996:VEP

[LJ96] R. Lai and A. Jirachiefpattana. Verifying Estelle protocol specifications using numerical Petri nets. International Journal of Computer Systems Science and Engineering, 11(1):15–33, January 1996. CODEN CSSEEI. ISSN 0267-6192.

Loh:1997:DLB

[LJ97] Peter K. K. Loh and Hsu Wen Jing. Dynamic load balancing on multiprocessor networks. International Journal of Computer Systems Science and Engineering, 12(6):369– 372, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Leu:2012:RAB

[LJL^{+}12] Fang-Yie Leu, Fuu-Cheng Jiang, Chih-Cheng Lien, Sen-Tarng Lai, and Shiuan-You Chiou. A rate-allocation based multi-path control scheme for event-driven wireless sensor networks on constant event packet rates. International Journal of Computer Systems Science and Engineering, 27(5):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Lei:2015:RLO

[LJW^{+}15] Jingsheng Lei, Teng Jiang, Kui Wu, Haizhou Du, and Lin Zhu. Robust local outlier detection with statistical parameter for big data. International Journal of Computer Systems Science and Engineering, 30(5):??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Li:2016:SIB

[LK16] Guest Editors: Zhiyang Li and Nei Kato. Special issue: BDCloud 2015 and FCST 2015. International Journal of Computer Systems Science and Engineering, 31(2): ??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Lee:2014:ABE

[LKK14] Yong-Hwan Lee, Bonam Kim, and Youngseop Kim. AAMbased emotion recognition using variance of facial feature points on mobile video stream. International Journal of Computer Systems Science and Engineering, 29(6): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Loganantharaj:1994:RRP

[LL94] Rasiah Loganantharaj and Ay-Hwa A. Liou. Recent results on parallel link resolution. International Journal of Computer Systems Science and Engineering, 9(4):220– 233, October 1994. CODEN CSSEEI. ISSN 0267-6192.

Li:1996:CAS

[LL96] Shuo-Yen Robert Li and Chu Man Lau. Concentrators in ATM switching. International Journal of Computer Systems Science and Engineering, 11(6):335–342, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Lin:2010:SIS

[LL10] Kai Lin and Kequi Li. Special issue: Selected topics on wireless sensor networks. International Journal of Computer Systems Science and Engineering, 25(6):??, November 2010. CODEN CSSEEI. ISSN 0267-6192.

Liu:2015:IUS

[LL15] Weijiang Liu and Zhiyang Li. ICA3PP/U-Science 2014. International Journal of Computer Systems Science and Engineering, 30(6): ??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Ling:2016:WRP

[LL16] Qiu Ling and Qi Linkai. water-related public organizations. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Lu:2006:RSC

[LLHG06] Z. Lu, S. Li, P. Hyland, and A. Ghose. Reliable service composition by adopting assumptions. International Journal of Computer Systems Science and Engineering, 21 (4):??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Landis:1991:OMR

[LM91] David L. Landis and Daniel C. Muha. Optimal maintenance resource allocation for multiprocessor systems. Computer Systems Science and Engineering, 6(1):54–??, January 1991. CODEN CSSEEI. ISSN 0267-6192.

Ligon:1992:PPT

[LM92] Walter B. Ligon III and Aditya P. Mathur. Parallel parsing on a transputer network. Computer Systems Science and Engineering, 7(3): 152–159, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Li:2009:SIA

[LM09] Keqin Li and Geyong Min.

Special issue: Advanced topics on grid, cluster and P2P computing. International Journal of Computer Systems Science and Engineering, 24 (5):??, September 2009. CO-DEN CSSEEI. ISSN 0267-6192.

Lam:1990:PAC

[LMI90] Y. M. Lam, H. T. Mouftah, and M. Ilyas. Performance analysis of CSMA-CD protocols in local area bus networks. Computer Systems Science and Engineering, 5 (2):105–110, April 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Lee:2012:SSPb

[LMPJ12] Jongsuk R. Lee, Don Mc-Nickle, Krzysztof Pawlikowski, and Hae-Duck J. Jeong. Selfsimilar properties of malicious teletraffic. International Journal of Computer Systems Science and Engineering, 27 (6):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Lopez:2004:PPP

[LNR04] Natalia López, Manuel Núñez, and Fernando Rubio. Predicting performance in the presence of stochastic information. International Journal of Computer Systems Science and Engineering, 19(6): ??, November 2004. CODEN CSSEEI. ISSN 0267-6192.

Li:2007:ECR

[LNS^{+}07] K. Li, T. Nanya, H. Shen, F. Chin, and W. Zhang. An efficient cache replacement algorithm for multimedia object caching. International Journal of Computer Systems Science and Engineering, 22(1– 2):??, January/March 2007. CODEN CSSEEI. ISSN 0267-6192.

Lovric:1996:DHF

[Lov96] Tomislav Lovric. Detecting hardware faults with systematic and design diversity: Experimental results. International Journal of Computer Systems Science and Engineering, 11(2):83–92, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Liu:2001:PEM

[LOZ01] C. Liu, M. A. Orgun, and K. Zhang. A parallel execution model for Chronolog. International Journal of Computer Systems Science and Engineering, 16(4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Lauwereins:1993:QTA

[LP93] Rudy Lauwereins and Jean Peperstraete. Queueing theoretical analysis of processor utilization in parallel computers. Computer Systems Science and Engineering, 8(1): 13–23, January 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Liu:2003:ISA

[LP03] D.-R. Liu and S.-C. Pan. Integrating software agents with multi-level metadata in support of heterogeneous product search. International Journal of Computer Systems Science and Engineering, 18(3): ??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Liu:1999:CCQ

[LPTH99] Ling Liu, Calton Pu, Wei Tang, and Wei Han. CON-QUER: a continual query system for update monitoring in the WWW. International Journal of Computer Systems Science and Engineering, 14 (2):99–112, March 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Lee:2003:PES

[LPY03] K. Lee, Y. Y. Park, and H. Y. Yeom. Pre-emptive but safe interval caching for realtime multimedia systems. International Journal of Computer Systems Science and Engineering, 18(2):??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Li:2014:P

[LQ14a] Kequi Li and Heng Qi. Preface. International Journal of Computer Systems Science and Engineering, 29(4): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Li:2014:SIM

[LQ14b] Kequi Li and Heng Qi. Special issue: Multimedia information processing and retrieval. International Journal of Computer Systems Science and Engineering, 29(4): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Liu:2011:NME

[LQLL11] Weijiang Liu, Wenyu Qu, Zhaobin Liu, and Kequi Li. A novel method for estimating flow length distributions from double-sampled flow statistics. International Journal of Computer Systems Science and Engineering, 26(5): ??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Li:2014:ESR

[LQX^{+}14] Zhiyang Li, Wenyu Qu, Yujie Xu, Junjie Cao, and Zhixun Su. Efficient shape representation and retrieval in large database. International Journal of Computer Systems Science and Engineering, 29(4): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Lenders:1994:MDF

[LS94] Patrick M. Lenders and Heiko Schroder. Minimizing data flow in mathematical morphology architectures. International Journal of Computer Systems Science and Engineering, 9(4):240–245, October 1994. CODEN CSSEEI. ISSN 0267-6192.

Lyuu:1997:NAM

[LS97] Yuh-Dauh Lyuu and Eugen Schenfeld. New algorithms for matrix operations with applications to a reconfigurable parallel architecture. International Journal of Computer Systems Science and Engineering, 12(1):13–24, January 1997. CODEN CSSEEI. ISSN 0267-6192.

Li:2004:PPP

[LS04] Keqiu Li and Hong Shen.

Proxy placement problem for coordinated en-route transcoding proxy caching. International Journal of Computer Systems Science and Engineering, 19(6):??, November 2004. CODEN CSSEEI. ISSN 0267-6192.

Li:2014:DTO

[LSBW14] Wuyungerile Li, Shunsuke Saruwatari, Masaki Bandai, and Takashi Watanabe. Discussions on trade-offs in data aggregation in wireless sensor networks. International Journal of Computer Systems Science and Engineering, 29 (1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Lilien:1995:RHS

[LSD95] L. Lilien, S. M. Shatz, and S. Damerla. Redistribution of hierarchically structured software in response to distributed system site crashes. International Journal of Computer Systems Science and Engineering, 10(3): 144–155, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Li:2009:SIN

[LSG09] Kequi Li, Yanming Shen, and Minyi Guo. Special issue: Network and parallel computing. International Journal of Computer Systems Science and Engineering, 24(3): ??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Lor:1996:DCM

[LSM96] Sam Lor, Hong Shen, and Piyush Maheshwari. Divideand-conquer minimal-cut bisectioning of task graphs. International Journal of Computer Systems Science and Engineering, 11(4):227–234, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Lam:1992:GAG

[LSR^{+}92] H. Lam, S. Y. W. Su, V. Ruhela, S. Pant, S. M. Ju, M. Sharma, and N. Prasad. GTOOLS: an active GUI toolset for an object-oriented KBMS. Computer Systems Science and Engineering, 7 (2):69–85, April 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Lee:1997:NBS

[LT97] Tony T. Lee and Philip P.

To. Non-blocking and selfrouting properties of sort-Clos networks. International Journal of Computer Systems Science and Engineering, 12(2): 159–169, March 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Layfield:2009:DIS

[LTKK09] Ryan Layfield, Bhavani Thuraisingham, Laatifur Khan, and Murat Kantarcioglu. Design and implementation of a secure social network system. International Journal of Computer Systems Science and Engineering, 24(2): ??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Li:2008:RDR

[LTLH08] Ruixuan Li, Zhuo Tang, Zhengding Lu, and Jinwei Hu. Request-driven role mapping framework for secure interoperation in multi-domain environments. International Journal of Computer Systems Science and Engineering, 23 (3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Laranjeiro:2008:DFT

[LV08] Nuno Laranjeiro and Marco Vieira. Deploying fault tolerant Web service compositions. International Journal of Computer Systems Science and Engineering, 23(5): ??, September 2008. CODEN CSSEEI. ISSN 0267-6192.

Lyuu:2010:GUS

[LW10] Yuh-Dauh Lyuu and Ming-Luen Wu. Group undeniable signatures with convertibility. International Journal of Computer Systems Science and Engineering, 25(5): ??, September 2010. CODEN CSSEEI. ISSN 0267-6192.

Lin:2004:HOT

[LWD04] W. W. K. Lin, A. K. Y. Wong, and T. S. Dillon. HBP: an optimization technique to shorten the control cycle time of the Neural Network Controller that provides dynamic buffer tuning to eliminate overflow at user level. International Journal of Computer Systems Science and Engineering, 19(2):??, March 2004. CODEN CSSEEI. ISSN 0267-6192.

Lee:2006:SIH

[LWHS06] Jae-Gil Lee, Kyu-Young Whang, Wook-Shin Han, and Il-Yeol Song. Special issue: Hippocratic XML databases: a model and an access control mechanism. International Journal of Computer Systems Science and Engineering, 21 (6):??, November 2006. CO-DEN CSSEEI. ISSN 0267-6192.

Liaw:1993:SDU

[LWL93] H. T. Liaw, S. J. Wang, and C. L. Lei. Secure dynamic user hierarchy scheme based on the knapsack problem. Computer Systems Science and Engineering, 8(4): 210–216, October 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Lo:2007:SNT

[LWL07] Jason C. T. Lo, Allan K. Y.

Wong, and Wilfred W. K. Lin. SDITPM: A novel transfer policy model to facilitate object mobility that shortens service roundtrip time by load balancing over the Internet. International Journal of Computer Systems Science and Engineering, 22(1–2):??, January/March 2007. CO-DEN CSSEEI. ISSN 0267-6192.

Lee:1992:AAP

[LY92] Yann-Hang Lee and Philip S.

Yu. Adaptive access path selection for relational database systems. Computer Systems Science and Engineering, 7 (1):52–61, January 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Lien:1995:SNC

[LY95] Hsiou-Mien Lien and Shyan-Ming Yuan. SGHC: a new class of optimally faulttolerant networks. International Journal of Computer Systems Science and Engineering, 10(1):57–64, January 1995. CODEN CSSEEI. ISSN 0267-6192.

Liang:1996:DPA

[LY96] Shih T. Liang and Maria C. Yuang. Departure process analysis for earliest-due-date scheduling discipline in ATM switches. International Journal of Computer Systems Science and Engineering, 11 (6):343–352, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Lee:2012:IPU

[LY12] Jae-Suk Lee and Sung-Bong Yang. An improved pathfinding under multiple exits with SOA in a double-layered MANET. International Journal of Computer Systems Science and Engineering, 27(2): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Li:2006:SBA

[LYA06] N. Li, T. Yu, and A. Antón. A semantics based approach to privacy languages. International Journal of Computer Systems Science and Engineering, 21(5):??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

Lou:2002:ESA

[LYC02] D-C Lou, T-L Yin, and M-C Chang. An efficient steganographic approach. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/ September 2002. CODEN CSSEEI. ISSN 0267-6192.

Lin:1995:DIM

[LYL95] Yen-Wen Lin, Shyan-Ming Yuan, and Deron Liang. Design and implementation of Moony: a fault tolerant distributed shared memory system. International Journal of Computer Systems Science and Engineering, 10(2):111– 119, April 1995. CODEN CSSEEI. ISSN 0267-6192.

Lee:2012:SSPa

[LYL12a] Sun-Ho Lee, Kangbin Yim, and In-Yeong Lee. A secure solution with a password recovery feature for USB flash drives. International Journal of Computer Systems Science and Engineering, 27(6): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Li:2012:UBR

[LYL^{+}12b] Yuhoing Li, Liang Yu, Tao Li, Yishan Li, and Yan Shi. Utility-based resource allocation in heterogeneous wireless environments using game theory. International Journal of Computer Systems Science and Engineering, 27(5): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Leff:1993:PIO

[LYW93] Avraham Leff, Philip S. Yu, and Joel L. Wolf. Performance issues in object replication for a remote caching architecture. Computer Systems Science and Engineering, 8(1):40–51, January 1993. CODEN CSSEEI. ISSN 0267-6192.

Mouftah:1997:SIA

[MA97] Hussein Mouftah and Mohammed Atiquzzaman. Special issue: ATM networks. International Journal of Computer Systems Science and Engineering, 12(2):??, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Maeda:2010:SSP

[Mae10] Hiroshi Maeda. Simulation of soliton propagation in slab waveguide by frequency dependent FDTD method. International Journal of Computer Systems Science and Engineering, 25(2):??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Miranda:2000:PRS

[MÁGA00] J. Miranda, A. Álvarez, F. Guerra, and S. Arévalo. Programming replicated systems in Drago. International Journal of Computer Systems Science and Engineering, 15 (1):49–??, January 2000. CO-DEN CSSEEI. ISSN 0267-6192.

Molnar:1988:SOF

[MB88] R. K. Molnar and S. C. Bruell. Sensitivity of operational formulae. Computer Systems Science and Engineering, 3 (2):51–66, April 1988. CO-DEN CSSEEI. ISSN 0267-6192.

Mittal:1995:SRD

[MC95] Ravi Mittal and Deepak Cherian. Self-routing double tree (DOT) network, its permutation capability and applications. International Journal of Computer Systems Science and Engineering, 10 (1):19–27, January 1995. CO-DEN CSSEEI. ISSN 0267-6192.

McAvoy:2015:OCR

[MCD^{+}15] Laura M. McAvoy, Liming Chen, Mark P. Donnelly, Chris D. Nugent, and Paul J. McCullagh. Ontological characterization and representation of context within smart environments. International Journal of Computer Systems Science and Engineering, 30 (1):??, January 2015. CO-DEN CSSEEI. ISSN 0267-6192.

Moh:1995:PCR

[MCMM95] W. Melody Moh, Yu-Jen Chien, Teng-Sheng Moh, and Charles U. Martel. Prioritized conflict resolution on multiple access broadcast networks: algorithms and performance. International Journal of Computer Systems Science and Engineering, 10(4): 234–243, October 1995. CO-DEN CSSEEI. ISSN 0267-

Militello:2010:EIR

[MCSV10] C. Militello, V. Conti, F. Sorbello, and S. Vitabile. An embedded IRIS recognizer for portable and mobile devices. International Journal of Computer Systems Science and Engineering, 25(2): ??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Militello:2014:FFT

[MCSV14] Carmelo Militello, Vincenzo Conti, Filippo Sorbello, and Salvatore Vitabile. A fast fusion technique for fingerprint and iris spatial descriptors in multimodal biometric systems. International Journal of Computer Systems Science and Engineering, 29(3): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Moga:2003:SIS

[MD03] A. Moga and M. Dubois. Scalability implications of software-implemented coherence. International Journal of Computer Systems Science and Engineering, 18(1): ??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Moga:2004:SIS

[MD04] A. Moga and M. Dubois. Scalability implications of software-implemented coherence. International Journal of Computer Systems Science and Engineering, 19(1): ??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Martinez:2003:XMS

[MdlFD03] M. M. Martínez, P. de la Fuente, and J.-C. Derniame. XML as a means to support information extraction from legal documents. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Megson:1990:CES

[Meg90] G. M. Megson. Complex eigenproblem solution by a parallel norm-reducing Jacobi method. Computer Systems Science and Engineering, 5 (3):159–168, July 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Meng:1999:DSI

[Men99] Xiannong Meng. Distributed simulations: Issues and implementations in a cluster of workstations environment. International Journal of Computer Systems Science and Engineering, 14(1):27– 37, January 1999. CODEN CSSEEI. ISSN 0267-6192.

Metin:2016:NUM

[Met16] Senem Kumova Metin. Neighbour unpredictability measure in multiword expression extraction. International Journal of Computer Systems Science and Engineering, 31 (3):??, May 2016. CODEN CSSEEI. ISSN 0267-6192.

Mbale:2002:AIU

[MFU02] J. Mbale, X. X. Fei, and D. Ursino. Acquisition of interoperability using intelligent binary schema matching technology. International Journal of Computer Systems Science and Engineering, 17 (6):??, November 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Mohsin:1995:FTP

[MG95] M. Mohsin and B. Gupta. Fault-tolerance in pyramid tree network architecture. International Journal of Computer Systems Science and Engineering, 10(3):164–172, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Ma:2002:SCM

[MG02] M. Ma and E. Gunawan. Single-code and multi-code DS/CDMA MAC protocols for video/voice integration. International Journal of Computer Systems Science and Engineering, 17(6): ??, November 2002. CODEN CSSEEI. ISSN 0267-6192.

Munoz:2015:SCU

[MGOB15] Diego Muñoz, Francisco J. Gutierrez, Sergio F. Ochoa, and Nelson Baloian. Social connector: a ubiquitous system to ease the social interaction among family community members. International Journal of Computer Systems Science and Engineering, 30(1):

??, January 2015. CODEN CSSEEI. ISSN 0267-6192.

Morovan:2010:SIM

[MH10] Frank Morovan and Abdelkader Hameurlain. Special issue: Mobile data management: Models, methodologies and services. International Journal of Computer Systems Science and Engineering, 25 (3):??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Miller:1993:MSD

[MHH93] P. M. Miller, A. R. Hurson, and R. H. Hettmansperger. Modular scheme for designing associative memories. Computer Systems Science and Engineering, 8(3):166–181, July 1993. CODEN CSSEEI. ISSN 0267-6192.

Miller:1992:CDP

[MHHP92] L. L. Miller, A. R. Hurson, J. K. Huang, and S. H. Pakzad. Comparing database processor configurations in the database machine ASLM. Computer Systems Science and Engineering, 7(1):17– 24, January 1992. CODEN CSSEEI. ISSN 0267-6192.

Majzik:1996:MCU

[MHPS96] I. Majzik, W. Hohl, A. Pataricza, and V. Sieh. Multiprocessor checking using watchdog processors. International Journal of Computer Systems Science and Engineering, 11 (5):301–310, September 1996.

CODEN CSSEEI. ISSN 0267-6192.

Mejia:2014:IAI

[MJRIV14] Andrés Mejía, Reyes Juárez-Ramírez, Sergio Inzunza, and Rocio Valenzuela. Implementing adaptive interfaces: a use model for the development of usability in interactive systems. International Journal of Computer Systems Science and Engineering, 29 (1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Mahajan:1990:EPI

[MK90] Milind Mahajan and V. K. Prasanna Kumar. Efficient parallel implementation of RETE pattern matching. Computer Systems Science and Engineering, 5(3):187–192, July 1990. CODEN CSSEEI. ISSN 0267-6192.

Moon:2010:MSP

[MK10] Yang-Sae Moon and Jinho Kim. Mbr-safe piecewise aggregate approximation for time-series subsequence matching. International Journal of Computer Systems Science and Engineering, 25(5): ??, September 2010. CODEN CSSEEI. ISSN 0267-6192.

Mahmood:2011:IGA

[MK11] Ali Mirza Mahmood and Mrithyumjaya Rao Kuppa. Increasing generalization accuracy by using multivariate statistical method. International Journal of Computer Systems Science and Engineering, 26(2):??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Mai:2015:IOR

[MKK15] Hai Tanh Mai, Jaeho Kim, and Myoung Ho Kim. Integrating outlier removal into existing histogram construction methods for geographic data. International Journal of Computer Systems Science and Engineering, 30(2): ??, March 2015. CODEN CSSEEI. ISSN 0267-6192.

Min:1986:PET

[ML86] Yinghua Min and Zhongcheng Li. Pseudo-exhaustive testing strategy for large combinational circuits. Computer Systems Science and Engineering, 1(4):213–220, October 1986. CODEN CSSEEI. ISSN 0267-6192.

Moon:2009:ESSa

[ML09a] Yang-Sae Moon and Woong-Kee Loh. Efficient stream subsequence matching algorithms for handheld devices on streaming time-series data. International Journal of Computer Systems Science and Engineering, 24(2): ??, March 2009. CODEN CSSEEI. ISSN 0267-6192.

Moon:2009:ESSb

[ML09b] Yang-Sae Moon and Woong-Kee Loh. Efficient stream subsequence matching algorithms for handheld devices on streaming time-series data. International Journal of Computer Systems Science and Engineering, 24(4):??, July 2009. CODEN CSSEEI. ISSN 0267-6192.

Miller:1994:SMP

[MLGC94] L. L. Miller, J. Leszczylowski, S. K. Gadia, and P. Chan. Storage model for processing semantic exceptions. International Journal of Computer Systems Science and Engineering, 9(3):195–204, July 1994. CODEN CSSEEI. ISSN 0267-6192.

Mai:2015:DFS

[MLK15] Hai Tanh Mai, Ki Yong Lee, and Myoung Ho Kim. Dynamic filtering of sensor readings for approximate query processing in sensor networks. International Journal of Computer Systems Science and Engineering, 30(3):??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Miura:1994:FVL

[MLLS94] Yukiya Miura, Jung Li, Wang Li, and Naito Sachio. Fourvalued logic system and its application to describing the specification and design verification of combinational circuits. Computer Systems Science and Engineering, 9(1): 38–45, January 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Mataoui:2015:DBA

[MM15] M’hamed Mataoui and Mohamed Mezghiche. A distance based approach for link analysis in XML information retrieval. International Journal of Computer Systems Science and Engineering, 30(3): ??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Michaelides:1999:EUA

[MMD99] Danius Michaelides, Luc Moreau, and David De Roure. Editorial: Uniform approach to programming the World Wide Web. International Journal of Computer Systems Science and Engineering, 14 (2):69–81, March 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Moumen:2016:EEA

[MMH16] Chiraz Soussi Moumen, Franck Morvan, and Abdelkader Hameurlain. Estimation error-aware query optimization: an overview. International Journal of Computer Systems Science and Engineering, 31(3):??, May 2016. CODEN CSSEEI. ISSN 0267-6192.

Mori:1985:ADC

[MMI85] Kinji Mori, Shyoji Miyamoto, and Hirokazu Ihara. Autonomous decentralized computer system and software structure. Computer Systems Science and Engineering, 1 (1):17–22, October 1985. CO-DEN CSSEEI. ISSN 0267-6192.

Min:2000:CHT

[MMO00] Y. Min, L. L. Miller, and M. M. Owrang. A calculus for handling time in complex objects. International Journal of Computer Systems Science and Engineering, 15(4): 253–??, July 2000. CODEN CSSEEI. ISSN 0267-6192.

Motus:1998:FTA

[MN98] L. Motus and T. Naks. Formal timing analysis of OMT designs using LIMITS. International Journal of Computer Systems Science and Engineering, 13(3):161–170, May 1998. CODEN CSSEEI. ISSN 0267-6192.

Mala:2014:PEI

[MN14] C. Mala and B. Nithya. Performance evaluation of IEEE 802.11 contention control mechanisms using polygonal sequences. International Journal of Computer Systems Science and Engineering, 29 (6):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Moghadam:2014:NRD

[MNK14] Marian Zomorodi Moghadam, Keivan Navi, and Mahmood Kalemati. A novel reversible design for double edge triggered flip-flops and new designs of reversible sequential circuits. International Journal of Computer Systems Science and Engineering, 29(3):

??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Moran:2015:CSG

[MOEMK^{+}15] Alberto L. Morán, Felipe Orihuela-Espina, Victoria Meza-Kubo, Ana I. Grimaldo, Cristina Ramírez-Fernández, Eloísa García-Canseco, Juan M. Oropeza-Salas, and Luis E. Sucar. Out of context serious games: Transversal reutilization of games across healthcare domains. International Journal of Computer Systems Science and Engineering, 30(1): ??, January 2015. CODEN CSSEEI. ISSN 0267-6192.

Min:2003:PAH

[MOK03] G. Min and M. Ould-Khaoua.

Performance analysis of hypercube networks under LRD traffic. International Journal of Computer Systems Science and Engineering, 18(4): ??, July 2003. CODEN CSSEEI. ISSN 0267-6192.

Morzy:2007:CBA

[Mor07]  Mikolaj Morzy.  Cluster-

based analysis and recommendation of sellers in online auctions. International Journal of Computer Systems Science and Engineering, 22(5): ??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Maly:2003:FDR

[MOZR03] K. Maly, S. Olariu, L. Zhang, and N. Rao. Fairness in DQDB revisited: a new solution. International Journal of Computer Systems Science and Engineering, 18(2): ??, March 2003. CODEN CSSEEI. ISSN 0267-6192.

Mancini:1988:HLD

[MP88] Luigi Mancini and Giuseppe Pappalardo. Hoare logic of distributed redundant systems. Computer Systems Science and Engineering, 3(4): 171–180, October 1988. CO-DEN CSSEEI. ISSN 0267-6192.

Marinos:1991:DIM

[MPC91] L. Marinos, M. P. Papazoglou, and D. Christodoulakis. Data integration methodology for an office environment. Computer Systems Science and Engineering, 6(3):143–??, July 1991. CODEN CSSEEI. ISSN 0267-6192.

Mansor:1991:DIE

[MRDF91] Dzaharudin Mansor, Peter Richardson, Tharam S. Dillon, and Kevin E. Forward. Design, implementation and evaluation of a fault-tolerant signalling network management system. Part II. Computer Systems Science and Engineering, 6(4):211–226, October 1991. CODEN CSSEEI. ISSN 0267-6192.

Muller:1998:PCE

[MRSW98] Henk L. Muller, Sanjay Raina, Paul W. A. Stallard, and David H. D. Warren. Parallel calibrated emulation as a technique for evaluating parallel architectures. International Journal of Computer Systems Science and Engineering, 13(1):17–25, January 1998. CODEN CSSEEI. ISSN 0267-6192.

Muzio:1991:DCB

[MS91] J. C. Muzio and M. Serra.

Data compaction for bridging faults. Computer Systems Science and Engineering, 6(3): 131–??, July 1991. CODEN CSSEEI. ISSN 0267-6192.

Manivannan:1996:DTG

[MS96] D. Manivannan and M. Singhal. Decentralized token generation scheme for tokenbased mutual exclusion algorithms. International Journal of Computer Systems Science and Engineering, 11(1): 45–54, January 1996. CO-DEN CSSEEI. ISSN 0267-6192.

Makedon:1997:HMP

[MS97] Fillia Makedon and Antonis Symvonis. Heuristic methods for permutation routing on meshes and tori. International Journal of Computer Systems Science and Engineering, 12 (6):373–380, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Mullen:1992:FFO

[Mul92] James G. Mullen. FBASE:

a federated objectbase system. Computer Systems Science and Engineering, 7(2): 91–99, April 1992. CODEN CSSEEI. ISSN 0267-6192.

Mizanian:2009:MAF

[MVA09] Kiarash Mizanian, Mehdi Vasef, and Morteza Analoui. The modified alternating fractal renewal process for modelling peer to peer traffic. International Journal of Computer Systems Science and Engineering, 24(6): ??, November 2009. CODEN CSSEEI. ISSN 0267-6192.

Murray:1990:WDO

[MW90] Kevin A. Murray and Andy J. Wellings. Wisdom: a distributed operating system for transputers. Computer Systems Science and Engineering, 5(1):13–20, January 1990. CODEN CSSEEI. ISSN 0267-6192.

Murakami:2010:DMS

[MYU10] Makoto Murakami, Motoi Yamagiwa, and Minoru Uehara. Decision model for a system to start communicating with a human. International Journal of Computer Systems Science and Engineering, 25(2): ??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Murata:2006:PNM

[MYYW06] T. Murata, J. Yim, H. Yin, and O. Wolfson. Petri-net model and minimum cycle time for updating a moving objects database. International Journal of Computer Systems Science and Engineering, 21(3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Manivannan:2003:PCI

[MZ03] D. Manivannan and C. Zhang.

Performance of communicationinduced checkpointing algorithms. International Journal of Computer Systems Science and Engineering, 18(3):??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Nanya:1994:ACD

[Nan94] Takashi Nanya. Anti-code-disjoint mapping for exception handling in self-checking systems hierarchy. Computer Systems Science and Engineering, 9(1):46–53, January 1994. CODEN CSSEEI. ISSN 0267-6192.

Newe:2003:FVL

[NC03] T. Newe and T. Coffey. Formal verification logic for hybrid secutity protocols. International Journal of Computer Systems Science and Engineering, 18(1):??, January 2003. CODEN CSSEEI. ISSN 0267-6192.

Newe:2004:FVL

[NC04] T. Newe and T. Coffey. Formal verification logic for hybrid security protocols. International Journal of Computer Systems Science and Engineering, 19(1):??, January 2004. CODEN CSSEEI. ISSN 0267-6192.

Nikraz:2006:MDM

[NCB06] Magid Nikraz, Giovanni Caire, and Parisa A. Bahri. A methodology for the development of multi-agent systems using the JADE platform. International Journal of Computer Systems Science and Engineering, 21(2):??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Napier:2000:FSV

[NCMH00] J. Napier, L. Chen, J. May, and G. Hughes. Fault simulation to validate faulttolerance in Ada. International Journal of Computer Systems Science and Engineering, 15(1):61–??, January 2000. CODEN CSSEEI. ISSN 0267-6192.

Nguyen:2007:ONR

[ND07] Kinh Nguyen and Tharam Dillon. Obligation nets: a rigorous object-oriented technique for modeling the behavioural semantics of information systems. International Journal of Computer Systems Science and Engineering, 22 (4):??, July 2007. CODEN CSSEEI. ISSN 0267-6192.

Nemeth:1996:STI

[Nem96]  Gibor Nemeth.  Schedul-

ing, timing and intractability in massively parallel systems. International Journal of Computer Systems Science and Engineering, 11(4):245– 254, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Nguyen:2014:IEG

[NHK14] Cuong Xuan Nguyen, Hung-Hsuan Huang, and Kyoji Kawagoe. Implementation and evaluations of graphical password using object-based image ranking. International Journal of Computer Systems Science and Engineering, 29 (4):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Naghibzadeh:2003:YFR

[NK03] M. Naghibzadeh and K. H. (Kane)

Kim. The yielding-first ratemonotonic scheduling approach and its efficiency assessment. International Journal of Computer Systems Science and Engineering, 18(3): ??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Nah:2007:CBT

[NKWK07] Y. Nah, K. H. (Kane)

Kim, T. Wang, and M. H. Kim. A cluster-based TMOstructured scalable approach for location information systems. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Narasimhan:2002:SCR

[NMMS02] P. Narasimhan, L. E. Moser, and P. M. Melliar-Smith.

Strongly consistent replication and recovery of faulttolerant CORBA applications. International Journal of Computer Systems Science and Engineering, 17(2): ??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Navarro:2010:OTO

[NPR10] Pedro Luis Mateo Navarro, Gregorio Martínez Pêrez, and Diego Sevilla Ruiz. OpenHMI-tester: an open and cross-platform architecture for GUI testing and certification. International Journal of Computer Systems Science and Engineering, 25 (4):??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Nakajima:2007:SIB

[NTI^{+}07] T. Nakajima, E. Tokunaga, H. Ishikawa, D. Ueno, K. Fujinami, M. Sugaya, and S. Oikawa. Software infrastructures for building ubiquitous computing environments. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Niu:2016:UCA

[NXG^{+}16] Yitong Niu, Mingming Xiong, Jianyi Guo, Cunli Mao, Yantuan Xian, and Zhengtao Yu. Using cross ambiguity model improves the effect of Vietnamese word segmentation. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Obermaisser:2007:MVD

[OESHK07] R. Obermaisser, C. El-Salloum, B. Huber, and H. Kopetz. Modeling and verification of distributed realtime systems using periodic finite state machines. International Journal of Computer Systems Science and Engineering, 22(6):??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Obermaisser:2008:MVD

[OESHK08] R. Obermaisser, C. El-Salloum, B. Huber, and H. Kopetz. Modeling and verification of distributed realtime systems using periodic finite state machines. International Journal of Computer Systems Science and Engineering, 23(4):??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Omondi:1991:DHP

[Omo91] Amos R. Omondi. Design of a high performance instruction pipeline. Computer Systems Science and Engineering, 6(1):13–29, January 1991. CODEN CSSEEI. ISSN 0267-6192.

Okaie:2012:NFG

[ON12] Yutaka Okaie and Tadashi Nakano. Network formation games in non-cooperative service overlay networks. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Oberholzer:2006:PCI

[OO06] Hendrick J. G. Oberholzer and Martin S. Olivier. Privacy contracts incorporated in a privacy protection framework. International Journal of Computer Systems Science and Engineering, 21(1): ??, January 2006. CODEN CSSEEI. ISSN 0267-6192.

Okajima:2012:HVM

[OO12] Seiji Okajima and Yoshihiro Okada. Hierarchical visual motion retrieval system for distributed motion DB and its user experiments. International Journal of Computer Systems Science and Engineering, 27(1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

ORyan:2002:PPC

[OSN02] C. O’Ryan, D. C. Schmidt, and J. R. Noseworthy. Patterns and performance of a CORBA event service for large-scale distributed interactive simulations. International Journal of Computer Systems Science and Engineering, 17(2):??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Olariu:1995:FAC

[OSZ95] S. Olariu, J. L. Schwing, and J. Zhang. Fast adaptive convex hull algorithm on twodimensional processor arrays with a reconfigurable bus system. International Journal of Computer Systems Science and Engineering, 10(3):131– 137, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Olariu:1997:TCO

[OSZZ97] Stephan Olariu, James L.

Schwing, Jingyuan Zhang, and Albert Zomaya. A timeand cost-optimal algorithm for channel assignment. International Journal of Computer Systems Science and Engineering, 12(3):205–211, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Praba:2010:MAC

[PA10] V. Lakshmi Praba and G. Arumugam. Message authentication code algorithm for IP-SEC. International Journal of Computer Systems Science and Engineering, 25 (5):??, September 2010. CO-DEN CSSEEI. ISSN 0267-6192.

Papadopoulos:1994:POR

[Pap94] Constantinos V. Papadopoulos. Provably optimal routing algorithms for signal routing. International Journal of Computer Systems Science and Engineering, 9(4):211– 219, October 1994. CODEN CSSEEI. ISSN 0267-6192. Papadopoulos:1998:MPM

[Pap98] George A. Papadopoulos. A multimedia programming model based on timed concurrent constraint programming. International Journal of Computer Systems Science and Engineering, 13(4): 195–205, July 1998. CODEN CSSEEI. ISSN 0267-6192. Papadopoulos:1999:OOT

[Pap99] George A. Papadopoulos. Object-oriented term graph rewriting. International Journal of Computer Systems Science and Engineering, 14(1): 39–50, January 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Pereira:2002:E

[PB02] C. E. Pereira and R. Baldoni. Editorial. International Journal of Computer Systems Science and Engineering, 17(2): ??, March 2002. CODEN CSSEEI. ISSN 0267-6192. Pentzaropoulos:1986:PMS

[PBS86] G. C. Pentzaropoulos, G. C. Barney, and W. Swindells. Performance modelling of Serc Prime computer system. Computer Systems Science and Engineering, 1(2):109– 113, January 1986. CODEN CSSEEI. ISSN 0267-6192.

Porenta:2013:CCI

[PC13] Jernej Porenta and Mojca Ciglaric. Comparing commercial IP reputation databases to open-source IP reputation algorithms. International Journal of Computer Systems Science and Engineering, 28 (1):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Pirmez:2010:ELD

[PCB10] Luci Pirmez, Luiz F. R. C.

Carmo, and Luiz F. H. Bacellar. Enhancing Levenshtein distance algorithm for assessing behavioral trust. International Journal of Computer Systems Science and Engineering, 25(1):??, January 2010. CODEN CSSEEI. ISSN 0267-6192.

Papazoglou:1995:UMF

[PDK95] Mike P. Papazoglou, Alex Denis, and Bernd J. Krämer. Use of middleware facilities in interoperable databases. International Journal of Computer Systems Science and Engineering, 10(4):195–206, October 1995. CODEN CSSEEI. ISSN 0267-6192.

Petersohn:1998:FSW

[PHPd98] Carsta Petersohn, Cornelis Huizing, Jan Peleska, and Willem-Paul de Roever. Formal semantics for Ward and Mellor’s transformation schemas and its application to fault tolerant systems. International Journal of Computer Systems Science and Engineering, 13(2):131– 136, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Park:2007:SSB

[PKC07] S. Park, S-W Kim, and W. W. Chu. SBASS: Segment based approach for subsequence searches in sequence databases. International Journal of Computer Systems Science and Engineering, 22(1–2):??, January/March 2007. CODEN CSSEEI. ISSN 0267-6192.

Pakzad:1988:INF

[PL88] Simin Pakzad and S. Lakshmivarahan. Interconnection networks and fault tolerance. Computer Systems Science and Engineering, 3(2): 91–99, April 1988. CODEN CSSEEI. ISSN 0267-6192.

Park:2000:FTM

[PLL00] J.-H. Park, H.-K. Lee, and C.- S. Lee. A fault tolerant multistage interconnection network and its fault tolerant diagnosis: the ring-banyan network. International Journal of Computer Systems Science and Engineering, 15(2):119– ??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Pombortsis:1994:MCR

[PM94] A. Pombortsis and A. Menos. Modelling a class of resource sharing interconnection networks using timed Petri nets. International Journal of Computer Systems Science and Engineering, 9 (3):155–163, July 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Polze:1999:RSC

[PM99] Andreas Polze and Miroslaw Malek. Responsive services with CORBA. International Journal of Computer Systems Science and Engineering, 14 (4):209–216, July 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Penna:2016:EVI

[PMO16] Giuseppe Della Penna, Daniele Magazzeni, and Sergio Orefice. Extending visual information extraction to biomedical applications. International Journal of Computer Systems Science and Engineering, 31 (5):??, September 2016. CO-DEN CSSEEI. ISSN 0267-6192.

Poursalidis:2007:TPC

[PN07] V. Poursalidis and C. Nikolaou. Towards a personcentric Identity Management Infrastructure (IMI). International Journal of Computer Systems Science and Engineering, 22(5):??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Ping:2009:ACN

[PN09] Yang Ping and Liao Ningbo.

Approach on complex neuralgenetic algorithm modeling for isomorphism identification in conceptual design of mechanism. International Journal of Computer Systems Science and Engineering, 24(6): ??, November 2009. CODEN CSSEEI. ISSN 0267-6192.

Poursakidis:2010:TPC

[PN10] V. Poursakidis and C. Nikolaou. Towards a personcentric Identity Management Infrastructure (IMI). International Journal of Computer Systems Science and Engineering, 25(1):??, January 2010. CODEN CSSEEI. ISSN 0267-6192.

Perez:2011:MTT

[PONA11] Jon Perez, Roman Obermaisser, Carlos F. Nicolas, and Iban Ayestaran. Modeling time-triggered real-time control systems using executable time-triggered model (E-TTM) and systemC-AMS. International Journal of Computer Systems Science and Engineering, 26(6):??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Posch:1998:MPC

[Pos98] Reinhard Posch. Massive parallelism on a chip: VLSI aspects involving dynamic logic. International Journal of Computer Systems Science and Engineering, 13(2):101– 107, March 1998. CODEN CSSEEI. ISSN 0267-6192.

Potdar:2014:GE

[Pot14] Vidyasagar Potdar. Guest editor. International Journal of Computer Systems Science and Engineering, 29(1):

??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Prema:2013:AMT

[PRB13] P. Prema, B. Ramadoss, and S. R. Balasundaram. An algorithm for merging two classification trees for test case generation. International Journal of Computer Systems Science and Engineering, 28(4): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Paulo:2013:TAE

[PRPS13] João Paulo, Pedro Reis, José Pereira, and António Sousa. Towards an accurate evaluation of deduplicated storage systems. International Journal of Computer Systems Science and Engineering, 28(6): ??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Polze:2001:TPC

[PRSM01] A. Polze, J. Richling, J. Schwarz, and M. Malek. Towards predictable CORBA-based services. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Pradeep:1995:SMC

[PS95] B. Pradeep and C. Siva Ram Murthy. Sorting on meshconnected computers with multiple broadcasting using N^{2/3}×N^{2/3} processors. International Journal of Computer Systems Science and Engineering, 10(2):100–110, April 1995. CODEN CSSEEI. ISSN 0267-6192.

Pezze:2001:E

[PS01] M. Pezzè and S. M. Shatz. Editorial. International Journal of Computer Systems Science and Engineering, 16(3): ??, May 2001. CODEN CSSEEI. ISSN 0267-6192.

Petrinja:2010:TFD

[PS10] Etiel Petrinja and Giancarlo Succi. Trustworthiness of the FLOSS development process. International Journal of Computer Systems Science and Engineering, 25(4): ??, July 2010. CODEN CSSEEI. ISSN 0267-6192.

Peng:2009:AAI

[PTX^{+}09] Chao Peng, Yasuo Tan, Naixue Xiong, Laurence T. Yang, and Hong Zhu. Approximation algorithms for inner-node weighted minimum spanning trees. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Pissinou:1996:ADR

[PV96] Niki Pissinou and Kanonkluk Vanapipat. Active database rules in distributed database systems: a dynamic approach to solving structural and semantic conflicts in distributed database systems. International Journal of Computer Systems Science and Engineering, 11(1):35–44, January 1996. CODEN CSSEEI. ISSN 0267-6192.

Posadas:2011:ETA

[PVRM11] H. Posadas, E. Villar, Dominique Ragot, and Marcos Martinez. Early, timeapproximate modeling of multi-OS Linux platforms in a systemC co-simulation environment. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Pi:2008:ICM

[PXQ08] Dechang Pi, Chuhua Xian, and Xiaolin Qin. An improved C-means clustering algorithm. International Journal of Computer Systems Science and Engineering, 23(1): ??, January 2008. CODEN CSSEEI. ISSN 0267-6192.

Ping:1993:XRA

[PXY93] Chen Ping, Cai Xiyao, and Jin Yimin. XDRC++: a reflective architecture in the C++ language. Computer Systems Science and Engineering, 8(2):109–120, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Perkins:2004:QAR

[PYHO04] Dmitri D. Perkins, Jinping Yang, Herman Hughes, and Charles Owen. A QoS-aware routing scheme for mobile ad hoc networks. International Journal of Computer Systems Science and Engineering, 19 (5):??, September 2004. CO-DEN CSSEEI. ISSN 0267-6192.

Qin:1993:AAD

[QA93] Bin Qin and Reda A. Ammar. Analytic approach to deriving time costs of parallel computations. Computer Systems Science and Engineering, 8(2):90–100, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Qin:1991:APP

[QAS91] Bin Qin, Reda A. Ammar, and Howard A. Sholl. Allocating processing power to minimize time costs in parallel software systems. Computer Systems Science and Engineering, 6(3):159–??, July 1991. CODEN CSSEEI. ISSN 0267-6192.

Quan:2009:ICS

[QAY09] Dang Ming Quan, Jörn Altmann, and Laurence T. Yang. Improving the capability of the SLA workflow broker with parallel processing technology. International Journal of Computer Systems Science and Engineering, 24(5): ??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Qu:2009:OLA

[QKSN09] Wenyu Qu, Masaru Kitsuregawa, Yanming Shen, and Takashi Nanya. An optimal lifetime-adaptive method for wireless sensor networks. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Qu:2007:TBR

[QKZ^{+}07] Wenyu Qu, Masaru Kitsuregawa, Hai Zhuge, Hong Shen, and Yinwei Jin. A trafficbased routing algorithm by using mobile agents. International Journal of Computer Systems Science and Engineering, 22(6):??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Qin:2013:SIB

[QL13] ChaoYong Qin and Chenhan Liao. Session identification based on linked referrers and Web log indexing. International Journal of Computer Systems Science and Engineering, 28(5):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Qin:2015:SIB

[QL15] ChaoYong Qin and Chenhan Liao. Session identification based on linked referrers and web log indexing. International Journal of Computer Systems Science and Engineering, 30(2):??, March 2015. CODEN CSSEEI. ISSN 0267-6192.

Qi:2014:HKI

[QLL14] Heng Qi, Guoyu Lan, and Kequi Li. A hierarchical kernel for image classification based on layered representation. International Journal of Computer Systems Science and Engineering, 29(4):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Qaroush:2014:ESF

[QWK14] Aziz Qaroush, Mahdi Washaha, and Ismail Khater. Efficient spam filtering based on informative features extracted from the header fields and the URLs in the message. International Journal of Computer Systems Science and Engineering, 29(1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Qu:2015:TDA

[QY15] Qingjun Qu and Yiping Yao. A task distribution algorithm compatible with background control framework for parallel simulation. International Journal of Computer Systems Science and Engineering, 30 (5):??, September 2015. CO-DEN CSSEEI. ISSN 0267-6192.

Qin:2010:CSC

[QYCG10] Guangcheng Qin, Panlong Yang, Guihai Chen, and Deke Guo. Characterizing the scaling capacity for multiple access wireless sensor networks. International Journal of Computer Systems Science and Engineering, 25(6): ??, November 2010. CODEN CSSEEI. ISSN 0267-6192.

Qu:2009:PIR

[QZK09] Wenyu Qu, Wanlei Zhou, and Masaru Kitsuregawa. A parallel information retrieval method for e-commerce. International Journal of Computer Systems Science and Engineering, 24(5):??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Rebhi:2015:FDL

[RAF15] A. Rebhi, S. Abid, and F. Fnaiech. Fabric defect localization using line variances of the local homogeneity images. International Journal of Computer Systems Science and Engineering, 30 (4):??, July 2015. CODEN CSSEEI. ISSN 0267-6192.

Ravikumar:1992:PDP

[Rav92] C. P. Ravikumar. Parallel r-dimensional placement on a vector minisupercomputer. Computer Systems Science and Engineering, 7(3): 147–151, July 1992. CODEN CSSEEI. ISSN 0267-6192.

Ravikumar:1995:PDP

[Rav95] C. P. Ravikumar. Parallel r-dimensional placement on a vector minisupercomputer. International Journal of Computer Systems Science and Engineering, 10(3): 138–143, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Raynal:2000:NBA

[Ray00] M. Raynal. Non-blocking atomic commitment in distributed systems: a tutorial based on a generic protocol. International Journal of Computer Systems Science and Engineering, 15(2): 77–??, March 2000. CODEN CSSEEI. ISSN 0267-6192.

Raynal:2002:TBS

[Ray02] M. Raynal. Token-based sequential consistency. International Journal of Computer Systems Science and Engineering, 17(6):??, November 2002. CODEN CSSEEI. ISSN 0267-6192.

Ravikumar:1998:VIS

[RC98] C. P. Ravikumar and S. Chakraverty. VLSI implementation of a strongly fault secure ALU. International Journal of Computer Systems Science and Engineering, 13(5):283–288, September 1998. CODEN CSSEEI. ISSN 0267-6192.

Rafferty:2015:GLO

[RCNL15] Joseph Rafferty, Liming Chen, Chris Nugent, and Jun Liu. Goal lifecycles and ontological models for intention based assistive living within smart environments. International Journal of Computer Systems Science and Engineering, 30(1):??, January 2015. CODEN CSSEEI. ISSN 0267-6192.

Ross:1992:CSP

[RD92] D. F. Ross and R. H. Davis. Comparative study of performance of concurrency control algorithms in a centralized database. Computer Systems Science and Engineering, 7(1):3–16, January 1992. CODEN CSSEEI. ISSN 0267-6192.

Rodriguez-Dominguez:2015:MDA

[RDRLG^{+}15] Carlos Rodríguez-Domínguez, Tomás Ruiz-López, José Luis Garrido, Manuel Noguera, and Kawtar Benghazi. A model-driven approach to service composition on the basis of the specification of BPMN choreographies. International Journal of Computer Systems Science and Engineering, 30 (1):??, January 2015. CO-DEN CSSEEI. ISSN 0267-6192.

Reddi:1988:SMA

[Red88] Arumalla V. Reddi. Secondary memory analysis of a mini/microcomputer in resource-sharing distributed information systems. Computer Systems Science and Engineering, 3(4):189–198, October 1988. CODEN CSSEEI. ISSN 0267-6192.

Reddi:1991:MPM

[Red91] Arumalla V. Reddi. Memory performance of microcomputers at large user nodes of distributed systems. Computer Systems Science and Engineering, 6(2):85–??, April 1991. CODEN CSSEEI. ISSN 0267-6192.

Ranran:2016:HGB

[RES^{+}16a] Liu Ranran, Zheng Enxing, Chang Shan, Bei Shaoyi, and Zhang Lanchun. Hierarchical gradient based parameter identification for nonuniformly sampling Wiener systems. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Ranran:2016:HSG

[RES^{+}16b] Liu Ranran, Zheng Enxing, Chang Shan, Bei Shaoyi, and Zhang Lanchun. Hierarchical stochastic gradient identification for non-uniformly sampling Hammerstein systems with colored noise. International Journal of Computer Systems Science and Engineering, 31(6):??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Rodriguez:2007:MAD

[RFMP07] Alfonso Rodríguez, Eduardo Fernández-Medina, and Mario Piattini. An MDA approach to develop secure business processes through a UML 2.0 extension. International Journal of Computer Systems Science and Engineering, 22(5): ??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Robach:1988:TMR

[RG88] C. Robach and S. Guibert. Testability measures: a review. Computer Systems Science and Engineering, 3(3): 117–126, July 1988. CODEN CSSEEI. ISSN 0267-6192.

Rahman:2014:DCA

[RHS^{+}14] Mohammad Osiur Rahman, Aini Hussain, Edgar Scavino, M. A. Hannan, and Hassan Basri. DNA computation algorithm for recyclable waste paper sorting. International Journal of Computer Systems Science and Engineering, 29 (5):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Richter:1990:MDV

[Ric90] Harald Richter. Multiprocessor with dynamically variable topology. Computer Systems Science and Engineering, 5 (1):29–35, January 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Rahwan:2006:ISM

[RJS06] Iyad Rahwan, Thomas Juan, and Leon Sterling. Integrating social modelling and agent interaction through goal-oriented analysis. International Journal of Computer Systems Science and Engineering, 21(2):??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Rundensteiner:1999:EES

[RKZ99] E. A. Rundensteiner, H. A.

Kuno, and Y. Zhou. Experimental evaluation of the SMX strategy for incremental materialized path view maintenance. International Journal of Computer Systems Science and Engineering, 14(6):331– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Raju:1998:PCM

[RM98] N. N. Raju and Ravi Mittal. Performance comparison of multicasting schemes for ATM switches under heterogeneous traffic. International Journal of Computer Systems Science and Engineering, 13 (6):359–368, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Rao:2002:GQP

[RM02] N. S. V. Rao and N. Manickam. General quickest paths and path-tables. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/ September 2002. CODEN CSSEEI. ISSN 0267-6192.

Ranganathan:1991:VAD

[RMB91] N. Ranganathan, Amar Mukherjee, and M. Bassiouni. VLSI algorithms for data compression. Computer Systems Science and Engineering, 6(4): 238–253, October 1991. CO-DEN CSSEEI. ISSN 0267-6192.

Richardson:1991:FTA

[RMDF91] Peter Richardson, Dzaharudin Mansor, T. S. Dillon, and K. E. Forward. Fault-tolerant approach to the design of a Signalling Transfer Point: Part 1. Computer Systems Science and Engineering, 6 (3):178–??, July 1991. CO-DEN CSSEEI. ISSN 0267-6192.

Raposo:2000:PNB

[RMR00] A. B. Raposo, L. P. Magalhães, and I. L. M. Ricarte. Petri nets based coordination mechanisms for multiworkflow environments. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:1994:RUC

[Rom94] Alexander B. Romanovsky.

Recovery unification for concurrent heterogeneous systems. International Journal of Computer Systems Science and Engineering, 9(4):234– 239, October 1994. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:1998:PTD

[Rom98] Alexander B. Romanovsky.

Predictable toleration of design faults: Recovery blocks in real time systems. International Journal of Computer Systems Science and Engineering, 13(6):369–377, November 1998. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:2000:DED

[Rom00] A. Romanovsky. The domino effect as a deadlock. International Journal of Computer Systems Science and Engineering, 15(3):147–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:2002:VSR

[Rom02] A. Romanovsky. On version state recovery and adjudication in class diversity. International Journal of Computer Systems Science and Engineering, 17(3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Ravikumar:1990:PIS

[RP90] C. P. Ravikumar and L. M. Patnaik. Performance improvement of simulated annealing algorithms. Computer Systems Science and Engineering, 5(2):111–115, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Rathnayaka:2013:PRF

[RPD^{+}13] A. J. Dinusha Rathnayaka, Vidyasagar M. Potdar, Tharam Dillon, Omar Hussain, and Samitha Kuruppu. Prosumer recruitment framework for prosumer community groups in smart-grid. International Journal of Computer Systems Science and Engineering, 28 (5):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Rathnayaka:2015:PRF

[RPDH15] A. J. Dinusha Rathnayaka, Vidyasagar M. Potdar, Tharam Dillon, and Omar Hussain. Prosumer recruitment framework for prosumer community groups in smart-grid. International Journal of Computer Systems Science and Engineering, 30(4):??, July 2015. CODEN CSSEEI. ISSN 0267-6192.

Rego:1988:CND

[RS88] Vernon Rego and Wojciech Szpankowski. Closed-network duals of multiqueues with application to token-passing systems. Computer Systems Science and Engineering, 3(3): 127–139, July 1988. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:1993:URN

[RS93] A. B. Romanovsky and I. V.

Shturtz. Unplanned recovery for non-program objects. Computer Systems Science and Engineering, 8(2):72–79, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:1996:DC

[RS96] A. B. Romanovsky and I. V.

Shturtz. Dynamic conversions. International Journal of Computer Systems Science and Engineering, 11(2):109– 116, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Raut:2014:SCP

[RS14] Manoj K. Raut and Arindama Singh. A survey on computing prime implicants and implicates in classical and nonclassical logics. International Journal of Computer Systems Science and Engineering, 29 (5):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

H:2014:RRD

[RSBW14] Ari Raptino H, Shunsuke Saruwatari, Masaki Bandai, and Takashi Watanabe. Rate and relay diversity in temporal spectrum sharing. International Journal of Computer Systems Science and Engineering, 29(1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Ruta:2010:HZB

[RSDD10] Michele Ruta, Floriano Scioscia, Tommaso Di Noia, and Eugenio Di Sciascio. A hybrid ZigBee/Bluetooth approach to mobile semantic grids. International Journal of Computer Systems Science and Engineering, 25(3): ??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Rishie:1996:LBM

[RSG96] Naphtali Rishie, Artyom Shaposhnikov, and Scott Graham. Load balancing in a massively parallel semantic database. International Journal of Computer Systems Science and Engineering, 11(4): 195–199, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Ravikumar:1991:PMC

[RSP91] C. P. Ravikumar, Sarma Sastry, and L. M. Patnaik. Parallel min-cut placement on reduced hardware SIMD architecture. Computer Systems Science and Engineering, 6 (1):3–??, January 1991. CO-DEN CSSEEI. ISSN 0267-6192.

Rogers:2000:IRC

[RW00] P. Rogers and A. J. Wellings.

An incremental recovery cache supporting software fault tolerance mechanisms. International Journal of Computer Systems Science and Engineering, 15(1):33–??, January 2000. CODEN CSSEEI. ISSN 0267-6192.

Richardson:2011:RRT

[RW11] T. Richardson and A. J.

Wellings. On the road to real-time OSGi: extending OSGi with an admission control protocol. International Journal of Computer Systems Science and Engineering, 26 (6):??, November 2011. CO-DEN CSSEEI. ISSN 0267-6192.

Romanovsky:1999:ECE

[RXR99] Alexander Romanovsky, Jie Xu, and Brian Randell. Editorial: Coordination exception handling in real-time distributed object systems. International Journal of Computer Systems Science and Engineering, 14(4):197–207, July 1999. CODEN CSSEEI. ISSN 0267-6192.

Romanovsky:2001:DCA

[RZ01] A. Romanovsky and A. Zorzo.

A distributed coordinated atomic action scheme. International Journal of Computer Systems Science and Engineering, 16(4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Son:1986:ACC

[SA86] Sang Hyuk Son and Ashok K. Agrawala. Algorithm for concurrency control in replicated databases. Computer Systems Science and Engineering, 1 (2):75–81, January 1986. CO-DEN CSSEEI. ISSN 0267-6192.

Synodinos:2003:LHE

[SA03] D. G. Synodinos and P. Avgeriou. Leveraging hypermedia engineering for multichannel access with the use of XML. International Journal of Computer Systems Science and Engineering, 18(5): ??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Saha:1995:ITA

[Sah95] D. Saha. Improved topology for achieving reliability in high-speed tree LANs. International Journal of Computer Systems Science and Engineering, 10(3):173–178, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Sang:1990:RDD

[San90] Hyuk Son Sang. Reconstruction of distributed databases. Computer Systems Science and Engineering, 5(4):215– 222, October 1990. CODEN CSSEEI. ISSN 0267-6192.

Serrano-Alvarado:2005:AMT

[SARAL05] Patricia Serrano-Alvarado, Claudia Roncancio, Michel Adiba, and Cyril Labbé. An adaptable mobile transaction system. International Journal of Computer Systems Science and Engineering, 20(2): ??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Srinivas:1996:FAD

[SB96a] Sampalli Srinivas and Nripendra N. Biswas. Fast algorithm for data exchange in reconfigurable tree structures. International Journal of Computer Systems Science and Engineering, 11(4):235–243, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Steyaert:1996:DPC

[SB96b] Bart Steyaert and Herwig Bruneel. Delay performance of CBR traffic in an ATM multiplexer with high-priority background traffic. International Journal of Computer Systems Science and Engineering, 11(6):393–399, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Srinivas:1992:PRA

[SBK^{+}92] S. Srinivas, A. Basu, K. G.

Kumar, A. Paulraj, and L. M. Patnaik. Pipelined ring algorithm for matrix multiplication on transputer networks. Performance analysis and estimation. Computer Systems Science and Engineering, 7 (1):42–51, January 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Shu:2009:MFV

[SC09] Gao Shu and Dingfang Chen. Matchmaking framework for visualization Web services. International Journal of Computer Systems Science and Engineering, 24(4): ??, July 2009. CODEN CSSEEI. ISSN 0267-6192.

Sumoza:2010:CCP

[SC10] Rodolfo Sumoza and Jose Aguilar Castro. A cache coherence protocol for distributed memory platforms. International Journal of Computer Systems Science and Engineering, 25 (5):??, September 2010. CO-DEN CSSEEI. ISSN 0267-6192.

Sumoza:2011:CCP

[SC11] Rodolfo Sumoza and Jose Aguilar Castro. A cache coherence protocol for distributed memory platforms. International Journal of Computer Systems Science and Engineering, 26 (1):??, January 2011. CO-DEN CSSEEI. ISSN 0267-6192.

Schroder:1988:ISA

[Sch88]  Heiko Schroder.  Instruc-

tion systolic array — tradeoff between flexibility and speed. Computer Systems Science and Engineering, 3(2): 83–90, April 1988. CODEN CSSEEI. ISSN 0267-6192.

Shimada:2011:SID

[SCKN11] Hiromasa Shimada, Alexandre Courbot, Yuki Kinebuchi, and Tatsuo Nakajima. A software infrastructure for dependable embedded systems. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Srikanthan:1997:TLD

[SCL97] T. Srikanthan, K. Y. Chan, and S. K. Leong. Two-layer dynamic load balancing for a transputer-based vehicle simulator. International Journal of Computer Systems Science and Engineering, 12(6):351– 358, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Souveyet:1999:OTP

[SD99] C. Souveyet and R. Deneckere. Overview of the TOOBIS project: from a temporal extension of ODMG to a patterns-based approach extending the OO model. International Journal of Computer Systems Science and Engineering, 14(6):389–??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Shi:1999:HDM

[SDC99] L. Shi, O. De Vel, and J. Cao. A hierarchical and distributed monitoring system in interprocess communications. International Journal of Computer Systems Science and Engineering, 14(5):317– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Suresha:2014:DTM

[SDN14] M. Suresha, Ajit Danti, and S. K. Narasimhamurthy. Decision trees to multiclass prediction for analysis of arecanut data. International Journal of Computer Systems Science and Engineering, 29 (1):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Spyropoulos:1985:PAP

[SE85] C. D. Spyropoulos and D. J. Evans. Performance analysis of priority-driven algorithm for multiprocessor. Computer Systems Science and Engineering, 1(1):36–46, October 1985. CODEN CSSEEI. ISSN 0267-6192.

Salah:2007:TDA

[SE07] K. Salah and K. Elbadawi. Throughput and delay analysis of interrupt-driven kernels under Poisson and bursty traffic. International Journal of Computer Systems Science and Engineering, 22(1–2):??, January/March 2007. CO-DEN CSSEEI. ISSN 0267-6192.

Squadrito:1999:ASP

[SED^{+}99] Michael Squadrito, Levon Esibov, Lisa Cingiser DiPippo, Victor Fay Wolfe, Gregory Cooper, Bhavani Thurasingham, Peter Krupp, Michel Milligan, Russell Johnston, and Ramachandra Bethmangalkar. The affected set priority ceiling protocols for realtime object-oriented concurrency control. International Journal of Computer Systems Science and Engineering, 14 (4):227–239, July 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Selic:2002:ERT

[Sel02] N. Selic. The emerging realtime UML standard. International Journal of Computer Systems Science and Engineering, 17(2):??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Sait:1996:FSV

[SEuH96] Sadiq M. Sait, Khaled M.

Elleithy, and Masud ul Hasan. Formal synthesis of VLSI layouts from algorithmic specifications. International Journal of Computer Systems Science and Engineering, 11(2): 67–81, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Sharafat:2002:MRA

[SF02] A. R. Sharafat and M. S.

Fallah. A measure of resilience against denial of service attacks in computer networks. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-

Shen:2011:EXQ

[SF11] Yun Shen and Ling Feng. Evaluation of XPath queries with predicates: an Eulerian cycle theory based sequencing approach. International Journal of Computer Systems Science and Engineering, 26 (4):??, July 2011. CODEN CSSEEI. ISSN 0267-6192.

Saksena:1999:CA

[SFLP99] M. Saksena, R. B. France, and M. M. Larrondo-Petrie. A characterization of aggregation. International Journal of Computer Systems Science and Engineering, 14(6):363– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Sarkar:1990:MCA

[SG90] Dilip Sarkar and Ratan Kumar Guha. Mapping a class of algorithms from binary tree machines to linear arrays. Computer Systems Science and Engineering, 5(4): 202–204, October 1990. CO-DEN CSSEEI. ISSN 0267-6192.

Sharma:1998:PSS

[SG98] Gopal Sharma and Bidyut Gupta. Parallel scheduler for a shared memory (tightlycoupled) multiprocessor system. International Journal of Computer Systems Science and Engineering, 13(4):241– 247, July 1998. CODEN CSSEEI. ISSN 0267-6192.

Shatz:1988:SDP

[Sha88] S. M. Shatz. Superprocesses:

a distributed program configuration concept supporting. Computer Systems Science and Engineering, 3(1):3– 12, January 1988. CODEN CSSEEI. ISSN 0267-6192.

Shafarenko:1996:RRP

[Sha96] A. V. Shafarenko. RETRAN:

a recurrent paradigm for data-parallel computing. International Journal of Computer Systems Science and Engineering, 11(4):201–209, July 1996. CODEN CSSEEI. ISSN 0267-6192.

Shatnawi:2009:DTM

[Sha09] Omar Shatnawi. Discrete time modelling in software reliability engineering: a unified approach. International Journal of Computer Systems Science and Engineering, 24(6): ??, November 2009. CODEN CSSEEI. ISSN 0267-6192.

S:2013:AMM

[SHGA13] Vimala S., Khanna Nehemiah H., Saranya G., and Kannan A. Analysis and modeling of multivalued attributes in entity relationship modeling: an approach for improved database design. International Journal of Computer Systems Science and Engineering, 28(4):??, ???? 2013. CODEN CSSEEI. ISSN 0267-

Singh:2015:MUM

[SIGC15] David E. Singh, Florin Isaila, Félix García, and Jesús Carretero. Magio: using mobile agents to enhance parallel I/O. International Journal of Computer Systems Science and Engineering, 30(2): ??, March 2015. CODEN CSSEEI. ISSN 0267-6192.

Shin:1986:NPM

[SK86] Kang G. Shin and C. M. Krishna. New performance measures for design and evaluation of real-time multiprocessors. Computer Systems Science and Engineering, 1(4): 179–192, October 1986. CO-DEN CSSEEI. ISSN 0267-6192.

Sun:1992:AMS

[SK92] Xian-He Sun and Nabil N. Kamel. Augmenting multikey searching structure for general database queries. Computer Systems Science and Engineering, 7(4):229–??, October 1992. CODEN CSSEEI. ISSN 0267-6192.

Sheu:1998:SIR

[SK98] Phillip C.-Y. Sheu and K. H. Kim. Special issue: Real-time object-oriented systems. International Journal of Computer Systems Science and Engineering, 13(3):??, May 1998. CODEN CSSEEI. ISSN 0267-6192.

Singh:2008:OCA

[SKR08] S. P. Singh, C. Kistanna, and A. R. Rao. Optimal capacitor allocation in distribution systems using genetic algorithm. International Journal of Computer Systems Science and Engineering, 23(6): ??, November 2008. CODEN CSSEEI. ISSN 0267-6192.

Shen:2012:SIS

[SKV12] Jun Shen, Soo Dong Kim, and Nalini Venkatasubramanian. Special issue: Service oriented computing and applications. International Journal of Computer Systems Science and Engineering, 27(2): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Shin:1988:PMD

[SL88] Kang G. Shin and Heungkyu Lee. Port manipulator for the distributed realization of an integrated manufacturing system. Computer Systems Science and Engineering, 3(1): 21–31, January 1988. CO-DEN CSSEEI. ISSN 0267-6192.

Semple:1997:DID

[SL97] C. J. Semple and V. Lakshmi Narasimhan. Design and implementation of a distributed system utilising task farming on IBM PC compatible computers and a Novell network. International Journal of Computer Systems Science and Engineering, 12(4):

275–283, July 1997. CODEN CSSEEI. ISSN 0267-6192.

Sharon:2002:RBTb

[SL02] O. Sharon and A. Likholat. On the relation between the throughput gain with slot reuse and the number of address bits in the dual bus configuration. Part 2: Linear and equal throughputs. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Shen:2009:DPV

[SL09] Yanming Shen and Kequi Li. On the design of P2P video streaming with layered coding. International Journal of Computer Systems Science and Engineering, 24(3): ??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Sanati:2012:OGS

[SL12] Farzad Sanati and Jie Lu. An ontology for e-government service integration. International Journal of Computer Systems Science and Engineering, 27(2):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Shin:2014:IDA

[SL14] Se Jung Shin and Won Suk Lee. Interactive discovery of association rules over data streams. International Journal of Computer Systems Science and Engineering, 29(5): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Sousa:2011:AMD

[SLA^{+}11] Helaine Sousa, Denivaldo Lopes, Zair Abdelouahab, Daniela Barreiro Claro, and Slimane Hammoudi. An approach for model driven testing: framework, metamodels and tools. International Journal of Computer Systems Science and Engineering, 26 (4):??, July 2011. CODEN CSSEEI. ISSN 0267-6192.

Seok:2002:CAS

[SLPK02] S-J Seok, S-H Lee, J. Park, and C-H Kang. Characteristics of assured service and an alternative RIO scheme in differentiated services networks. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Srimani:1990:PED

[SM90] Pradip K. Srimani and Shivakant Mishra. Performance evaluation of distributed mutual exclusion algorithms for computer networks. Computer Systems Science and Engineering, 5 (2):59–64, April 1990. CO-DEN CSSEEI. ISSN 0267-

Sadi:2010:CCA

[SMSJ10] Muhammad Sheikh Sadi, D. G. Myers, Cesar Ortega Sanchez, and Jan Jurjens. Component criticality analysis to minimize soft errors risk. International Journal of Computer Systems Science and Engineering, 25(5): ??, September 2010. CODEN CSSEEI. ISSN 0267-6192.

Schlager:2007:EAA

[SN07] Christian Schläger and Thomas Nowey. On the effects of authentication and authorisation infrastructures on ecommerce activities. International Journal of Computer Systems Science and Engineering, 22(5):??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Schemmer:2009:AHR

[SN09] Stefan Schemmer and Edgar Nett. Achieving high reliable and timely task execution in mobile embedded applications. International Journal of Computer Systems Science and Engineering, 24 (1):??, January 2009. CO-DEN CSSEEI. ISSN 0267-6192.

Son:1987:MRC

[Son87] Sang Hyuk Son. On multiversion replication control in distributed systems. Computer Systems Science and Engineering, 2(2):76–84, April 1987. CODEN CSSEEI. ISSN 0267-6192.

Song:2016:AHB

[SP16] Doohee Song and Kwangjin Park. Adaptive hierarchical bitmap-based index for wireless broadcast environments in an emergency. International Journal of Computer Systems Science and Engineering, 31(4):??, July 2016. CODEN CSSEEI. ISSN 0267-6192.

Surendran:2011:PAR

[SPB11] D. Surendran, T. Purusothaman, and R. A. Baclachandar. Performance analysis of a resource aggregator in a grid of grids environment. International Journal of Computer Systems Science and Engineering, 26 (4):??, July 2011. CODEN CSSEEI. ISSN 0267-6192.

Surendran:2012:EES

[SPBK12] D. Surendran, T. Purusothaman, R. A. Balachandar, and G. Kousalya. ES-GIA: Extensible service based grid information aggregator. International Journal of Computer Systems Science and Engineering, 27(4):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Spiliopoulou:1999:LWD

[Spi99] Myra Spiliopoulou. The laborious way from data mining to Web log mining. International Journal of Computer Systems Science and Engineering, 14 (2):113–125, March 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Saxena:2003:CCC

[SR03] P. C. Saxena and J. Rai. Communication cost of kcoteries. International Journal of Computer Systems Science and Engineering, 18(6): ??, December 2003. CODEN CSSEEI. ISSN 0267-6192.

Srimani:1991:HDA

[Sri91] Pradip K. Srimani. Heuristic deadlock avoidance algorithm in a distributed system. Computer Systems Science and Engineering, 6(3): 170–??, July 1991. CODEN CSSEEI. ISSN 0267-6192.

Sur:1993:SDP

[SS93] Sumit Sur and Pradip K. Srimani. Self-diagnosability properties of hypercubes and star graphs and diagnosis algorithms. Computer Systems Science and Engineering, 8 (4):217–223, October 1993. CODEN CSSEEI. ISSN 0267-6192.

Stankovic:1999:AOM

[SS99] John A. Stankovic and Sang H. Son. An architecture and object model for distributed object-oriented realtime databases. International Journal of Computer Systems Science and Engineering, 14 (4):251–259, July 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Stefanovic:2013:IIS

[SS13] Nenad Stefanovic and Dusan Stefanovic. Integrated and interactive software solution for knowledge-based supply network design. International Journal of Computer Systems Science and Engineering, 28 (1):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Sheikh:2007:QFM

[SSB^{+}07] W. Sheikh, B. Shafiq, S. Baqai, R. A. Paul, M. F. Khan, and H. Ghafoor. A QoS framework for multimedia applications in mobile ad hoc networks. International Journal of Computer Systems Science and Engineering, 22(3):??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Shebalin:1991:SSA

[SSC91] Paul V. Shebalin, Sang H.

Son, and Chun-Hyon Chang. Software safety analysis in distributed systems. Computer Systems Science and Engineering, 6(2):102–??, April 1991. CODEN CSSEEI. ISSN 0267-6192.

Sinha:1997:NAS

[SSC97] Amitabha Sinha, Arunava Sinha, and Subrata Choudhury. New approach for splitting a class of real-time algorithms for pipelined processing. International Journal of Computer Systems Science and Engineering, 12(3): 213–220, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Simons:1999:UDP

[SSH99] A. J. H. Simons, M. Snoeck, and K. S. Y. Hung. Using design patterns to reveal the competence of object-oriented methods in system-level design. International Journal of Computer Systems Science and Engineering, 14(6):343– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Sarangi:2003:SEG

[SSP03] S. Sarangi, P. N. Sireesh, and S. P. Pal. A scalable, efficient and general Monte Carlo scheme for generating synthetic Web request streams. International Journal of Computer Systems Science and Engineering, 18(3): ??, May 2003. CODEN CSSEEI. ISSN 0267-6192.

Shigeta:2001:ARC

[SSS01] S. Shigeta, K. Shimizu, and M. Sowa. Access route control by an extended key/lock scheme. International Journal of Computer Systems Science and Engineering, 16(5): ??, September 2001. CODEN CSSEEI. ISSN 0267-6192.

Stafylopatis:1993:PPC

[Sta93] Andreas Stafylopatis. Performance of parallel computations of triangular structure. Computer Systems Science and Engineering, 8(1):

24–32, January 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Stotts:1990:BPE

[Sto90] David P. Stotts. Bounding procedure execution times in a synchronous Petri net computation model. Computer Systems Science and Engineering, 5(4):205–214, October 1990. CODEN CSSEEI. ISSN 0267-6192.

Simao:2013:ADQ

[SV13] José Simão and Luís Veiga.

Adaptability driven by quality of execution in high level virtual machines for shared cloud environments. International Journal of Computer Systems Science and Engineering, 28(6):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Sharon:2002:RBTa

[SVL02] O. Sharon, A. Vainshtein, and A. Likholat. On the relation between the throughput gain with slot reuse and the number of address bits in the dual bus configuration. Part 1: Maximum throughput. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CO-DEN CSSEEI. ISSN 0267-6192.

Sharif:1996:DAS

[SVN96] Hamid R. Sharif, Hamid Vakilzadian, and Sreekanth Nagisetty. Design and analysis of a serial link interconnection network architecture. International Journal of Computer Systems Science and Engineering, 11(5):287– 299, September 1996. CO-DEN CSSEEI. ISSN 0267-6192.

Skalka:2006:TVA

[SW06] C. Skalka and X. Sean Wang. Trust but verify: Authorization for Web services. International Journal of Computer Systems Science and Engineering, 21(5):??, September 2006. CODEN CSSEEI. ISSN 0267-6192.

Schonberger:2012:CAS

[SW12] Andreas Schönberger and Guido Wirtz. Configurable analysis of sequential multiparty choreographies. International Journal of Computer Systems Science and Engineering, 27(2):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Sun:2011:SIR

[SWJF11] Yi Sun, Lizhe Wang, Wei Jie, and Cheng Fu. Special issue — research in escience: current status and future direction. International Journal of Computer Systems Science and Engineering, 26 (3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

SeethaLakshmi:1990:EPP

[SY90] M. Seetha Lakshmi and Philip S. Yu. Effectiveness of parallel processing in database systems. Computer Systems Science and Engineering, 5(2):73–81, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Seng:2004:TMG

[SY04] Jia-Kang Seng and Jing Yu.

Toward a more generic design of XML benchmark workload. International Journal of Computer Systems Science and Engineering, 19(5): ??, September 2004. CODEN CSSEEI. ISSN 0267-6192.

Sait:1999:TDG

[SYNB99] Sadiq M. Sait, Habib Youssef, Khalid Nassar, and Muhammad S. T. Benten. Timing driven genetic placement. International Journal of Computer Systems Science and Engineering, 14(1):3–14, January 1999. CODEN CSSEEI. ISSN 0267-6192.

Seok:2010:HMP

[SYY^{+}10] Seung-Joon Seok, Bongsoo You, Sungwan Youm, Kyung-Hoe Kim, and Chul-Hee Kang. A heuristic multi-path routing scheme for online traffic in MPLS networks. International Journal of Computer Systems Science and Engineering, 25(1):??, January 2010. CODEN CSSEEI. ISSN 0267-6192.

Schreiner:1991:PPD

[SZ91] Franz Schreiner and Gerhard Zimmermann. Pesa-I: a parallel and distributed architecture for production systems. Computer Systems Science and Engineering, 6(2): 67–??, April 1991. CODEN CSSEEI. ISSN 0267-6192.

Sharifi:2003:ORQ

[SZ03] M. Sharifi and B. Zolfaghari. Optimizing the ready queue structure for dynamic scheduling strategies in real-time operating systems. International Journal of Computer Systems Science and Engineering, 18 (6):??, December 2003. CO-DEN CSSEEI. ISSN 0267-6192.

Schantz:2001:OLG

[SZM^{+}01] R. Schantz, J. Zinky, J. Megquier, J. Loyall, D. Karr, and D. Bakken. An object leel gateway supporting quality of service. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Tam:2008:DSB

[Tam08] Francis Tam. On the development of standards based carrier grade platforms. International Journal of Computer Systems Science and Engineering, 23(5):??, September 2008. CODEN CSSEEI. ISSN 0267-6192.

Tan:1997:DLB

[Tan97]  Kian-Lee Tan.  Decou-

pling load-balancing and optimization issues: a twophase query processing framework for shared-nothing systems. International Journal of Computer Systems Science and Engineering, 12(1):25– 36, January 1997. CODEN CSSEEI. ISSN 0267-6192.

Tolksdorf:2004:TBM

[TBEH04] R. Tolksdorf, C. Bizer, R. Eckstein, and R. Heese. Trustable B2C markets on the Semantic Web. International Journal of Computer Systems Science and Engineering, 19 (3):??, May 2004. CODEN CSSEEI. ISSN 0267-6192.

Teyseyre:2005:ERV

[TC05] Alfredo Raúl Teyseyre and Marcelo Campo. Early requirements validation with 3D worlds. International Journal of Computer Systems Science and Engineering, 20 (1):??, January 2005. CO-DEN CSSEEI. ISSN 0267-6192.

Tiba:2006:AOA

[TC06] Fabiano K. T. Tiba and Miriam A. M. Capretz. Agent-oriented architecture for monitoring and diagnosis in supervisory systems. International Journal of Computer Systems Science and Engineering, 21(2):??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Talevski:2007:DRV

[TCD07] A. Talevski, E. Chang, and T. S. Dillon. A dynamically reconfigurable Voice User Interface (VUI) framework and platform. International Journal of Computer Systems Science and Engineering, 22(3): ??, May 2007. CODEN CSSEEI. ISSN 0267-6192.

Tan:1992:ISR

[TD92] P. L. Tan and T. S. Dillon. Incorporating semantics, relationships and exceptions into object oriented data models. Computer Systems Science and Engineering, 7 (2):122–135, April 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Tahir:1997:CED

[TDGNH97] Jamel M. Tahir, Satnam S. Dlay, Raouf N. Gorgui-Naguib, and Oliver R. Hinton. Concurrent error detection in fast Fermat number transform networks. International Journal of Computer Systems Science and Engineering, 12(3): 221–226, May 1997. CODEN CSSEEI. ISSN 0267-6192.

Tang:2011:DFN

[TEQ11] Zhe Tang, Meng Joo Er, and Fang Qi. Dynamic fuzzy neural network for the intelligent control of a humanoid robot. International Journal of Computer Systems Science and Engineering, 26(3): ??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Thomasian:2006:ADA

[TF06] Alexander Thomasian and Gang Fu. Anticipatory disk arm placement to reduce seek time. International Journal of Computer Systems Science and Engineering, 21(3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Torbjornsen:1994:AAM

[TH94] Øystein Torbjørnsen and Svein-Olaf Hvasshovd. Application of an abstract machine supporting fault-tolerance in a parallel database server. Computer Systems Science and Engineering, 9(2):134– 141, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Tsang:1997:EIC

[TH97] Rose P. Tsang and David D. C. Hu. Efficient interconnections for cascading multiple ATM switches. International Journal of Computer Systems Science and Engineering, 12(2):143–157, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Tan:2008:SAD

[THDC08] Henry Tan, Fedja Hadzic, Tharam S. Dillon, and Elizabeth Chang. State of the art of data mining of tree structured information. International Journal of Computer Systems Science and Engineering, 23(4):??, July 2008. CODEN CSSEEI. ISSN 0267-6192.

Theel:1994:MAN

[The94] Oliver Theel. Meeting the application’s needs: a design study of a highly customized replication scheme and its implementation. Computer Systems Science and Engineering, 9(2):89–97, April 1994. CODEN CSSEEI. ISSN 0267-6192.

Thomasian:1997:AAF

[Tho97] Alexander Thomasian. Approximate analyses for fork/ join synchronization in RAID5. International Journal of Computer Systems Science and Engineering, 12(5):329–337, September 1997. CODEN CSSEEI. ISSN 0267-6192.

Thomasian:2013:ISS

[Tho13] Alexander Thomasian. Improved storage system performance by disk scheduling. International Journal of Computer Systems Science and Engineering, 28(2):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Tanaka:1998:OBC

[THT98] Katsuya Tanaka, Hiroaki Higaki, and Makoto Takizawa. Object-based checkpoints in distributed systems. International Journal of Computer Systems Science and Engineering, 13(3):179–185, May 1998. CODEN CSSEEI. ISSN 0267-6192.

Tseng:2001:GGO

[TJ01] Y.-M. Tseng and J.-K. Jan.

Generalized group-oriented cryptosystem with authenticated sender. International Journal of Computer Systems Science and Engineering, 16 (5):??, September 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Taivan:2015:WBA

[TJS15] Constantin Taivan, Rui José, and Bruno Silva. Web-based applications for open display networks: developers’ perspective. International Journal of Computer Systems Science and Engineering, 30(1): ??, January 2015. CODEN CSSEEI. ISSN 0267-6192.

Tu:2011:MBM

[TJY^{+}11] Xuping Tu, Hai Jin, Zhibin Yu, Jie Chen, Yabin Hu, and Xie Xia. MT-BTRIMER: A master-slave multi-threaded dynamic binary translator. International Journal of Computer Systems Science and Engineering, 26(5): ??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Tsichiya:2000:DEW

[TK00] T. Tsichiya and T. Kikuno.

Dependable evaluation of the weighted voting scheme in the presence of node and link failures. International Journal of Computer Systems Science and Engineering, 15(3): 181–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Taniar:2004:UHO

[TKCR04] David Taniar, Hui Yee Khaw, Haorianto Cokrowijoyo, and Johanna Wenny Rahayu. The use of hints in objectrelational query optimization. International Journal of Computer Systems Science and Engineering, 19(6): ??, November 2004. CODEN CSSEEI. ISSN 0267-6192.

Tan:1995:SHS

[TL95] Kian-Lee Tan and Hongjun Lu. Skew handling strategies for pipelined processing of multi-join queries in sharednothing systems. International Journal of Computer Systems Science and Engineering, 10(1):3–18, January 1995. CODEN CSSEEI. ISSN 0267-6192.

Tran:2006:MFO

[TLB06] Quynh-Nhu Numi Tran, Graham Low, and Ghassan Beydoun. A methodological framework for ontology centric oriented software engineering. International Journal of Computer Systems Science and Engineering, 21(2): ??, March 2006. CODEN CSSEEI. ISSN 0267-6192.

Tohma:1994:AHG

[TM94] Yoshiro Tohma and Yukichi Matsunaga. Application of hyper-geometric distribution model to the hardware debugging process. Computer Systems Science and Engineering, 9(1):25–30, January 1994. CODEN CSSEEI. ISSN 0267-6192.

Toval:2008:EKI

[TMNL08] Ambrosio Toval, Begoña Moros, Joaquín Nicolás, and Joaquín Lasheras. Eight key issues for an effective reuse-based requirements process. International Journal of Computer Systems Science and Engineering, 23(6): ??, November 2008. CODEN CSSEEI. ISSN 0267-6192.

Thomasian:1993:SRT

[TN93] Alexander Thomasian and Behzad Nadji. State representation tradeoffs in Markov chains of serialization delays in computer systems. Computer Systems Science and Engineering, 8(3):154–165, July 1993. CODEN CSSEEI. ISSN 0267-6192.

Tohma:1994:RFT

[Toh94] Yoshiro Tohma. Review of fault-tolerant computing for safely critical applications in Japan. Computer Systems Science and Engineering, 9 (1):3–10, January 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Troger:2009:OPM

[TP09] Peter Tröger and Andreas Polze. Object and process migration in .NET. International Journal of Computer Systems Science and Engineering, 24(1):??, January 2009. CODEN CSSEEI. ISSN 0267-6192.

Taniar:2002:PGQ

[TR02a] D. Taniar and J. Wenny Rahayu. Parallel group-by query processing in a cluster architecture. International Journal of Computer Systems Science and Engineering, 17(1): ??, January 2002. CODEN CSSEEI. ISSN 0267-6192.

Taniar:2002:PSM

[TR02b] D. Taniar and J. Wenny Rahayu. Parallel sort-merge object-oriented collection join algorithms. International Journal of Computer Systems Science and Engineering, 17 (3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Thyagharajan:2006:OBR

[TR06] K. K. Thyagharajan and V. Ramachandran. Optimal buffering requirement analysis for jitter-free variable bit rate video streaming. International Journal of Computer Systems Science and Engineering, 21(3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Torres-Rojas:2002:RTB

[TRAR02] F. J. Torres-Rojas, M. Ahamad, and M. Raynal. Real-time based strong consistency for distributed objects. International Journal of Computer Systems Science and Engineering, 17(2):??, March 2002. CODEN CSSEEI. ISSN 0267-6192.

Taniar:2010:MHN

[TRLD10] David Taniar, Wenny Rahayu, Vincent Lee, and Olena Daly. Mining hierarchical negative association rules. International Journal of Computer Systems Science and Engineering, 25(2):??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Taniar:2004:PAS

[TRT04] D. Taniar, J. W. Rahayu, and R. Boon-Noi Tan. Parallel algorithms for selection query processing involving index in parallel database systems. International Journal of Computer Systems Science and Engineering, 19(2):??, March 2004. CODEN CSSEEI. ISSN 0267-6192.

Tanca:1999:SIS

[TS99] Letizia Tanca and Dan Suciu. Special issue: Semantics on the World Wide Web. International Journal of Computer Systems Science and Engineering, 14(2):??, March 1999. CODEN CSSEEI. ISSN 0267-6192.

Taboada:2010:PAM

[TTD10] Guillermo L. Taboada, Juan Touriño, and Ramón Doallo.

Performance analysis of messagepassing libraries on highspeed clusters. International Journal of Computer Systems Science and Engineering, 25 (1):??, January 2010. CO-DEN CSSEEI. ISSN 0267-6192.

Tupakula:2008:HMA

[TV08] Udaya Kiran Tupakula and Vijay Varadharajan. A hybrid model against TCP SYN and reflection DDoS attacks. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Thomasian:2016:DAH

[TX16] Alexander Thomasian and Jun Xu. Data allocation in a Heterogeneous Disk Array (HDA) with multiple RAID levels for database applications. International Journal of Computer Systems Science and Engineering, 31(5): ??, September 2016. CODEN CSSEEI. ISSN 0267-6192.

Tan:1997:FSP

[TY97] Kian-Lee Tan and Jeffrey X. Yu. Framework for scheduling partition-based joins. International Journal of Computer Systems Science and Engineering, 12(4):263–273, July 1997. CODEN CSSEEI. ISSN 0267-6192.

Tang:2009:CQM

[TYW09] Siliang Tang, Jianhua Yang, and Zhaohui Wu. A context quality management infrastructure for complex ubiquitous environment. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Tang:2012:GTM

[TYY^{+}12] Feilong Tang, Ilsun You, Shui Yu, Huakang Li, and Cho-Li Wang. Grid transaction management and an efficient development kit. International Journal of Computer Systems Science and Engineering, 27 (5):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Tang:2009:PSP

[TZC09] Yuxin Tang, Yunquan Zhang, and Hu Chen. A parallel shortest path algorithm based on graph-partitioning and iterative correcting. International Journal of Computer Systems Science and Engineering, 24(5):??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Tse:1996:PFI

[TZL96] Philip W. Tse, Moshe Zuckerman, and François Le Faucheur. Performance and fairness issues related to the EFCI rate-based flow control for the ABR service in ATM networks. International Journal of Computer Systems Science and Engineering, 11 (6):383–392, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Ulieru:2006:PSS

[UI06] Mihaela Ulieru and Dan Ionescu. Privacy and security shield for health information systems (e-health). International Journal of Computer Systems Science and Engineering, 21(3):??, May 2006. CODEN CSSEEI. ISSN 0267-6192.

Ugarte-Pedrero:2013:CCP

[UPSL^{+}13] Xabier Ugarte-Pedrero, Igor Santos, Carlos Laorden, Borja Sanz, and Pablo G. Bringas. Collective classification for packed executable identification. International Journal of Computer Systems Science and Engineering, 28(1):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Varadharajan:1987:NBS

[VB87] V. Varadharajan and K. D. Baker. Net-based system design representation and analysis. Computer Systems Science and Engineering, 2(4): 167–178, October 1987. CO-DEN CSSEEI. ISSN 0267-6192.

vanderAalst:2001:HHD

[vdA01] W. M. P. van der Aalst. How to handle dynamic change and capture management information? an approach based on generic workflow models. International Journal of Computer Systems Science and Engineering, 16(5): ??, September 2001. CODEN CSSEEI. ISSN 0267-6192.

vanderAalst:2000:DWC

[vdAJ00a] W. M. P. van der Aalst and S. Jablonski. Dealing with workflow change: identification of issues and solutions. International Journal of Computer Systems Science and Engineering, 15(5): ??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

vanderAalst:2000:E

[vdAJ00b] Wil van der Aalst and Stefan Jablonski. Editorial. International Journal of Computer Systems Science and Engineering, 15(5):??, September 2000. CODEN CSSEEI. ISSN 0267-6192.

Vardanega:2000:LLI

[VG00] T. Vardanega and J. Gaisler.

Lessons learned from the implementation of on-board tolerance to physical faults in Ada. International Journal of Computer Systems Science and Engineering, 15(1): 19–??, January 2000. CODEN CSSEEI. ISSN 0267-6192.

Vo:2008:SMA

[VK08] Duc-Liem Vo and Kwangjo Kim. A secure mutual authentication scheme with key agreement using smart card from bilinear pairings. International Journal of Computer Systems Science and Engineering, 23(3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Varadhaganapathy:2015:HBI

[VKAR15] S. Varadhaganapathy, V. Krishnaveni, G. Arumugam, and R. R. Rajalaxmi. Harmony and bio inspired harmony search optimization algorithms for feature selection in classification. International Journal of Computer Systems Science and Engineering, 30 (4):??, July 2015. CODEN CSSEEI. ISSN 0267-6192.

Vlachogiannis:2007:ACD

[Vla07] John G. Vlachogiannis. Automatic construction of a decision tree for the analysis of traffic load in computer networks. International Journal of Computer Systems Science and Engineering, 22(1–2):??, January/March 2007. CO-DEN CSSEEI. ISSN 0267-6192.

Vieira:1998:OST

[VM98] Sibelius Lellis Vieira and Mauricio Ferreira Magalhães. Online sporadic task scheduling in hard real-time systems. International Journal of Computer Systems Science and Engineering, 13(4): 249–258, July 1998. CODEN CSSEEI. ISSN 0267-6192.

vonRenteln:2011:IEA

[vRB11] Alexander von Renteln and Uwe Brinkschulte. Implementing and evaluating the AHS organic middleware. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Vargas-Solar:2008:IQA

[VSZMCV08] G. Vargas-Solar, J-L Zechinelli-Martini, and V. Cuevas-Vicenttín. Integrating and querying astronomical data on the e-GrOV data grid. International Journal of Computer Systems Science and Engineering, 23(2):??, March 2008. CODEN CSSEEI. ISSN 0267-6192.

Vuong:1995:SET

[VT95] Son T. Vuong and George K.

Tsiknis. Significant event temporal logic: a method for protocol specification and verification. International Journal of Computer Systems Science and Engineering, 10(1): 41–49, January 1995. CO-DEN CSSEEI. ISSN 0267-6192.

Wallace:1990:PRG

[Wal90] C. S. Wallace. Physically random generator. Computer Systems Science and Engineering, 5(2):82–88, April 1990. CODEN CSSEEI. ISSN 0267-6192.

Wang:1997:APC

[Wan97]  Hu-Jun Wang.  Analysis

of prioritized cluster-based crossbar systems. International Journal of Computer Systems Science and Engineering, 12(4):255–261, July 1997. CODEN CSSEEI. ISSN 0267-6192.

Wang:2003:ELA

[Wan03] J.-L. Wang. An effective link arrangement scheme for distributed systems using Markov-chain based reliability analysis. International Journal of Computer Systems Science and Engineering, 18 (6):??, December 2003. CO-DEN CSSEEI. ISSN 0267-6192.

Ward:2007:BIA

[War07] Jeremy Ward. Building the information assurance community of purpose. International Journal of Computer Systems Science and Engineering, 22(5):??, September 2007. CODEN CSSEEI. ISSN 0267-6192.

Wang:1994:PAC

[WBA94] Hongying Wang, Douglas M. Blough, and Leon Alkalaj. Practical approach to comparison-based fault diagnosis in multiprocessor systems. Computer Systems Science and Engineering, 9(1): 11–20, January 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Won:2013:ODS

[WBY^{+}13] Jun-Im Won, Hyun Back, Jee-Hee Yoon, Sanghyun Park, and Sang-Wook Kim. Onedimensional spatial join processing using a DOT-based index structure. International Journal of Computer Systems Science and Engineering, 28 (2):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Wu:2001:CKA

[WC01] T.-C. Wu and C.-C. Chang.

Cryptographic key assignment scheme for hierarchical access control. International Journal of Computer Systems Science and Engineering, 16 (1):??, January 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Wang:2008:ATS

[WC08a] Jin-Long Wang and Chen-Hsing Chen. Adaptive two-stage QoS provisioning schemes for CDMA networks. International Journal of Computer Systems Science and Engineering, 23(1): ??, January 2008. CODEN CSSEEI. ISSN 0267-6192.

Wang:2008:CID

[WC08b] S. Y. Wang and C. L.

Chou. On the characteristics of information dissemination paths in vehicular ad hoc networks on the move. International Journal of Computer Systems Science and Engineering, 23(6):??, November 2008. CODEN CSSEEI. ISSN 0267-6192.

Wang:1999:MFC

[WCD99] Jin Chin Wang, Jacqueline Chame, and Michel Dubois. Modeling finite cache effects in write-invalidate snooping protocols. International Journal of Computer Systems Science and Engineering, 14(1): 15–25, January 1999. CO-DEN CSSEEI. ISSN 0267-6192.

Weidong:2009:NDT

[WCJ09] Liu Weidong, Lin Chuang, and Wang Jirong. A novel data transmission scheme for mobile sensor network. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Wang:2015:RBC

[WCL15] Ping Wang, Kuo-Ming Chao, and Chi-Chun Lo. A reputation-based computing approach in trusted web service provisions. International Journal of Computer Systems Science and Engineering, 30 (4):??, July 2015. CODEN CSSEEI. ISSN 0267-6192.

Wang:2004:AHR

[WCLC04] Y.-H. Wang, C. M. Chung, C.-J. Lee, and C.-C. Chuang. Ad hoc routing protocol based on setup on-demand backup node. International Journal of Computer Systems Science and Engineering, 19(2): ??, March 2004. CODEN CSSEEI. ISSN 0267-6192.

Wang:1990:PCC

[WD90] Jin-Chin Wang and Michel Dubois. Performance comparison of cache coherence protocols based on the access burst model. Computer Systems Science and Engineering, 5(3):147–158, July 1990. CODEN CSSEEI. ISSN 0267-6192.

Wong:1987:CRT

[WDF87] C. Y. Wong, T. S. Dillon, and K. E. Forward. Concurrent, real-time systems: a systematic approach using timed Petri nets. Computer Systems Science and Engineering, 2(3):117–124, July 1987. CODEN CSSEEI. ISSN 0267-6192.

Wei:1995:CRN

[Wei95] Wei Wei. Constructing reliable neural network systems using less reliable neuron network components. International Journal of Computer Systems Science and Engineering, 10(1):28–32, January 1995. CODEN CSSEEI. ISSN 0267-6192.

Wu:1992:CCA

[WF92] Jie Wu and Eduardo B. Fernandez. Comparison connection assignments for diagnosis of multiprocessor systems under a two-fault assumption. Computer Systems Science and Engineering, 7 (3):199–201, July 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Wu:1993:FTD

[WF93] Jie Wu and Eduardo B. Fernandez. Fault-tolerant distributed broadcast algorithm for cube-connected cycles. Computer Systems Science and Engineering, 8(4): 224–233, October 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Wolf:2011:RSE

[WGK^{+}11] Julian Wolf, Mike Gerdes, Florian Kluge, Sascha Uhrig, Jörg Mische, Stefan Metzlaff, Christine Rochange, Hugues Cassé, Pascal Sainrat, and Theo Ungerer. RTOS support for execution of parallelized hard real-time tasks on the MERASA multi-core processor. International Journal of Computer Systems Science and Engineering, 26(6): ??, November 2011. CODEN CSSEEI. ISSN 0267-6192.

Wu:1997:DSA

[WH97] Tzong-Chen Wu and Bor-Ron Hwang. On the design of a secure anonymous conference key distribution system. International Journal of Computer Systems Science and Engineering, 12(4):239– 244, July 1997. CODEN CSSEEI. ISSN 0267-6192.

Wang:2011:EEB

[WH11] Guoli Wang and Chao Huang.

Energy-efficient beaconless real-time routing protocol for wireless sensor networks. International Journal of Computer Systems Science and Engineering, 26(3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Wallace:1985:TCM

[WK85] Christopher S. Wallace and David Koch. TTL-compatible multiport bus. Computer Systems Science and Engineering, 1(1):47–52, October 1985. CODEN CSSEEI. ISSN 0267-6192.

Wang:2007:FAI

[WKT07] L. Wang, L. Khan, and B. Thuraisingham. A framework for automated image annotation. International Journal of Computer Systems Science and Engineering, 22(1– 2):??, January/March 2007. CODEN CSSEEI. ISSN 0267-6192.

Wang:1991:PAC

[WL91] Chihping Wang and Victor O. K. Li. Precedenceagreement concurrency control algorithm for distributed database systems. Computer Systems Science and Engineering, 6(4):227–237, October 1991. CODEN CSSEEI. ISSN 0267-6192.

Wang:2001:EBA

[WL01] J.-L. Wang and L.-T. Lee. Effective bandwidth allocation for ATM networks. International Journal of Computer Systems Science and Engineering, 16(1):??, January 2001. CODEN CSSEEI. ISSN 0267-6192.

Wang:1993:OSS

[WLC93] Chhping Wang, Victor O. K. Li, and Arbee L. P. Chen. One-shot semi-join execution strategies for processing distributed join queries. Computer Systems Science and Engineering, 8(4):245– 253, October 1993. CODEN CSSEEI. ISSN 0267-6192.

Wu:2007:IFR

[WLC07] C.-L. Wu, D.-C. Lou, and S.-Y. Chen. Integer factorization for RSA cryptosystem under a PVM environment. International Journal of Computer Systems Science and Engineering, 22(1–2):??, January/March 2007. CO-DEN CSSEEI. ISSN 0267-6192.

Wei:2011:LPV

[WLG^{+}11] Xiaohui Wei, Hongliang Li, Qingnan Guo, Na Jiang, and Liang Hu. LimeVI: A platform for virtual cluster live migration over WAN. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Wang:2015:QOT

[WLXD15] Tingting Wang, Weijiang Liu, Yujie Xu, and Mianxiong Dong. QoS oriented task scheduling based on generic algorithm in cloud computing. International Journal of Computer Systems Science and Engineering, 30(6): ??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Wang:2014:CDC

[WLY^{+}14] Zhixiao Wang, Yanxiao Liu, Wenyao Yan, Wei Wei, Junhuai Li, Deyun Zhang, and Omar Alfandi. Cheating detection and cheater identification in (t, n) secret sharing scheme. International Journal of Computer Systems Science and Engineering, 29(1): ??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Woods:1995:SET

[WMFK95] W. A. Woods, H. D. Moser, O. Frieder, and Paul B. Kantor. Simulation evaluation of task scheduling on reconfigurable multicomputer architectures. International Journal of Computer Systems Science and Engineering, 10(4): 244–254, October 1995. CO-DEN CSSEEI. ISSN 0267-6192.

Weiss:2006:MDF

[WOE06] M. Weiss, A. Oreshkin, and B. Esfandiari. Method for detecting functional feature of interactions of Web services. International Journal of Computer Systems Science and Engineering, 21(4): ??, July 2006. CODEN CSSEEI. ISSN 0267-6192.

Wang:1995:IIR

[WP95] Tao Wang and Wuxu Peng. Improving Internet routing information protocol. International Journal of Computer Systems Science and Engineering, 10(4):207–213, October 1995. CODEN CSSEEI. ISSN 0267-6192.

Wang:1993:AIL

[WSL93] Jing-Long Wang, Meng-Hong Shie, and Liang-Teh Lee. Asymmetrically-initiated load balancing policy for distributed systems. Computer Systems Science and Engineering, 8(4):187–194, October 1993. CODEN CSSEEI. ISSN 0267-6192.

Waluyo:2005:RLD

[WST05] Agustinus Borgy Waluyo, Bala Srinivasan, and David Taniar. Research on locationdependent queries in mobile databases. International Journal of Computer Systems Science and Engineering, 20 (2):??, March 2005. CODEN CSSEEI. ISSN 0267-6192.

Wei:2015:IRM

[WSZG15] Yan Wei, Shi Songshan, Liu Zhenggang, and Li Guoxiang.

Investigation on remote monitoring system for hybrid electric bulldozer. International Journal of Computer Systems Science and Engineering, 30 (5):??, September 2015. CO-DEN CSSEEI. ISSN 0267-6192.

Wu:2001:DSM

[Wu01] T.-C. Wu. Digital signature/ multisignature schemes giving public key verification and message recovery simultaneously. International Journal of Computer Systems Science and Engineering, 16(6): ??, November 2001. CODEN CSSEEI. ISSN 0267-6192.

Wu:2016:EED

[Wu16] Jun Wu. Energy efficient dual execution mode scheduling for real-time tasks with shared resources. International Journal of Computer Systems Science and Engineering, 31(3):??, May 2016. CODEN CSSEEI. ISSN 0267-6192.

Wong:2001:CAE

[WW01] A. K. Y. Wong and J. H. C.

Wong. A convergence algorithm for enhancing the performance of distributed applications running on sizeable networks. International Journal of Computer Systems Science and Engineering, 16 (4):??, July 2001. CODEN CSSEEI. ISSN 0267-6192.

Wu:2004:RNR

[WWD04] Richard S. L. Wu, Allan K. Y. Wong, and Tharam S. Dillon. RDCT: a novel reconfigurable dynamic cache size tuner to shorten information retrieval time over the Internet. International Journal of Computer Systems Science and Engineering, 19(6): ??, November 2004. CODEN CSSEEI. ISSN 0267-6192.

Wong:2009:NAA

[WWDL09] Jackei H. K. Wong, Allan K. Y. Wong, Tharam Dillon, and Wilfred W. K. Lin. A novel approach to achieve real-time TCM (traditional Chinese medicine) telemedicine through the use of ontology and clinical intelligence discovery. International Journal of Computer Systems Science and Engineering, 24 (4):??, July 2009. CODEN CSSEEI. ISSN 0267-6192.

Wu:1995:DAC

[WWH95] Tzong-Chen Wu, Tzong-Sun Wu, and Wei-Hua He. Dynamic access control scheme based on the Chinese remainder theorem. International Journal of Computer Systems Science and Engineering, 10 (2):92–99, April 1995. CO-DEN CSSEEI. ISSN 0267-6192.

Wei:2016:PRR

[WWS16] Zhe Wei, Fang Wang, and Le Sun. A practical reliable reputation based data fusion method for ad-hoc wireless sensor networks. International Journal of Computer Systems Science and Engineering, 31(3):??, May 2016. CODEN CSSEEI. ISSN 0267-6192.

Wu:1993:CSB

[WY93] Tzong-Chen Wu and Yi-Shiung Yeh. Cryptosystem for selectively broadcasting separate secrets. Computer Systems Science and Engineering, 8(2):121–124, April 1993. CODEN CSSEEI. ISSN 0267-6192.

Wong:1996:RFE

[WY96] Allan K. Y. Wong and Daniel S. Yeung. RHS — a framework for exploiting distributed parallelism efficiently. International Journal of Computer Systems Science and Engineering, 11(3): 177–184, May 1996. CODEN CSSEEI. ISSN 0267-6192.

Wu:2002:CRH

[WY02] K-L Wu and P. S. Yu.

Controlled replication for hash routing-based Web caching. International Journal of Computer Systems Science and Engineering, 17(4/5):??, July/September 2002. CODEN CSSEEI. ISSN 0267-6192.

Wang:2016:CRM

[WYLW16] En Wang, Yongjian Yang, Wenbin Liu, and Fan Wang.

Clustering routing method based on social characteristic in campus DTN. International Journal of Computer Systems Science and Engineering, 31(2):??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Wakahara:2012:IPD

[WYO12] Toshihiko Wakahara, Noriyasu Yamamoto, and Hiroki Ochi. Image processing of dotted picture in QR code of cellular phones. International Journal of Computer Systems Science and Engineering, 27 (1):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Wu:2001:TCA

[WYT01] K.-L. Wu, P. S. Yu, and J. Z. Teng. Thrashing control and avoidance for concurrent mergesorts using parallel prefetching. International Journal of Computer Systems Science and Engineering, 16 (6):??, November 2001. CO-DEN CSSEEI. ISSN 0267-6192.

Wei:2015:BCS

[WZWD15] Xiao Wei, Daniel Dajun Zeng, Wei Wu, and Yeming Dai. Building the concept semantic space for large text database. International Journal of Computer Systems Science and Engineering, 30(5): ??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Xu:2007:MCB

[XC07] Haiping Xu and Yi-Tsung Cheng. Model checking bidding behaviors in Internet concurrent auctions. International Journal of Computer Systems Science and Engineering, 22(4):??, July 2007. CODEN CSSEEI. ISSN 0267-6192.

Xu:2015:SAK

[XC15] Zheng Xu and Haiyan Chen.

The semantic analysis of knowledge map for the traffic violations from the surveillance video big data. International Journal of Computer Systems Science and Engineering, 30(5):??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Xue:2011:MSI

[XCHY11] Gang Xue, Tony Chen, Gang He, and Shaowen Yao. Modeling service interaction patterns. International Journal of Computer Systems Science and Engineering, 26(2): ??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Xie:1999:PDC

[XFA99] H. A. Xie, K. E. Forward, and K. M. Adams. A programmable database coprocessor in a PSQL system. International Journal of Computer Systems Science and Engineering, 14(5):299– ??, 1999. CODEN CSSEEI. ISSN 0267-6192.

Xiang:2009:PAW

[XFZ09] Yang Xiang, Xiang Fan, and Wen Tao Zhu. Propagation of active worms: A survey. International Journal of Computer Systems Science and Engineering, 24(3):??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Xu:2016:APM

[XLH16] Yangyang Xu, Zhaobin Liu, and Zhonglian Hu. Ascending partition method for differentially private histogram publication. International Journal of Computer Systems Science and Engineering, 31(2): ??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Xu:2010:EEP

[XLSM10] Junfeng Xu, Kequi Li, Yanming Shen, and Geyong Min. An energy efficient packet aggregation scheme for wireless sensor networks. International Journal of Computer Systems Science and Engineering, 25(6):??, November 2010. CODEN CSSEEI. ISSN 0267-6192.

Xu:2011:IBA

[XLW11] Zheng Xu, Xiangfeng Luo, and Lizhe Wang. Incremental building association link network. International Journal of Computer Systems Science and Engineering, 26(3): ??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Xiong:1997:PMA

[XM97] Yijun Xiong and Lorne Mason. Performance of multistage ATM switching networks with multicast capability. International Journal of Computer Systems Science and Engineering, 12(2):103– 115, March 1997. CODEN CSSEEI. ISSN 0267-6192.

Xu:2011:MAB

[XQZ11] Baoshu Xu, Wenyu Qu, and Wanlei Zhou. A mobile agentbased routing algorithm and some theoretical analysis. International Journal of Computer Systems Science and Engineering, 26(1):??, January 2011. CODEN CSSEEI. ISSN 0267-6192.

Xie:2002:DCL

[XS02] X. Xie and S. M. Shatz. Development of class-level and instance-level design models for distributed-object software. International Journal of Computer Systems Science and Engineering, 17(3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Xu:2016:SIC

[XSZ16] Guest Editors: Zheng Xu, Vijayan Sugumaran, and Hui Zhang. Special issue: Cyber-physical systemrelated technologies and applications. International Journal of Computer Systems Science and Engineering, 31 (6):??, November 2016. CO-DEN CSSEEI. ISSN 0267-6192.

Xiao:2011:HLM

[XZ11] Jing Xiao and Min Zhu. HyperMonitor: A lightweight multi-platform monitor based on hardware virtualization. International Journal of Computer Systems Science and Engineering, 26(5):??, September 2011. CODEN CSSEEI. ISSN 0267-6192.

Yuan:1997:ECS

[YA97] Shyan-Ming Yuan and Ashok K. Agrawala. An efficient communication structure for decentralized algorithms with fault tolerance. International Journal of Computer Systems Science and Engineering, 12 (6):343–349, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Yang:1993:MCT

[Yan93] Chyan Yang. Multigauge computers and their applications. Computer Systems Science and Engineering, 8(1): 33–39, January 1993. CO-DEN CSSEEI. ISSN 0267-6192.

Yang:2000:EAF

[Yan00] S.-J. Yang. Efficient algorithms for finding cores on weighted tree networks. International Journal of Computer Systems Science and Engineering, 15(4):233–??, July 2000. CODEN CSSEEI. ISSN 0267-6192.

Yang:2002:DAE

[Yan02] S-J Yang. A distributed algorithm of efficient core discipline on weighted bipartite networks. International Journal of Computer Systems Science and Engineering, 17(1): ??, January 2002. CODEN CSSEEI. ISSN 0267-6192.

You:2012:Ea

[YBT12] Ilsun You, Leonard Barolli, and Feilong Tang. Editorial. International Journal of Computer Systems Science and Engineering, 27(5): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Ye:2013:BMM

[YC13] Lei Ye and Henry C. B. Chan.

B2B magics: a mobile agentbased Internet commerce system for B2B e-commerce. International Journal of Computer Systems Science and Engineering, 28(2):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Yiyu:2009:IFS

[YCFX09] Tan Yiyu, Yau Chihang, Anthony S. Fong, and Yang Xiaojian. An instruction folding solution for a Java processor. International Journal of Computer Systems Science and Engineering, 24(3): ??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Yak:1987:MET

[YDF87] Y. W. Yak, T. S. Dillon, and K. E. Forward. Modelling the effect of transient faults in fault-tolerant computer systems. Computer Systems Science and Engineering, 2(4): 161–166, October 1987. CO-DEN CSSEEI. ISSN 0267-6192.

Yin:2009:SSV

[YDR^{+}09] Hao Yin, ChangLai Du, Chao Ren, Zhijia Chen, Heyong Min, and Chuang Lin. A secure and scalable video conference system based on peerassisted content delivery networks. International Journal of Computer Systems Science and Engineering, 24(5): ??, September 2009. CODEN CSSEEI. ISSN 0267-6192.

Ye:2015:SGS

[Ye15] Xinfeng Ye. Security games for service computing applications. International Journal of Computer Systems Science and Engineering, 30(6): ??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Yen:2008:RID

[Yen08] William Chung-Kung Yen. Restricted independent domination problems on graphs. International Journal of Computer Systems Science and Engineering, 23(6):??, November 2008. CODEN CSSEEI. ISSN 0267-6192.

Yoneda:1994:IFT

[YET94] Tomohiro Yoneda, Toru Egashira, and Yoshihiro Tohma. Improving a faulttolerant clock synchronization algorithm by overcorrection. Computer Systems Science and Engineering, 9(1): 54–64, January 1994. CO-DEN CSSEEI. ISSN 0267-6192.

Yuan:1997:DID

[YH97] Shyan-Ming Yuan and Yu-Kwen Hsu. Design and implementation of a distributed monitor facility. International Journal of Computer Systems Science and Engineering, 12 (1):43–51, January 1997. CO-DEN CSSEEI. ISSN 0267-6192.

Ye:2013:BBC

[YHLC13] Lei Ye, Adam S. K. Ho, Solomon K. S. Leung, and Henry C. B. Chan. A business-to-business e-commerce system using software agents and RFID. International Journal of Computer Systems Science and Engineering, 28 (3):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

Yoo:2011:FNS

[YLK11] Sanghyun Yoo, Ki Yong Lee, and Myoung Ho Kim. Fast k-NN search using precomputed l-NN sets. International Journal of Computer Systems Science and Engineering, 26(4):??, July 2011. CODEN CSSEEI. ISSN 0267-

Yue:2013:DSA

[YLQW13] Kun Yue, Wei-Yi Liu, Wen-Hua Qian, and Xiao-Feng Wang. Discovering semantic associations between Web services based on the electrostatic force theory. International Journal of Computer Systems Science and Engineering, 28(5):??, ???? 2013. CODEN CSSEEI. ISSN 0267-6192.

You:2012:DDS

[YLS12] Ilsun You, Jong-Hyouk Lee, and Kouichi Sakurai. DSSH: Digital signature based secure handover for network-based mobility management. International Journal of Computer Systems Science and Engineering, 27(3):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Yue:2011:AKE

[YLZ11] Kun Yue, Wei-Yi Liu, and Li-Ping Zhou. Automatic keyword extraction from documents based on multiple content-based measures. International Journal of Computer Systems Science and Engineering, 26(2):??, March 2011. CODEN CSSEEI. ISSN 0267-6192.

Yen:1998:EOR

[YMST98] Cheng-Huang Yen, L. L. Miller, A. Sirjani, and J. Tenner. Extending the objectrelational interface to support an extensible view system for multidatabase integration and interoperation. International Journal of Computer Systems Science and Engineering, 13 (4):227–240, July 1998. CO-DEN CSSEEI. ISSN 0267-6192.

Yang:2012:ESW

[YOM^{+}12] Tao Yang, Tetsuya Oda, Gjergji Mino, Leonard Barolli, Arjan Durresi, and Fatos Xhafa. Energy-saving in wireless sensor networks considering mobile sensor nodes, single and multi events. International Journal of Computer Systems Science and Engineering, 27(5):??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Young:2001:NMM

[You01] M. Young. Neat models of messy problems: Notes on the interplay between solutionand problem-centered disciplines, and more particularly on the interaction between Petri net research and software engineering research. International Journal of Computer Systems Science and Engineering, 16(3):??, May 2001. CODEN CSSEEI. ISSN 0267-6192.

Yang:1995:RAE

[YQ95] Cui-Qing Yang and Yaoshuang Qu. Reliability analysis for the execution of remote jobs in a workstation-based environment. International Journal of Computer Systems Science and Engineering, 10(2): 120–128, April 1995. CODEN CSSEEI. ISSN 0267-6192.

Yang:1990:FTM

[YS90] Shih-Chian Yang and John A. Silvester. Fault-tolerant multistage interconnection networks: performance/ reliability tradeoffs. Computer Systems Science and Engineering, 5(4):233–242, October 1990. CODEN CSSEEI. ISSN 0267-6192.

Youssef:1999:TDG

[YS99] Habib Youssef and Sadiq M. Sait. Timing-driven global routing for standard-cell VLSI design. International Journal of Computer Systems Science and Engineering, 14(3): 175–185, May 1999. CODEN CSSEEI. ISSN 0267-6192.

You:2012:Eb

[YS12] Ilsun You and JinShu Su. Editorial. International Journal of Computer Systems Science and Engineering, 27(6): ??, ???? 2012. CODEN CSSEEI. ISSN 0267-6192.

Youssef:1996:HLS

[YSAMB96] Habib Youssef, Sadiq M. Sait, A. S. Al-Mulhelm, and M. S. T. Benten. Highlevel synthesis from purely behavioral descriptions. International Journal of Computer Systems Science and Engineering, 11(5):259–273, September 1996. CODEN CSSEEI. ISSN 0267-6192.

Yang:2016:TRE

[YSS^{+}16] Tingting Yang, Zhonghua Sun, Zhou Su, Hailong Feng, and Ruilong Deng. Towards risk evaluation algorithm of PSC ship-selecting system based on hierarchic analysis and grey integrated clustering method. International Journal of Computer Systems Science and Engineering, 31(2): ??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Yamagiwa:2010:RMU

[YUM10] Motoi Yamagiwa, Minoru Uehara, and Makoto Murakami. A reconstruction method for ultrasonic deterioration image by the combination of constructive solid geometry and strongly typed genetic programming. International Journal of Computer Systems Science and Engineering, 25 (2):??, March 2010. CODEN CSSEEI. ISSN 0267-6192.

Yeung:1998:EIF

[YW98a] Daniel S. Yeung and A. K. Y.

Wong. An evaluation identifying features of unified distributed programming. International Journal of Computer Systems Science and Engineering, 13(5):311–321, September 1998. CODEN CSSEEI. ISSN 0267-6192.

Yeung:1998:OCF

[YW98b] Daniel S. Yeung and Allan K. Y. Wong. OORHS: a conceptual framework that provides easy and reversible distributed programming. International Journal of Computer Systems Science and Engineering, 13(5):289–301, September 1998. CODEN CSSEEI. ISSN 0267-6192.

Yu:2000:DPG

[YWD00] P. S. Yu, K.-L. Wu, and A. Dan. A dynamic parity grouping for efficient parity buffering for RAID-5 disk arrays. International Journal of Computer Systems Science and Engineering, 15(3): 155–??, May 2000. CODEN CSSEEI. ISSN 0267-6192.

Yu:2003:AMN

[YWL^{+}03] J. X. Yu, G. Wang, H. Lu, G. Yu, J. Lv, and B. Sun. Automata match: a new XML query processing approach. International Journal of Computer Systems Science and Engineering, 18(5): ??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Yang:2006:EAP

[YWS06] Zhiqiang Yang, Rebecca N. Wright, and Hiranmayee Subramaniam. Experimental analysis of a privacypreserving scalar product protocol. International Journal of Computer Systems Science and Engineering, 21(1): ??, January 2006. CODEN CSSEEI. ISSN 0267-6192.

Yuan:1996:SXW

[YWW96] Shyan-Ming Yuan, Jeng-Dar Wu, and Dong-Lieng Wang. SHDS — an X-Window based semi-hypermedia document system. International Journal of Computer Systems Science and Engineering, 11(1): 55–64, January 1996. CO-DEN CSSEEI. ISSN 0267-6192.

You:1997:PIM

[YZP97] J. You, W. P. Zhu, and E. Pissaloux. Parallel image matching on a distributed system. International Journal of Computer Systems Science and Engineering, 12(6):381– 385, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2015:HHI

[ZCG^{+}15] Aiping Zhou, Guang Cheng, Xiaojun Guo, Dinh Tu Truong, and Chengang Zhu. Heavy hitter identification based on adaptive sampling with mapreduce. International Journal of Computer Systems Science and Engineering, 30(6):??, November 2015. CODEN CSSEEI. ISSN 0267-6192.

Zang:2011:MDS

[ZCK11] Tianyi Zang, Radu Calinescu, and Marta Kwiatkowska. Metamodel-driven SOA for collaborative e-science application. International Journal of Computer Systems Science and Engineering, 26(3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2003:FSA

[ZCL03] H. Zhou, W. Cao, and Z. Lin. Formal semantics analysis for VHDL RTL synthesis. International Journal of Computer Systems Science and Engineering, 18(6):??, December 2003. CODEN CSSEEI. ISSN 0267-6192.

Zeng:2007:BDC

[ZD07] Qingtian Zeng and Hua Duan. Behavior description for complex flexible manufacturing system based on decomposition of Petri net. International Journal of Computer Systems Science and Engineering, 22(6):??, November 2007. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2011:ASI

[ZEZF11] Peng Zhang, Sabu Emmanuel, Yanning Zhang, and Cheng Fu. Auto-scaled ISL tracking for region based control infrastructure and applications in video surveillance. International Journal of Computer Systems Science and Engineering, 26(3):??, May 2011. CODEN CSSEEI. ISSN 0267-6192.

Zowghi:2005:E

[ZG05] Didar Zowghi and Vincenzo Gervasi. Editorial. International Journal of Computer Systems Science and Engineering, 20(1):??, January 2005. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2016:SSM

[Zha16] Cui-Xia Zhang. The state space model of remanufacturing carbon footprint for lowcarbon. International Journal of Computer Systems Science and Engineering, 31(6): ??, November 2016. CODEN CSSEEI. ISSN 0267-6192.

Zhenhua:2003:BTS

[Zhe03] L. Zhenhua. A bus-type software model based on CORBA and XML. International Journal of Computer Systems Science and Engineering, 18 (6):??, December 2003. CO-DEN CSSEEI. ISSN 0267-6192.

Zhu:1996:FPA

[Zhu96] Yahui Zhu. Fast processor allocation and dynamic scheduling for mesh multiprocessors. International Journal of Computer Systems Science and Engineering, 11(2): 99–107, March 1996. CODEN CSSEEI. ISSN 0267-6192.

Zhao:2015:RDP

[ZHZ^{+}15] Chihang Zhao, Jie He, Xiaoqin Zhang, Xingzhi Qi, and Aiwen Chen. Recognition of driving postures by nonsubsampled contourlet transform and k-nearest neighbor classifier. International Journal of Computer Systems Science and Engineering, 30(3):

??, May 2015. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2002:TBI

[ZJ02] W. Zhou and W. Jia. A token-based independent update protocol for managing replicated objects. International Journal of Computer Systems Science and Engineering, 17(3):??, May 2002. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2014:MSD

[ZKWM14] Guan-Wen Zhang, Jien Kato, Yu Wang, and Kenji Mase. Multi-stage deep convolutional learning for people reidentification. International Journal of Computer Systems Science and Engineering, 29 (4):??, ???? 2014. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2015:ANN

[ZL15] Qingyuan Zhou and Jianjian Luo. Artificial neural network based grid computing of e-governance scheduling for emergency management. International Journal of Computer Systems Science and Engineering, 30(5): ??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2016:SQE

[ZL16] Qingyuan Zhou and Juan Luo. The service quality evaluation of ecologic economy systems using simulation computing. International Journal of Computer Systems Science and Engineering, 31 (6):??, November 2016. CO-DEN CSSEEI. ISSN 0267-6192.

Zhang:2015:GAK

[ZLL^{+}15] Shunxiang Zhang, Kui Lu, Wenjuan Liu, Xiaobo Yin, and Guangli Zhu. Generating associated knowledge flow in large-scale web pages based on user interaction. International Journal of Computer Systems Science and Engineering, 30(5):??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Lin:1995:DED

[zLTC95] Ch. zong Lin, Ch.-Ch. Tseng, and Kw.-H. Chi. Design of an efficient data-driven pipelined computer architecture. International Journal of Computer Systems Science and Engineering, 10(3):179–186, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Zheng:2016:OMH

[ZLZL16] Xiaowei Zheng, Jiaxuan Li, Yong Zhang, and Qingkun Liu. An optimization model of hadoop cluster performance prediction based on Markov process. International Journal of Computer Systems Science and Engineering, 31(2): ??, March 2016. CODEN CSSEEI. ISSN 0267-6192.

Zhao:2004:DIC

[ZMMS04] W. Zhao, L. E. Moser, and P. M. Melliar-Smith. Design and implementation of a consistent time service for fault-tolerant distributed systems. International Journal of Computer Systems Science and Engineering, 19(5): ??, September 2004. CODEN CSSEEI. ISSN 0267-6192.

Zomaya:1998:SIS

[Zom98] Albert Y. Zomaya. Special issue: Simulation in parallel and distributed computing environments. International Journal of Computer Systems Science and Engineering, 13 (1):??, January 1998. CO-DEN CSSEEI. ISSN 0267-6192.

Zolfaghari:2006:ISA

[ZS06] Behrouz Zolfaghari and Mohsen Sharifi. The impact of skewed associativity on the replacement traffic in COMA multiprocessors. International Journal of Computer Systems Science and Engineering, 21 (6):??, November 2006. CO-DEN CSSEEI. ISSN 0267-6192.

Zhang:2003:XRR

[ZT03] H. Zhang and F. Tompa. XQuery rewriting at the relational algebra level. International Journal of Computer Systems Science and Engineering, 18(5):??, September 2003. CODEN CSSEEI. ISSN 0267-6192.

Zhou:2008:SIN

[ZXG08] Wanlei Zhou, Yang Xiang, and Minyi Guo. Special issue: Network attacks and defense systems. International Journal of Computer Systems Science and Engineering, 23 (3):??, May 2008. CODEN CSSEEI. ISSN 0267-6192.

Zhao:2010:LCK

[ZXTS10] Geng Zhao, Kefeng Xuan, David Taniar, and Bala Srinivasan. LookAhead continuous KNN mobile query processing. International Journal of Computer Systems Science and Engineering, 25(3): ??, May 2010. CODEN CSSEEI. ISSN 0267-6192.

Zhang:1996:ARB

[ZYM96] Hongyi Zhang, Oliver W.

Yang, and H. Mouftah. Adaptive rate-based congestion control in ATM switching networks. International Journal of Computer Systems Science and Engineering, 11(6):361– 367, November 1996. CODEN CSSEEI. ISSN 0267-6192.

Ziane:1995:PQO

[ZZQ95] Mikal Ziane, Mohamed Zait, and Hong Hoang Quang. Parallelism and query optimization. International Journal of Computer Systems Science and Engineering, 10(1):50– 56, January 1995. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2015:SDR

[ZZWD15] Xiang Zhang, HuaiXiang Zhang, Yifan Wu, and Guojun Dai. A SoPC design of a real-time highdefinition stereo matching algorithm. International Journal of Computer Systems Science and Engineering, 30(5): ??, September 2015. CODEN CSSEEI. ISSN 0267-6192.

Zhang:2015:EEP

[ZZZZ15] Shengqing Zhang, Jiakuo Zuo, Li Zhao, and Cairong Zhu. Energy-efficient power allocation for MIMO twoway cognitive relay networks. International Journal of Computer Systems Science and Engineering, 30(5): ??, September 2015. CODEN CSSEEI. ISSN 0267-6192.